A Complete Bibliography of Publications in SIAM slash ASA Journal on Uncertainty Quantification

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

17 October 2023 Version 1.14

Title word cross-reference

2 [Kue15]. 4 [SSJ^{+}13]. _{2} [HBC^{+}17, PKH22]. A [FH16]. ℓ_{1} [HSH^{+}22]. f [Rah16]. L^{2}(R^{d}, γ_{d}) [SZ23]. p [EEHM14]. R^{d} [LSW17].

-Optimal [FH16]. -Quantiles [EEHM14]. -Regularized [HSH^{+}22]. -Sensitivity [Rah16].

Above [DS17]. Accelerated [CKR20, DMS23, HPS16, JCN16, PM18]. Accelerating [GJWZ16]. Accelerator [Ade19]. Acceptance [DMS23, LDF^{+}23]. Accuracy [CCD16, SAY22, ST15, Spo23]. Accurate [CLW17, CMT18, LG14, MPL16, VM19].

Acoustic [BTG14, FLL15]. Across [CCS17]. Active [CCZW22, CLS^{+}19, Ede23, HNP23, Lee19]. Active-Subspace [Ede23]. Adaptation [BHR18, BBD20, TST18]. Adapted [KJH22, ZKA22]. Adaptive [BCWZ14, BR18, BPR21, CDGR19, CCR19, CKR20, DBK^{+}18, DDS19, EMN16, EM16, EMM20, FH16, GBC^{+}14, HKT20, KS16, MG21, MCWG22, Peh19, TS23, ZCK19, ZHYL17]. Adaptive-Domain [BCWZ14]. Adaptivity [BPW15]. Additive [FBG^{+}20, GCB18, HZ21, KLS15]. Adjustments [Tuo19]. Advanced [BRPP18]. Advection [BST18, Cha15, CCH^{+}20, DEKR18, JT13, JT15, SK22, TBR18]. Advection-Diffusion [Cha15, JT13, JT15]. Advective [GMC22].

Aerodynamic [LLSS17]. Agent [FHC^{+}18]. Agent-Based [FHC^{+}18]. Aided [YLT18]. Airfoil [LLSS17]. Algebraic [KW16]. Algorithm [CGHM14, DC16, DKST15, DKST19, DS19, JCN16, Kou14, LvLP21, TMM20, ZHYL17]. Algorithms [BRPP18, BR18, BCMN18, CMT18, Hos19, MCNT23, RKSA15]. Allen [Kue15]. Allocation [PKNW19]. Alternating [DS19]. American [SWGE22]. Analysis [ABPS14, AUH17, AF19, BR22, BN14, BMP18, BBCM22, BTG14, CCD16, Cha15, CEGT15, CMT18, DKPP16, ESS15, FKL21, GKL18, GKS21, GSM22, HBC19, HLO^{+}18, HSZ19, JLK17, KBS13, KSS^{+}19, Le 13, LCI14, LLS20, Ma20, MCNT23, MS20, MvOCP13, MSDO14, Mor18, NHM^{+}16, OV14, PRB21, SU21, SLH19, SNS16, SL22, SLD19, SS15, Ver17, WR16, WSB16, YY19, dRLI19]. Analytic [SZ23]. Analyzing [Pul13]. Anchored [AKB^{+}23]. Angles [UBD^{+}22]. Anisotropic [DEH^{+}18]. Annealing [CKP23]. ANOVA [AKB^{+}23, BMP18, OH21, Rah14]. Any [Soi17]. APIK [CCZW22]. Appearing [LG14]. Application [BGV16, BW16, BBCM22, BPA20, CRG^{+}15, CKR20, GGS23, LG17, OH21, TWW20, Wal17, WB14, YY19]. Applications [BIM^{+}21, BKM18, CLS^{+}19, DKST15, DKST19, GK19, HPR18, NP17, PG22, Tec20, WYHW22]. Applied [DEH^{+}18]. Approach [AKSAS22, BCWZ14, BGHK20, BPA20, Döl20, LCI14, MDK20, MOM18, OHK^{+}20, SA18, WR16, ZLR19]. Approaches [LLSS17, RW21]. Approximate [EEHM14, GSM22, GNW14, NT18, PG22, PB20, PB21, Spi20, SG15]. Approximating [DG13, RDGS22]. Approximation [AP23, BG15, BKM18, BBDM21, BTG14, CG21, CS23, CLNR15, CDM^{+}18, CFGS20, CLYM20, GP14a, GHT16, GNR15, GM16, HZ21, HLY22, KJ16, KWT^{+}20, KGL22, KKNT15, KN18, LLBDR18, SP16, Spo23, TMM20]. Approximations [ABJN18, APSG17, BCWZ14, BPW15, KLS15, KS16, LSW17, Min22, SNM17]. Arbitrary [vdBSBvB20]. Arising [CKR20, HK18, SP16]. Aspects [ZNX14]. Assembly [WYHW22]. Assessing [RKSA15]. Assimilation [BGV16, BCD^{+}17, Che20, SSJ^{+}13, ST15, WWLL20, YWT20]. Asymptotic [AF19, BCH15, CDGR19, GCB18, HSH^{+}22, SU21]. Asymptotical [LNW21]. Asymptotics [SAS15]. Asynchronous [Ari19]. Atlantic [PAS14]. Atmospheric [BW16, HBC^{+}17, PKH22]. Attention [EDCMD22]. Automata [KMW14]. Automating [SBB^{+}14]. Average [Lee19]. Averse [APSG17, KS18, KS22, ZKA22]. Aware [AP23]. Awareness [HP14].

Background [BGHK20]. Backward [BGM16, BCMZ16, BS13b]. Balance [SLH19]. Balancing [AP23, SAY22]. Based [AMV^{+}17, AKB^{+}23, BCWZ14, BC21, BLS15, BBR^{+}19, BGV16, BLSZ18, BPA20, CSD17, CA23, CPBP20, DM16, FHC^{+}18, GKL18, GP14a, HGS13, HBC^{+}17, LW17, MCNT23, MWP15, MGBL18, MvOCP13, Mor18, MRS^{+}20, OCMB17, PO21, SBBA19, STW23, UPMS21, WSB16, CG22, TMM20, TS23, vdBSBvB20, KSS^{+}19, MCWG22]. Bases [PB23]. Basis [CQ14, CQR17, EL13, HUW13, HL15, JCN16, MPL16, NP17, PMQ17, PSDF14, TST18, TBR18, ZKA22]. Bathymetry [LG17]. Bayes [CCD16, WGD22]. Bayesian [AD20, APSS21, AP23, BLV17, BD23, BS21, BGHK20, BSV18, BTG14, CCD16, CLW17, CHKP19, DMS23, Dia23, DY21, EFF^{+}23, ESS15, EDA23, GP18, GBC^{+}14, GGS23, GK19, HS21, HSAY20, HQS21, HN17, Hos17, HKZ18a, HKZ18b, KLLU19, KBS13, LLS22, Lat20, LZ22, Le 13, LCI14, LST18, Ma20, MPL16, MvOCP13, OR21, PMM16, PM22, PB20, PB21, PZ20, RHH^{+}15, RW21, RHBH20, SBBA19, SN17, SCN21, Spi20, SS15, ST17, Tan15, Tan20, Tec20, TGH17, WCL22, WB14, WCG23, XLWZ21, ZGS22]. Benchmark [LMS21]. Bernstein [GK20]. Best [SU20, SU21]. Between [GM16, Maz21, Min22]. Beyond [CCS17, OVHW19]. BGK [HPW21]. Bias [CS23]. Bias-free [CS23]. Biharmonic [LW22]. Biological [KLLU19]. Bismut [YY19]. Block [BBCM22, KP20]. Block-Diagonal [BBCM22]. Boltzmann [HLPR18, HPW21, ZHY21]. Both [XMLD21]. Bound [HLO^{+}18, LHO^{+}21, TGH17]. Bound-to-Bound [HLO^{+}18, LHO^{+}21]. Boundaries [ADH23, CBE18, HG17]. Boundary [AEST15, KSS^{+}19, PB23, SWGE22, VJC19]. Bounds [ACD15, DKPP16, EMN16, KP20, PS20, RDGS22, Spo23]. Bramble [MUL19]. Brownian [YZAJ20]. Bucy [CJY22]. Build [OGH22]. Burgers [LLS20].

Cahn [KLLU19, Kue15]. Calculation [BS13a]. Calibrating [FHC^{+}18, ZLR19]. Calibration [BCKW22, CD21, DBK^{+}18, GW18, GXW22, KSW14, LYWD23, MvOCP13, RHH^{+}15, RHBH20, SBW22, TW16, Tuo19, TWW20, Wan22]. Can [HS21]. Canonical [MvOCP13]. Carlo [DKST19, BLOP21, AUH17, BC21, BJL^{+}18, BKM18, CS23, CHKW22, CGRF18, DM15, DDS19, DMS23, DLS16, DTVW20, DKST15, DNP21, EMN16, EHM16, GKS21, GP18, GNR15, GSM22, GMN22, GKK^{+}21, HMR^{+}13, HGGT20, HQS21, HPW21, KBJ14, KRS21, KN18, LWZ22, MCNT23, MD19, PM18, Peh19, PB21, QPO^{+}18, SST17, VV19]. Case [CCH^{+}20, EDCMD22, KBJ14, OVHW19]. Categorical [RPD^{+}20]. Cell [ABJR22]. Cellular [KMW14]. Certain [RErB^{+}22]. Certified [EFF^{+}23]. Chain [BC21, DTVW20, DKST15, DKST19, HQS21, LWZ22, MCNT23, PM18]. Chains [HHS22]. Chance [CG21]. Chaos [BGR22, CFGS20, DKLM15, ESS15, FMS17, HSS^{+}18, JCN16, JLK17, LMS21, MS17, OB16, Rah20, Soi15, Tan15, TI14, Ver17]. Chaotic [DS17, FG21, SAS15, WB14, XSW19]. Characterizing [CRG^{+}15, RV20]. Chemical [WR16]. Choice [WGD22]. Choosing [Che20]. Circuit [XSW19]. Circular [PR16]. Circulation [PAS14]. Class [CSD17, HB18, MCNT23]. Classical [DZ21]. Classification [BLSZ18, BKM18, OH21]. Climate [HGS13, TGG13, WB14]. Closures [JT13, JT15]. Cloud [HG17]. Clustering [MRS^{+}20, OHK^{+}20]. CO [HBC^{+}17, PKH22]. Coarse [DG13]. Coarse-Grid [DG13]. Codes [CSD17, DBK^{+}18, Le 13, LCI14, Ma20]. Coefficient [HPS16, WGD22]. Coefficients [AUH17, BST16, BS18, BOS15, CHYZ13, DKLM15, EL13, Kou14, SP16, TST18, VV19]. Cokriging [Ma20]. Collaboration [HLO^{+}18, LHO^{+}21]. Collocation [ABJN18, BST16, DEH^{+}18, EL13, GJWZ16, GIMS19, JS22, Kou14, MMRT16, TJWG15, ZNX14]. Combinations [RKSA15]. Combined [OV14]. Combining [MPL16]. Community [RHH^{+}15]. Comparative [BGR22]. Comparing [OC16]. Comparison [FPGD^{+}16, HLPR18, JT13, JT15, KBS13, KPR22, LLSS17, OCMB17]. Complete [BW23]. Complex [CLS^{+}19, Hes14, VJC19, Wal17, WGD22]. Complexity [BLS15, BDKR22, BJP16, CS23]. Component [MS20]. Components [DGVE19, Owe13]. Composite [WYHW22]. Compressed [ABJN18, MS20, SZ22]. Compression [BQS18]. Compressive [HSS^{+}18, YLT18]. Computation [MD19, PB20, PB21, RA21, SCN21, SNS16, Spi20, VM19]. Computational [AJHSZ20, BET^{+}14, DLS22, EF21, RW21, WCG23, ZNX14]. Computationally [OCMB17]. Computer [BG16, BCKW22, BMP18, CLW17, DBK^{+}18, GBCC15, GW18, HWM18, HC18, HML20, KGL22, KBW18, LCI14, LYWD23, MG21, MT16, SN17, SCN21, SWTA23, SBW22, Tan20, TW16, Tuo19, TWW20, VJC19, Wan22, WGD22, XMLD21, ZLR19]. Computer-Model [KGL22]. Computing [Ari19, EHM16, PRB21, SST17, VCNGP16]. Condition [SWGE22]. Conditional [HKTW18, HKT20, VM19]. Conditional-Value-at-Risk [HKTW18]. Conditioning [GTP14]. Conditions [KS18, KS22, PB23]. Conductivity [HL15, KP18]. Confidence [BPA20, KBS13]. Confinement [OVHW19]. Conjugate [MUL19, OR21]. Connections [GM16]. Conservation [DP16, KRS21, MRST16]. Consistency [HLO^{+}18]. Constrained [CG21, CPBP20, KS18, KS22, KS16, MN21, NvdGW20, SWTA23, WB16, YWT20, ZKA22]. Constraint [CQ14]. Constraints [CDS22, LLBDR18, Sta15]. Constructing [NP19]. Construction [Ade19, BS13a, CLS^{+}19, HKT20, SBB^{+}14]. Contaminant [OV14]. Content [SSJ^{+}13]. Context [ADH23, AP23, SSJ^{+}13, SW23]. Context-Aware [AP23]. Continuation [Dia23, Kue15]. Continuous [DDS19, SS15]. Continuum [GMC22]. Control [APSG17, AUH17, CQ14, EM16, GKK^{+}21, HLR15, KS16, MN21, PG22]. Controlling [HWM18]. Convection [ZHY21]. Converged [PG13]. Convergence [BLS15, BGA^{+}17, BC22, HZ21, KLS15, KM15, Lás16, NvdGW20, Tec20, TW16, TWW20, dHKNS23]. Convergent [ZCK19]. Convex [JKS^{+}21]. Correctability [BFKRB22]. Correction [SHA18]. Corrector [GHT16]. Correlated [MS15]. Correlation [CLW17, KKNT15, MvOCP13].

Corrigendum [KS22]. Corruptions [ABJN18]. Cost [ST15]. Cost-Accuracy [ST15]. Costs [AP23]. Coupled [AEST15, ASG13]. Coupling [CGRF18, DLZ21, KBW18, NP19]. Covariance [BQS18, BBCM22, CS23, GMC22, HS21, Hes13, Min22, MD19, YZAJ20]. Covariances [BGA^{+}17]. Cramér [GKL18]. Criterion [Che20, GP14a]. Cross [DS19, EFF^{+}23, Héa20, HSS^{+}18, PKW18, UPMS21, MCWG22]. Cross-Entropy [EFF^{+}23, Héa20, PKW18]. Cross-Entropy-Based [UPMS21]. Cross-Validation [HSS^{+}18, MCWG22]. Crossing [DPP18]. Cubature [HPR18]. Cubic [BC22, Kue15]. Cubic-Quintic [Kue15]. Cumulative [BST18]. curl [RSW16].

D [Kue15, SSJ^{+}13]. D-Var [SSJ^{+}13].

Damped [CCH^{+}20]. Darcy [BN14]. Data [BGV16, BJP16, BLSZ18, BR18, BCKW22, BCD^{+}17, Che20, CHYZ13, CG22, EMN16, GJWZ16, HK18, HGS13, HLO^{+}18, HBC^{+}17, HKZ18a, HKZ18b, KRS21, LG14, LHO^{+}21, NT18, SLH19, SSJ^{+}13, ST15, SRS17, SPHJ21, TJWG15, WWLL20, XSW19, YWT20, dHKNS23]. Data-Driven [CHYZ13]. Datasets [HLO^{+}18]. Dealing [EF21, RHBH20]. Decision [OV14, TH18]. Decomposing [Mor18]. Decomposition [BPW15, BW23, Rah14, Rah18, RJ22, TST18]. Deep [BCKW22, DMS23, Ede23, LLS22, Sca22, SZ23]. Defensive [WWLL20]. Defined [PSDF14, Soi17]. Deformation [KSW14]. Degenerate [BGA^{+}17]. Deglaciation [SWGE22]. Delay [Ver17]. Delayed [DMS23, LDF^{+}23]. Delayed-Acceptance [DMS23]. Densities [EDA23, GNR15, MDK20, RDGS22]. Density [BLOP21, BST18, DFS20, WBSC^{+}15]. Dependent [BG15, BIM^{+}21, GIMS19, OP17, PB23, Rah14]. Derivative [DM16, MGBL18]. Derivative-Based [DM16, MGBL18]. Deriving [SO14]. Descent [JZZ21, WCL22]. Description [MS15]. Design [APSS21, BG16, BBR^{+}19, BS21, CRG^{+}15, FMS17, FH16, GM16, HLR15, HL17, IDL17, KJH22, MG21, PZ20, SHA18, SN17, SCN21, SPHJ21, TH18, TGL23, WGD22, WCG23, ZLR19]. Designs [DBK^{+}18, GP14a, MGBL18, PR16]. Detection [DC16, dRLI19]. Determinant [SSZR19]. Deterministic [BW23, CFN^{+}23, FPGD^{+}16, HWM18, Hes13, KWT^{+}20]. Deviation [TS23]. Diagnostics [VW20]. Diagnostics-Driven [VW20]. Diagonal [BBCM22]. Difference [KRS21]. Differential [APV17, AAV18, BGM16, BCMZ16, BS18, CG21, CCZW22, CD21, CPBP20, DLZ21, DKLM15, EL13, GJWZ16, GR17, HUW13, HPR18, HQS21, JKS^{+}21, JLK17, MM21, MSDO14, Mot19, OB16, RML18, TJWG15, TST18, Ver17, VCNGP16]. Diffusion [AF19, BS18, BOS15, Cha15, DEH^{+}18, DEKR18, HPS16, HSZ19, HLR15, IDL17, JT13, JT15, NT18, PM22, SK22, SP16, TMM20, UP15b, ZHY21]. Diffusions [CFJ^{+}21, GB18]. Dimension [BS13a, BBR^{+}19, BJL^{+}18, CLS^{+}19, EFF^{+}23, GK19, LG17, OH21, PS17, SZ23, SYP19, Spo23, UPMS21, Vol15]. Dimension-Independent [BJL^{+}18, Vol15]. Dimensional [AMV^{+}17, BLSZ18, BTG14, CG21, CCH^{+}20, ITG19, KBJ14, LLBDR18, Rah14, Rah18, RJ22, RErB^{+}22, Spo23, WCL22, WCG23, dRLI19]. Dimensions [BCL16, CMT18, Hos19, RJ22, TS23]. Directed [BFKRB22]. Disc [DMS23]. Discontinuities [HML20]. Discontinuous [MDK20]. Discrepancies [SO14]. Discrepancy [LHO^{+}21]. Discrete [MUL19, OHK^{+}20]. Discretely [CFJ^{+}21]. Discretization [EMM20, MM21, YY19]. Discretized [HJLT23]. Dispersion [BW16, BC22]. Displacement [RM18]. Distance [BBB^{+}15, GKL18]. Distances [Min22]. Distributed [DPP18, HPS16]. Distribution [BRGT17, BBD20, CBE18, GNR15, HG19, HP14, IT17, SAS15, vAH17]. Distributions [BS13a, BMP18, Oli17]. Disturbance [MJP13]. Divergence [ACD15, TGG13]. Divergences [HK18]. Divisible [Hos17]. Domain [AJHSZ20, BCWZ14, GP14b, TST18]. Domains [CBE18, Dju21, Döl20, EMM20, GP14b, PR16, vWGBS15]. Dominated [TBR18]. Don’t [APSS21]. Doubly [BGM16, BCMZ16]. Drag [SGH^{+}19]. Driven [CHYZ13, VW20]. Dynamic [PAS14, ZLR19]. Dynamical [CRR16, DGVE19, FH16, MS15, OB16, Pul13, PA19, RA21, SAS15, SHA18, SRS17]. Dynamics [ADSG22, Ari19, BPA20, BW23, DKPP16, RPMC21, SHL14].

Early [BHR18]. Easy [XMLD21]. Easy-to-Interpret [XMLD21]. Edge [UBD^{+}22]. Edge-Preserving [UBD^{+}22]. Editors [BEG13]. Effect [YZAJ20]. Effective [SZ22]. Effectively [SNM17]. Effects [BBD20, BBDM21, BBCM22, EDCMD22, KSS^{+}19, PAS14, PRB21, SNS16]. Efficiency [WR16]. Efficient [AMV^{+}17, BR18, CBE18, CMT18, CGRF18, FLL15, GHA19, LG14, NP17, OH21, PMQ17, RA21, SBBA19, TGH17, ZCK19]. Eigenvalue [LS18]. Eigenvector [BQS18]. Electrical [BSS^{+}15, HL15]. Electromagnetic [Döl20]. Electromagnetics [AJHSZ20]. Element [AEST15, BPR21, CSD17, DEKR18, EM16, GKS21, HL15, KLS15, SAY22, SE16, Tan19, UP15b]. Elements [ADSG22, ANSS14]. Elliptic [AD20, AUH17, BS18, BR18, CBE18, CQ14, CHYZ13, GHT16, GP14b, HZ21, HSAY20, RSW16, SST17]. Elworthy [YY19]. Embedded [DEH^{+}18]. Emission [GBD23]. Empirical [BGM16, VCNGP16, ZGS22]. Emulating [SGH^{+}19]. Emulation [BG16, BW16, CLW17, FHC^{+}18, GBCC15, HWM18, KGL22, LG17, MG21, MT16, OC16, OGH22, VJC19, WGD22, ZS21]. Emulator [MWP15, SBB^{+}14]. Emulators [FSK22, HC18, KBW18, OGH22, RS15b, SWTA23, VW20, WB14]. Energy [PZ20]. Enhanced [HC18]. Enough [SO14]. Ensemble [BW23, CJY22, CRR16, CCR19, DEH^{+}18, ESS15, FBG^{+}20, KM15, LWZ22, PMQ17, PG22, STW23, WPU22]. Ensemble-Based [STW23]. Ensembles [HGS13]. Ensembling [BB16]. Entropic [Min22]. Entropy [BS13a, EFF^{+}23, Héa20, PKW18, UPMS21]. Equation [BCWZ14, BST18, CKP23, CD21, CCH^{+}20, CHKW22, HLPR18, HPW21, JKS^{+}21, Kue15, LLS20, LW22, MM21, PS20, RSW16, SK22, XLL14]. Equations [APV17, AAV18, AF19, BGM16, BCMZ16, BST16, BS18, BOS15, BC22, CMT18, CG21, CCZW22, CPBP20, DLS16, DEKR18, DLZ21, DKLM15, EL13, GJWZ16, GR17, GIMS19, HUW13, HZ21, HPR18, HQS21, HKZ18a, HKZ18b, JS22, JT13, JLK17, KP18, KBJ14, KLS15, LES19, LW17, MSDO14, Mot19, MUL19, NT18, OB16, RML18, SP16, TJWG15, TST18, Ver17, VCNGP16, ZHY21, JT15]. Equations-Based [CPBP20]. Equilibration [EM16]. Equilibria [Pul13]. Ergodic [AAV18]. Erratum [JT15, DKST19]. Error [AEST15, BPR21, BGA^{+}17, BPW15, DC15, EM16, FBG^{+}20, GCB18, Lás16, LvLP21, MPL16, MM21, PSDF14, RS15a, SSZR19, WYHW22]. Errors [EF21, SBW22]. Estimate [DLZ21, SSJ^{+}13]. Estimated [Tec20]. Estimates [AEST15, BS13b, Gri17, RS15a]. Estimating [BGHK20, HBC^{+}17, TI14, WB16]. Estimation [BLOP21, ABJR22, BBB^{+}15, BPR21, BCMN18, BBD20, BBDM21, BBCM22, CBE18, CGHM14, CA23, CDMN22, DG19, DM16, DFS20, GB18, GHA19, HKTW18, HKT20, HJLT23, HGGT20, HKZ18a, IDL17, KWT^{+}20, LvLP21, MM21, MD19, OH21, PMQ17, PS17, PKW18, Peh19, QPO^{+}18, SK22, Tan15, TW16, UP15a, WPU22, XS17]. Estimators [BCH15, EM16, Maz21, NvdGW20, PG22, SU20, SU21]. Euler [CCH^{+}20]. Evaluate [TGG13]. Evaluating [AGZ23]. Evaluations [WGD22]. Event [DGVE19, PKW18, UPMS21, WPU22, XLL14]. Events [AGZ23, MS15, TS23, UP15a]. Evidence [SCN21]. Evolution [CEGT15, KLS15, OB16]. Evolve [GIMS19]. Evolve-Filter-Relax [GIMS19]. Evolving [DEKR18, WB14]. Exact [Min22]. Excited [MS15]. Excursion [ABCG16, BPA20, RA21]. Exhibiting [SHL14]. Existence [HLPR18, KS18, KS22]. Exit [GB18, HMR^{+}13]. Expansion [CFGS20, DKLM15, HUW13, Rah20, Soi15]. Expansions [BGR22, FMS17, HSS^{+}18, JLK17, LMS21, MS17, OB16, TI14, Ver17]. Expectation [LvLP21, VM19]. Expectations [SST17]. Expected [GB18, HGGT20, SO14, Tan20, TGH17]. Expensive [CDM^{+}18, OCMB17]. Experience [OVHW19]. Experiment [FMS17]. Experimental [BS21, CRG^{+}15, FH16, GM16, HLR15, KJH22, PR16, SHA18, SPHJ21, TH18, TGL23, WCG23]. Experimentation [TGH17]. Experiments [BG16, BMP18, GBCC15, HWM18, HC18, HML20, OVHW19, SN17, SCN21, SGH^{+}19, Tan20, XMLD21]. Explicit [AAV18, SW23]. Explorations [Maz21]. Exponential [HS21, HN17]. Expression [SZ23]. Extended [CRR16]. Extrapolated [DLS22]. Extremal [RKSA15]. Extreme [DGVE19, Gri16, Gri17, MS15, RA21]. EzGP [XMLD21].

Factors [CCD16, XMLD21]. Failure [DG19, EHM16, PKW18, UPMS21]. Failure-Informed [UPMS21]. Fast [CD21, FN16, KGL22, LYWD23, WCG23]. Fault [ABJN18]. Fault-Tolerant [ABJN18]. Feedback [TMM20]. FEM [SP16]. Fictitious [GP14b]. Fidelity [Peh19, PKNW19]. Field [ABCG16, SG15, XS17]. Fields [APSG17, BK21, DG13, KSS^{+}19, KKNT15, SZ22]. Filling [PZ20]. Filter [BGA^{+}17, CRR16, CCH^{+}20, GIMS19, KM15, LvLP21, PMQ17, RErB^{+}22, TMM20, WPU22]. Filtering [BCWZ14, BD23, HB18, HKZ18a, HKZ18b, RErB^{+}22, SAS15, WWLL20]. Filters [CJY22, CLYM20, ESS15]. Fine [MSDO14]. Fine-Scale [MSDO14]. Finite [ADSG22, AEST15, ANSS14, BPR21, CSD17, DEKR18, EM16, GKS21, HL15, KRS21, KLS15, LLBDR18, Min22, PSDF14, SAY22, SE16, Tan19, UP15b]. Finite-Dimensional [LLBDR18]. First [BCMZ16, GHT16, IDL17]. First-Order [GHT16]. Flexible [CLW17, MCTI16]. Floes [CDS22]. Flow [DKST15, LvLP21, NP17, SHL14, DKST19]. Flows [AS22, BH15, BGHK20, BU21, HHS22, MWP15, SC20]. Fluid [RPMC21, SHL14]. Flux [MRST16]. Fluxes [DMS23]. Fokker [CMT18, RW21]. Forced [CCH^{+}20]. Forcing [OB16, PAS14]. Forecast [GCB18]. Forecasting [HG17]. Forecasts [AGZ23]. Form [HK18]. Format [DKLM15, RDGS22]. Formations [SC20]. Formula [RA21, YY19]. Formulae [vAH17]. Formulas [HPR18]. Formulating [SS13]. Formulations [DP16, GP14b]. Forward [LST18]. Fractional [HZ21, HSZ19, KRS21, NT18]. Fracture [OHK^{+}20]. Fragments [Ari19]. Framework [BB16, GXW22, HWM18, HLY22, JKS^{+}21, MCTI16, TW16, WCG23, XLWZ21]. Frameworks [FPGD^{+}16]. Free [CLS^{+}19, Jah23, Spo23, CS23]. Frequency [HKZ18a, HKZ18b, KP18, MMRT16]. Frequentist [KBS13, PKH22]. Full [DY21]. Function [BST18, MSDO14, TMM20, WBSC^{+}15, ZHYL17]. Functional [BRB20, BMP18, FSK22, Kre23, Mor18, NHM^{+}16, Tan19]. Functionals [ACD15, GB18]. Functions [BGR22, CS23, CLNR15, GNR15, GTP14, HPR18, KWT^{+}20, KKNT15, KN18, LK22, MGBL18, MRST16, SZ23, SYP19, TGG13, WB16, vAH17]. Fusion [OVHW19]. Fuzzy [Mot19]. Fuzzy-Stochastic [Mot19].

Gain [TMM20, TGH17]. Galerkin

[AKB^{+}23, BPR21, EM16, EMM20, HL15, KP20, KS16, LCE18, Pul13, PA19, SHL14, SL22, UP15b]. Gamma [AKSAS22, RKSA15]. GAN [PO21]. GAN-Based [PO21]. Gauss [HP14, TI14]. Gaussian [AS22, ABCG16, BBR^{+}19, BQS18, BSDGH23, BBDM21, CLW17, DG13, DM16, FN16, GM16, GNW14, GW18, GXW22, HWM18, HS21, HZ21, HC18, HQS21, Hos19, KWT^{+}20, KKNT15, LK22, LG17, LLBDR18, LSW17, MGBL18, MS20, MG21, Min22, MCWG22, MT16, OR21, PR16, PG13, RDGS22, RPD^{+}20, SAY22, SZ22, SYP19, Soi17, SWTA23, Tan19, Tec20, TGL23, Vol15, WB16, WYHW22, XMLD21, XS17, vAH17]. General [DTVW20, HLY22, Jah23, MCTI16, vWGBS15]. Generalized [BR22, CKP23, CHKP19, GGS23, JCN16, Owe13, PM22, Rah14, ZS21]. Generalizing [Ari19]. Generated [SC20]. Generating [vdBSBvB20]. Generation [DPP18, SZ22]. Generative [AS22]. Geological [SC20]. Geometric [MDK20]. Geometrical [ITG19]. Geometry [LLSS17, Spi20]. Geometry-Induced [LLSS17]. Geophysical [MWP15, SBB^{+}14]. Gfrerer [JM13]. Gfrerer-Type [JM13]. Gibbs [ABPS14, DY21, UBD^{+}22]. Girsanov [WR16]. Global [BR22, DM16, FKL21, HBC19, HGS13, LCI14, Maz21, SNS16, SLD19, WSB16]. Globally [ZCK19]. Goal [BBR^{+}19, EDCMD22, JCN16]. Goal-Oriented [BBR^{+}19, EDCMD22, JCN16]. Good [APSS21, SO14]. Governed [APSG17, CG21]. Gradient [BIM^{+}21, BGV16, BH15, CRG^{+}15, CLS^{+}19, HC18, JZZ21, MUL19, STW23, WCL22, vAH17]. Gradient-Enhanced [HC18]. Gradient-Free [CLS^{+}19]. Gradients [CD21, Lee19]. Graph [BLSZ18]. Graph-Based [BLSZ18]. Graphical [BFKRB22]. Greedy [BCMN18]. Green [MSDO14]. Grid [BCWZ14, DG13]. Grids [ZCK19]. Groundwater [NP17]. Group [RPD^{+}20]. Grouping [DEH^{+}18]. Growth [KLLU19]. Guaranteed [EM16, KP20].

Hamiltonian [DMS23]. Harmonic [KP18]. Hastings [Hos19, MCNT23, MJP13]. Having [CS23]. Hazard [SBB^{+}14, SS13]. Heat [DMS23]. Heavy [Hos17]. Heavy-Tailed [Hos17]. Heights [LG17]. Helmholtz [PS20, SW23]. Heterogeneous [GKS21, Gri14, MGBL18, SC20]. Heteroscedastic [SBW22]. Hidden [CFJ^{+}21, CGHM14]. Hierarchical [ABPS14, BG15, BC21, BSV18, CA23, DKST15, DKST19, Le 13, Ma20, OC16, SBBA19]. Hierarchies [DDS19]. High [AGZ23, AMV^{+}17, BS13a, BLSZ18, CG21, KBJ14, KKNT15, MMRT16, PKNW19, RJ22, RErB^{+}22, SZ23, Spo23, TS23, WCL22, WCG23]. High-Dimensional [CG21, KBJ14, RErB^{+}22, Spo23, WCL22, WCG23]. High-Frequency [MMRT16]. High-Impact [AGZ23]. High-Order [KKNT15]. Higher [DLS16, Döl20, EB21, EM16, GP18]. Higher-Order [EM16, GP18]. Highly [DM15, VM19]. Hilbert [DZ21, GR17].

Hilliard [KLLU19]. History

[AMV^{+}17, CG22, DNP21]. Hitting [IDL17]. HIV [AMV^{+}17, WSB16]. hoc [BHTG21]. Holomorphic [DLS16]. Holomorphy [SW23]. Homogenization [AD20]. Hybrid [ASG13, BCWZ14, CRR16, DS19, UBD^{+}22, ZHYL17]. Hyper [Tec20]. Hyper-Parameters [Tec20]. Hyperbolic [RML18, SLH19]. Hyperpriors [AKSAS22]. Hypocoercivity [LW17].

Ice [BPA20, CDS22]. Ice-Sheet [BPA20]. Identification [HSH^{+}22, Hes14, KLLU19, LG14]. Identifying [JKS^{+}21, OHK^{+}20]. II [HKZ18b]. III [BET^{+}14]. Illustrated [WSB16]. Image [CG22, GGS23]. Impact [AGZ23, ST15]. Impacts [RV20]. Impedance [BSS^{+}15, HL15]. Implications [BMP18]. Implicit [ST17]. Implosion [OVHW19]. Importance [AP23, CCR19, CKR20, FG21, OP17, PG22, SA18, SSZR19, TS23, UPMS21]. Improved [BDW13, HP14, Oli17, TWW20]. Improvement [SO14]. IMSE [GP14a]. Inaccuracy [HWM18]. Inadequacy [MOM18]. Including [YZAJ20]. Inclusion [ADH23]. Incompressible [CCH^{+}20]. Inconsistent [HLO^{+}18]. Independence [BSV18]. Independent [BJL^{+}18, MCNT23, Soi17, Vol15]. Index [RB19, Rah16]. Indicator [BGR22]. Indices [HG19, KPR22, Maz21, MD19, Owe13, Owe14, QPO^{+}18, Tan15]. Individual [AMV^{+}17]. Individual-Based [AMV^{+}17]. Induced [LLSS17, Maz21]. Inequalities [BRB20, BU21]. Inequality [JM13, LLBDR18, Wal17, vAH17]. Inertial [OVHW19]. Inexact [BGV16, LS18, SPHJ21, SBW22]. Inference [AKSAS22, BBB^{+}15, CCS17, CFJ^{+}21, DMS23, Dia23, EDA23, HKZ18a, HKZ18b, PMM16, PM22, RW21, STW23, SSZR19, WCL22, ZGS22]. Infinite [BTG14, Hos19, HML20]. Infinitely [Hos17]. Influence [LG17]. Influences [HUW13]. Influential [LG14]. Information [BG16, BU21, Che20, DKPP16, HK18, HGGT20, SA18, SSJ^{+}13, Spi20, SG15, TGH17]. Information-Theoretic [SG15]. Informative [DM15, OGH22]. Informed [BJL^{+}18, CCZW22, UPMS21, ZGS22]. Initial [BJP16, SLH19]. Input [BBD20, GJWZ16, LG14, LW17, TJWG15, WBSC^{+}15, WYHW22]. Inputs [EEHM14, Mor18, NHM^{+}16, OP17, RPD^{+}20, Tan19, TBR18]. Instability [CGHM14, HC18]. Instances [EF21]. Integrated [GM16]. Integration [DLS16, DLS22]. Integrator [AAV18]. Interaction [BC22, RB19]. Interest [BET^{+}14, EMN16, Sca22]. Intermediate [OGH22]. Intermittent [CGHM14]. Intermittently [MS15]. Internal [OGH22, SO14, Tan20]. Interpolation [GP14a, KJ16, RM18, VCNGP16]. Interpolatory [DG13]. Interpret [XMLD21]. Intrusive [DP16]. Inverse [AD20, ABPS14, AKSAS22, APSS21, AP23, BCL16, BC21, BQS18, BHR18, BSV18, BTG14, BET^{+}14, CHKP19, CPBP20, CKR20, ESS15, FN16, GK20, HHS22, HSAY20, HN17, Hos17, JKS^{+}21, Jah23, JZZ21, KBJ14, LLS22, Lat20, LW22, LST18, LNW21, MPL16, NT18, RS15a, SBBA19, SST17, SN17, SE16, SRS17, ST17, Tec20, TGL23, Vol15, YLT18]. Inversion [BW23, DY21, GP18, HQS21, OR21]. Issue [HC18]. Iteration [BLS15, SE16]. Iterative [CEGT15, FBG^{+}20, Lás16, PP23].

Jacobian [SSZR19]. Joint [SK22]. Jump [BS18].

Kalman [BD23, BGA^{+}17, BW23, CJY22, ESS15, FBG^{+}20, KM15, PMQ17, WPU22]. Karhunen [HUW13]. Kernel [AGZ23, BR22, DZ21, GP14a, GR17, HS21, HSAY20, KJ16, Tuo19, TWW20, VW20, Wan22]. Kernel-Based [GP14a]. Kernels [GM16, MG21, RPD^{+}20]. Kinetic [AF19, DP16, LW17]. Kinetics [WR16]. Know [APSS21]. Knowledge [CRG^{+}15]. Known [VJC19]. Korteweg [XLL14]. Kriging [CSD17, CCZW22].

Lagged [RErB^{+}22]. Lagrange [BS13a, SG15]. Lagrangian [BU21]. Lambda [ZS21]. Land [RHH^{+}15]. Landmark [FSK22]. Landmark-Warped [FSK22]. Langevin [ADSG22, CKP23]. Laplace [Spo23]. Large [APSS21, CMT18, KM15, PP23, RKSA15, SGH^{+}19, TS23, TGL23, WB14, WCG23]. Large-Scale [PP23, TGL23, WCG23, APSS21]. Latent [LK22]. Lattice [DLS22, ZHY21]. Law [vWGBS15]. Laws [DP16, KRS21, MRST16, SLH19]. Learning [BCKW22, CRG^{+}15, GGS23, GBD23, HNP23, PO21, SZ23, dHKNS23]. Least [CLNR15, DS19, HNP23, LCE18, NvdGW20, PP23, RKSA15, SNM17]. Least-Squares [CLNR15, LCE18]. Level [AKB^{+}23, BPR21, DDS19, Jah23]. Level-Set-Based [AKB^{+}23]. Levenberg [BGV16, BDKR22]. Lévy [KLS15]. Li [YY19]. Lie [CCH^{+}20]. Lifted [YZAJ20]. Likelihood [BJL^{+}18, CA23, KWT^{+}20, Oli17, XS17]. Likelihood-Informed [BJL^{+}18]. Likelihoods [DS17, LST18]. Limit [AF19, BKM18, CCS17, KM15, SLH19]. Limited [SPHJ21, WGD22]. Limits [CFGS20]. Line [SC20]. Line-Sources [SC20]. Linear [APSS21, AF19, BQS18, BD23, BBDM21, BW23, DG19, Dju21, FH16, FN16, GK20, Jah23, JZZ21, KLS15, Lás16, LCE18, LW17, LLBDR18, LNW21, RKSA15, SU20, SU21, SP16, SL22, SRS17, TGL23, WGD22, dHKNS23]. Linear-Gaussian [FN16]. Linked [MG21].

Linking [KBW18]. Literature [LMS21]. Local [CDM^{+}18, EM16, GMN22, HL15]. Locally [ZKA22]. Locating [HML20]. Location [WYHW22]. Locations [Che20]. Loève [HUW13]. Log [BN14, HPS16, LST18, UP15b]. Log-Likelihoods [LST18]. Log-Normal [BN14]. Log-Normally [HPS16]. Log-Transformed [UP15b]. Long [OB16, SAS15]. Long-Time [SAS15]. Loss [DY21, GMC22, Tan20]. Low [BOS15, BQS18, BSV18, CLNR15, ITG19, KP18, KKNT15, LES19, Peh19]. Low-Dimensional [ITG19]. Low-Fidelity [Peh19]. Low-Frequency [KP18]. Low-Rank [BOS15, BQS18, BSV18, KKNT15, LES19]. Lower [TGH17]. Lurking [dRLI19].

Machlup [Kre23]. Macroscale [DC16]. Managing [CFN^{+}23]. Manifolds [HSAY20]. Many [BJP16]. Map [CKR20, PM18, TMM20]. Map-based [TMM20]. Maps [PMM16, SBB^{+}14]. Marginal [HG19, MJP13, WWLL20]. Marginalization [SBBA19]. Marginalization-Based [SBBA19]. Markov [DKST19, BC21, BK21, BRB20, CFN^{+}23, CFJ^{+}21, DTVW20, DKST15, HHS22, HQS21, LWZ22, MCNT23, NP19, PM18, TH18]. Marquardt [BGV16, BDKR22]. Mass [MWP15]. Massive [LYWD23]. Massively [GNW14, HLO^{+}18]. Matching [AMV^{+}17, DNP21]. Material [Gri16]. Matérn [KSS^{+}19, MG21, TWW20]. Matérn-Type [TWW20]. Mathematical [Rah18]. Matrices [BQS18, CA23]. Matrix [GTP14]. Maxima [Gri17]. Maximization [LvLP21]. Maximum [BS13a, CA23, KWT^{+}20, Oli17, XS17]. Maxwell [KP18]. McDiarmid [Wal17]. MCMC [BIM^{+}21, CGHM14, CDM^{+}18, LDF^{+}23, SBBA19, Vol15, ZHYL17]. Mean [APSG17, HMR^{+}13, OH21]. Mean-Variance [APSG17]. Measure [BCH15, BET^{+}14]. Measure-Theoretic [BET^{+}14]. Measurement [RHBH20, SSZR19]. Measurements [BCMN18]. Measures [DM16, Hos17, Hos19, HB18, LSW17, Maz21, Rah14]. Measuring [OP17]. Media [Cha15, FLL15, GKS21, GK19, PS20]. Mercer [GM16]. Meridional [PAS14]. Meshes [CGRF18, vWGBS15]. Message [BEG13]. Metamodel [DM16]. Metamodeling [NHM^{+}16]. Metamodels [RPD^{+}20]. Method [AD20, ASG13, AKB^{+}23, BGM16, BR22, BDKR22, BPR21, BET^{+}14, CQ14, CHYZ13, CS23, CLNR15, Dia23, DTVW20, DC15, DNP21, Ede23, EFF^{+}23, EMN16, EHM16, FLL15, GIMS19, GKK^{+}21, Héa20, HMR^{+}13, HL15, ITG19, JS22, KW16, KMW14, LG14, Lás16, LCE18, MUL19, OC16, PKW18, SHL14, SSJ^{+}13, TJWG15, VV19, VCNGP16, WBSC^{+}15, WWLL20, YLT18, ZCK19, ZHY21, ZKA22]. Methodology [HG17, MvOCP13]. Methods [APV17, AEST15, BST16, BST18, BC21, BGV16, CQR17, CHKW22, DG13, DEH^{+}18, DEKR18, DLZ21, DPP18, EM16, EL13, GJWZ16, HUW13, HSAY20, HLPR18, HPW21, JCN16, KBJ14, KBS13, KRS21, LS18, MPL16, NP17, OCMB17, Pul13, SBBA19, SST17, STW23, SE16, SL22, ST17, TBR18, UP15b]. Methods-Accelerated [JCN16]. Metrics [BRPP18]. Metropolis [Hos19, MCNT23, MJP13]. Metropolized [Oli17]. MICE [BG16]. Microcracks [DC16]. Microstructure [Gri16]. Microstructures [Gri14]. Milky [BBB^{+}15]. Mini [RS15b]. Mini-Minimax [RS15b]. Minimax [RS15b]. Minimization [PZ20]. Minimizers [Kre23]. Minimum [BBB^{+}15]. Misaligned [FSK22]. Mises [GKL18, GK20, HP14]. Misspecified [SHA18]. Mixed [AEST15, ANSS14, UP15b]. Mixtures [VW20]. Model [Ade19, APSS21, ABJR22, AMV^{+}17, BBB^{+}15, BCH15, BG16, BB16, BCKW22, BFKRB22, BD23, CCS17, CFJ^{+}21, CLW17, CCZW22, DDS19, DFS20, DC15, FHC^{+}18, FBG^{+}20, GK19, GCB18, GW18, HK18, HBC19, HKT20, Hes13, HKZ18a, HKZ18b, HPW21, KGL22, KBS13, KSW14, LHO^{+}21, LvLP21, Ma20, MvOCP13, Mor18, MOM18, PG13, RHH^{+}15, RV20, RML18, RHBH20, SCN21, SO14, SPHJ21, Tan15, VCNGP16, WSB16, WGD22, XS17, YWT20, ZCK19]. Model-Form [HK18]. Model-Variance [VCNGP16]. Modeling [Ade19, AP23, AS22, BCD^{+}17, DC15, HGS13, Ker23, OC16, PB23, RSW16, Tan19, XSW19]. Models [BRGT17, BGV16, BDKR22, BCMN18, BFKRB22, BDW13, CD21, CDMN22, CLS^{+}19, CDM^{+}18, CPBP20, DG19, EEHM14, FSK22, GP14a, Gri16, Gri17, Héa20, HGS13, HKTW18, HJLT23, Hes14, KLLU19, KP18, KBW18, LST18, LYWD23, MS17, MPL16, MM21, Maz21, MG21, MCWG22, MJP13, OR21, Peh19, PKNW19, RPMC21, RErB^{+}22, SHA18, ST15, SWTA23, SG15, SS15, SBW22, SSZR19, Tan19, TGG13, TW16, Tuo19, TWW20, VJC19, Wan22, WB14, WGD22, XMLD21, YZAJ20, ZNX14, ZS21]. Modes [CHKP19, Kre23]. Modified [Lee19]. Molding [PT17]. Moment [DZ21]. Monotone [GBCC15, RM18]. Monotonic [BKM18]. Monte [DMS23, DTVW20, DKST19, PM18, BLOP21, AUH17, BC21, BJL^{+}18, BKM18, CS23, CHKW22, CGRF18, DM15, DDS19, DLS16, DKST15, DNP21, EMN16, EHM16, GKS21, GP18, GNR15, GSM22, GMN22, GKK^{+}21, HMR^{+}13, HGGT20, HQS21, HPW21, KBJ14, KRS21, KN18, LWZ22, MCNT23, MD19, Peh19, PB21, QPO^{+}18, SST17, VV19]. Most [LG14]. Moving [Dju21]. Multi [BB16]. Multi-Model [BB16]. Multifidelity [AP23, BPA20, Ker23, Le 13, LCI14, Ma20, PKW18, Peh19, PKNW19, PB20, PB21, QPO^{+}18, ST15, ZNX14, PG22]. Multilevel [AUH17, BLS15, BJL^{+}18, BPR21, CDGR19, CJY22, CS23, CGRF18, DKST15, DKST19, EMN16, EHM16, GNR15, GB18, GHA19, GSM22, GMN22, HPS16, HMR^{+}13, HGGT20, HQS21, HPW21, JS22, KRS21, Kou14, KN18, LDF^{+}23, MCNT23, MD19, SU20, SU21, SST17, TJWG15, UP15a, VV19]. Multimodal [Oli17, Soi15]. Multioutput [LK22]. Multiphysics [MCTI16]. Multiple [BMP18, BET^{+}14, HGS13]. Multipliers [BS13a, SG15]. Multipoint [IT17]. Multiresponse [Tan20]. Multiscale [APV17, AD20, CKR20, GHT16, MSDO14, PS17, PMM16]. Multisymmetric [HPR18]. Multivariate [BW16, CLNR15]. Mutual [BG16].

Nanoemulsion [CRG^{+}15]. Natural [SS13]. Navier [GIMS19, KBJ14, LES19]. Nearly [CS23]. Necessary [SA18]. Negative [LK22]. Nested [BSDGH23, GHA19, GSM22, KPR22, OC16, vdBSBvB20]. Net [OH21]. Network [Sca22, SZ23, SS15]. Networks [CKR20, HNP23, LLS22, OHK^{+}20]. Neural [LLS22, OH21, Sca22, SZ23]. Neutron [CHKW22]. Neutronics [CLS^{+}19]. Newton [SHL14]. Nodes [TI14]. Noise [BC22, CGRF18, HZ21, Jah23, KLS15, LNW21, Tan20]. Noises [vWGBS15]. Noisy [HSH^{+}22, dHKNS23]. Non [AS22, Hos19, MS20, Soi17, Vol15]. Non-Gaussian [AS22, Hos19, MS20, Soi17, Vol15]. Nonasymptotic [Spo23]. Nonautonomous [CEGT15]. Nonintrusive [Ade19, HLPR18, JLK17]. Nonlinear [AS22, BCWZ14, BST16, BC21, CRR16, CDMN22, HLPR18, IT17, NT18, PMQ17, RSW16, RKSA15, VM19]. Nonlinearity [BC22]. Nonlocal [BC22]. Nonnested [CGRF18]. Nonparametric [BCH15, BH15, GBD23, SHA18, XLWZ21]. Nonprobabilistic [OV14]. Nonsmooth [Sca22]. Nonstandard [vWGBS15]. Nonstationary [MT16, PG13, VW20]. Normal [BN14, GSM22]. Normality [CDGR19]. Normalizing [HHS22]. Normally [HPS16]. North [SWGE22]. Novel [HG17, KMW14]. Nugget [YZAJ20]. Nuisance [RHBH20]. Numerical [AD20, BTG14, Cha15, DBK^{+}18, FLL15, HZ21, Kue15, MRST16, NHM^{+}16, SAY22].

Objective [Ma20, PKH22]. Observables [DC16]. Observation [Che20, EF21, HGS13]. Observation-Based [HGS13]. Observations [DM15, LYWD23]. Observed [CFJ^{+}21, PS17, SAS15, SK22, TH18]. Observing [BHTG21]. ODEs [CCD16, WBSC^{+}15]. Off [Maz21]. Online [LZ22, SK22]. Onsager [Kre23]. onto [BGA^{+}17]. Operator [DLS16, MOM18, PM22]. Operators [Min22, dHKNS23]. Optima [GBC^{+}14]. Optimal [AAV18, APSG17, APSS21, AUH17, BLS15, BCMN18, BHR18, BS21, CCD16, CQ14, CRG^{+}15, CS23, CLYM20, EF21, FH16, GP14a, GHT16, GKK^{+}21, HNP23, HZ21, IDL17, KJH22, MN21, NP19, PSDF14, SK22, SP16, Soi17, WCG23]. Optimality [KS18, KS22]. Optimization [AKB^{+}23, BC21, CG21, JKS^{+}21, Kou14, KS18, STW23, SLD19, Tan20, VV19, ZCK19, ZKA22, KS22]. Optimization-Based [BC21]. Optimize [BSS^{+}15]. Optimizers [GBC^{+}14]. Oracle [JM13]. Order [AAV18, BCMZ16, DG19, DLS16, Döl20, DC15, EB21, EM16, GP18, GHT16, GIMS19, HKTW18, HKT20, KKNT15, PB23, RPMC21, YY19]. Ordinary [MM21]. Oriented [BBR^{+}19, EDCMD22, JCN16]. Oscillatory [MS17]. Other [GB18]. Outer [HB18]. Output [BG15, BS21]. Output-Weighted [BS21]. Outputs [Ker23, VCNGP16]. Overlapping [MRS^{+}20]. Overturning [PAS14]. Owen [RB19].

Parabolic

[AEST15, BST16, Dju21, HQS21]. Parallel [Ari19, CDM^{+}18, GNW14]. Parameter [APSG17, ABJR22, BG15, CCS17, CGHM14, CDMN22, JKS^{+}21, KLLU19, PMQ17, PB21, SK22, SSZR19, WSB16]. Parameter-Dependent [BG15]. Parameterization [CDS22]. Parameterized [GP14b, HUW13, LCE18]. Parameters [BBB^{+}15, CG21, HBC19, LG14, MN21, MvOCP13, RHBH20, Tec20]. Parametric [BPW15, DLS16, DS19, GR17, HSZ19, SL22, SW23, WR16]. Parametrization [TW16]. Parametrized [TBR18, VCNGP16]. Partial [BS18, CG21, CCZW22, DKLM15, EL13, GJWZ16, GR17, HUW13, HPR18, HQS21, JKS^{+}21, MSDO14, Mot19, RML18, TJWG15, TST18, VCNGP16]. Partially [PS17, PG13, SAS15, SK22, TH18]. Particle [Ade19, BJP16, CRR16, CCH^{+}20, LvLP21, MJP13, RW21, RErB^{+}22, STW23, TMM20, WWLL20]. Partition [Soi17]. Parts [WYHW22]. Pasciak [MUL19]. Passive [BGHK20]. Path [DKPP16]. Path-Space [DKPP16]. Pathwise [FG21]. PDE [KS22, ANSS14, CQ14, GHT16, HK18, HSH^{+}22, KSS^{+}19, KS18, KS16, MN21, NvdGW20, ZKA22]. PDE-Based [KSS^{+}19]. PDE-Constrained [KS22, KS18, KS16, MN21, ZKA22]. PDEs [APSG17, AUH17, BG15, BR18, BPW15, CHYZ13, DS19, GP14b, HPS16, Kou14, LG14, PMQ17, PB23, VV19]. PDMPs [Ari19]. Penalization [ANSS14]. Penalized [NvdGW20, Wan22]. Periodic [SHL14]. Permeability [BN14]. Perturbation [BN14, Döl20]. Perturbed [EMM20, NT18]. Petrov [AKB^{+}23, LCE18]. Phase [CG22, IT17]. Phenomenon [JZZ21].

Photoacoustics [RV20]. Physical [EEHM14, LYWD23, MRS^{+}20]. Physics [CSD17, CCZW22, CDS22, CA23, MWP15]. Physics-Based [CSD17, CA23, MWP15]. Physics-Informed [CCZW22]. Piecewise [CFN^{+}23]. Piecewise-Deterministic [CFN^{+}23]. Placement [SK22]. Planck [CMT18, RW21]. Plate [BW16]. Point [NP17]. Poisson [HLPR18]. Polar [PR16]. Policies [SS13]. Policy [BLS15, CRG^{+}15]. Polynomial [BGR22, DLS22, DKLM15, ESS15, FMS17, GM16, HSS^{+}18, JCN16, JLK17, LMS21, MS17, OB16, Rah18, SNM17, Soi15, Tan15, TI14, Ver17]. Polynomial-Type [BGR22]. Polynomials [DZ21]. Porous [Cha15]. Posed [HN17, Hos17]. Posedness [PS20, Lat20]. Positive [vdBSBvB20]. Post [BHTG21]. Posterior [BQS18, GBD23, Hos19, Kre23, SST17]. Posteriori [AEST15, BS13b, BPR21, CEGT15, PSDF14, RS15a]. Posteriors [DY21]. Potential [JS22]. Power [WBSC^{+}15, vWGBS15]. Power-Law [vWGBS15]. Practical [PP23, SSJ^{+}13]. Preconditioner [GP14b]. Preconditioning [KP20, PKW18]. Predicting [PAS14, XSW19]. Prediction [BRGT17, CLW17, Che20, GW18, GXW22, OVHW19]. Predictive [DC16]. Presence [FBG^{+}20, GCB18, LNW21]. Preservation [PA19]. Preserving [AF19, UBD^{+}22]. Primary [OHK^{+}20]. Principal [MS20]. Principles [BRB20]. Prior [ABCG16, HQS21, Hos17, OR21, ZGS22]. Priori [PS20]. Priors [HS21, HN17, Hos19, PO21, Sta15, TGL23, Vol15]. Probabilistic [BGV16, CD21, KMW14, MS15, OHK^{+}20, WBSC^{+}15, OV14]. Probabilities [EHM16, RA21, RKSA15]. Probability [BCH15, BST18, DG19, LSW17, Maz21, PKW18, Rah14, vAH17]. Problem [BN14, DZ21, FN16, JKS^{+}21, LW22, SWTA23, SRS17]. Problems [AD20, ABPS14, AKSAS22, APSS21, AP23, AEST15, ASG13, BCWZ14, BCL16, BC21, BQS18, BHR18, BSV18, BET^{+}14, CEGT15, CBE18, CQ14, CG21, CHKP19, CPBP20, CKR20, DEH^{+}18, DFS20, Dju21, Döl20, ESS15, GK20, HHS22, HSAY20, HN17, Hos17, Jah23, JZZ21, KBJ14, KP20, KS16, LLS22, Lat20, LS18, LST18, LNW21, MPL16, MN21, NT18, NP17, NvdGW20, RS15a, RKSA15, SBBA19, SST17, SN17, ST17, Tec20, TBR18, TGL23, UP15b, Vol15, VM19, ZKA22]. Process [BBR^{+}19, BSDGH23, CLW17, DM16, GM16, GNW14, GW18, GXW22, HWM18, HS21, HC18, KWT^{+}20, LG17, MG21, MvOCP13, MCWG22, PT17, PG13, RPD^{+}20, SWTA23, Tan19, Tec20, TGL23, WYHW22, XMLD21]. Process-Based [BBR^{+}19, MvOCP13]. Processes [BRB20, CFN^{+}23, HLR15, IDL17, LK22, MGBL18, Min22, MS15, MT16, OGH22, PR16, PS17, PSDF14, SAY22, TH18, WB16, WYHW22]. Profile [GBC^{+}14]. Progressive [DG19]. Projected [EDA23, Tuo19, Wan22, WCL22]. Projections [PA19]. Propagating [Hes13]. Propagation [BDW13, DP16, DFS20, EPS22, GKS21, GMC22, GK19, GMN22, KMW14, Lás16, MMRT16, MCTI16, YWT20]. Proper [TGG13]. Properties [BCH15, Rah18, TW16]. Proposals [BJL^{+}18].

Quadratic [APSG17, Tan20]. Quadrature [HPS16, KW16, PZ20, TI14, vdBSBvB20]. Quadratures [SNM17]. Qualitative [XMLD21]. Quality [DLZ21]. Quantification [Ade19, ADH23, AJHSZ20, BSS^{+}15, BH15, BLSZ18, BK21, BRB20, BS21, BU21, BHTG21, CBE18, CLW17, CQR17, CDS22, CG22, CFGS20, DGVE19, DLS22, DKST15, DKST19, DKPP16, EEHM14, FPGD^{+}16, GK20, GMN22, HS21, HSZ19, HBC^{+}17, HPW21, HLY22, KP18, KWT^{+}20, KW16, LLS22, LLSS17, MPL16, MM21, MRS^{+}20, OCMB17, PKH22, PP23, RB19, RJ22, RS15b, Sca22, SC20, SW23, TGL23, XLWZ21, YLT18]. Quantified [OHK^{+}20]. Quantifying [ABCG16, CFN^{+}23, PO21, RPMC21, SWGE22]. Quantile [BCKW22, BPA20, EDCMD22, FHC^{+}18]. Quantile-Based [BPA20, FHC^{+}18]. Quantile-Oriented [EDCMD22]. Quantiles [EEHM14]. Quantitative [RV20, XMLD21]. Quantities [BG15, BET^{+}14, EMN16, Sca22]. Quasi [BLOP21, DLS16, GKS21, GP18, GKK^{+}21, SST17]. Quasi-Monte [BLOP21, DLS16, GKS21, GP18, GKK^{+}21, SST17]. Quintic [Kue15].

Random [APSG17, AUH17, ABCG16, BCL16, BST16, BDKR22, BJP16, BR18, BK21, BPA20, CG21, CHYZ13, DG13, DGVE19, DEKR18, Dju21, DKLM15, Döl20, EL13, FLL15, GJWZ16, GKS21, GSM22, GK19, Gri14, HK18, JKS^{+}21, JS22, KSS^{+}19, KRS21, KKNT15, LG14, LW17, LW22, LST18, MRST16, MUL19, OB16, PS20, Pul13, RKSA15, SLH19, SP16, Soi15, Soi17, SG15, SSZR19, TJWG15, TST18, TBR18, UP15b, VV19, XS17, vAH17]. Randomize [BSS^{+}15]. Randomize-Then-Optimize [BSS^{+}15]. Randomized [BLOP21, Oli17, PP23]. Randomly [EMM20, NT18]. Range [OVHW19, SWTA23]. Range-Constrained [SWTA23]. Rank [BOS15, BQS18, BSV18, CLNR15, KKNT15, LES19, RDGS22]. Ranking [HL17]. Rare [PKW18, TS23, UP15a, UPMS21, WPU22, XLL14]. Rare-Event [XLL14]. Rates [BC22, NvdGW20, SZ23, TWW20, dHKNS23]. Raus [JM13]. Ray [ADH23]. Reaction [BST18, CKR20]. Realism [HP14]. Realizations [PSDF14, Soi17]. Rearrangement [RM18]. Reconstruction [CG22, GBC^{+}14, GGS23, UBD^{+}22]. Recovery [EF21]. Reduced [BCD^{+}17, BCMN18, CQ14, CQR17, CDMN22, DG19, DC15, EL13, GIMS19, HUW13, Héa20, HKTW18, HKT20, HPW21, JCN16, MPL16, NP17, PMQ17, PB23, PSDF14, RPMC21, TBR18, ZKA22]. Reduced-Basis [ZKA22]. Reduced-Order [HKTW18, HKT20, PB23]. Reduced-Order-Model [DC15]. Reducible [APSS21]. Reduction [BIM^{+}21, BBR^{+}19, BJP16, BBD20, CG22, CLS^{+}19, DTVW20, EFF^{+}23, GK19, LG17, MPL16, PS17, RML18, SYP19, UPMS21, VCNGP16, ZCK19]. Refinement [DG19]. Regime [PKNW19]. Regional [HGS13]. Regression [ANSS14, BGM16, BCKW22, BSDGH23, GM16, GNW14, KJ16, LZ22, NvdGW20, Tec20, TWW20, YLT18]. Regularity [LW17, RSW16]. Regularization [Jah23, LNW21]. Regularized [HSH^{+}22]. Related [BBB^{+}15, RPMC21]. Relax [GIMS19]. Relaxation [SLH19]. Reliability [SS15]. Remediation [OV14]. Remote [BHTG21]. Rényi [ACD15]. Repetitions [Maz21]. Replica [Ari19]. Replicate [DPP18]. Repolarization [ABJR22]. Representations [ITG19, SAY22]. Representing [LHO^{+}21, MOM18]. Reproducing [DZ21, GR17]. Resin [PT17]. Resource [YWT20]. Resource-Constrained [YWT20]. Response [BPW15, FSK22, Gri14, Gri16, Gri17, HL17]. Results [BDKR22, Lás16]. Retrievals [PKH22]. Reversal [RML18]. Rice [RA21]. Ridge [SYP19, TWW20]. Rigorous [CMT18]. Rising [DS17]. Risk [APSG17, ACD15, EPS22, GHA19, HKTW18, HKT20, KS18, KS22, KJH22, ZKA22]. Risk-Adapted [KJH22]. Risk-Averse [APSG17, KS18, KS22, ZKA22]. Risk-Sensitive [ACD15]. Robust [ACD15, AKB^{+}23, BD23, HK18, SPHJ21, VV19]. Robustness [HG19, SRS17]. ROMES [DC15]. Root [KM15]. Rotated [YLT18]. Rotating [DMS23]. Rotational [HLY22]. Rule [DLS22, JM13]. Rules [vdBSBvB20].

Saddle [NP17]. SAGA [MN21]. Sample [DDS19, DLZ21, HGGT20, Min22, SA18, vdBSBvB20]. Sample-Adaptive [DDS19]. Sampler [ABPS14, MJP13, UBD^{+}22]. Samplers [BSV18]. Sampling [AP23, ASG13, BSS^{+}15, BS21, CG22, CCR19, CKR20, CGRF18, DG13, FG21, FN16, Hes13, HSS^{+}18, KSS^{+}19, MCWG22, NHM^{+}16, NP19, Oli17, PG22, PB21, SA18, STW23, ST17, TS23, UPMS21, Vol15]. Sampling-based [CG22]. Sampling/Spectral [ASG13]. SAR [CG22]. Satellite [HBC^{+}17, SGH^{+}19]. Saturation [IT17, JZZ21]. Scalable [CA23, MRS^{+}20, WCG23]. Scalar [BGHK20, DP16, MRST16]. Scale [MSDO14, PP23, RKSA15, TGL23, WCG23, APSS21]. Scaled [GW18, GXW22, KGL22]. Scales [CCS17]. Scaling [LLS22]. Scattering [BCL16, BTG14, Döl20, FLL15]. Scenario [DPP18]. Scheme [AF19, BCMZ16, BC22]. Schemes [CLYM20, NP19]. Schrödinger [BC22, JS22]. Scores [AGZ23]. SDEs [BS13b]. Sea [CDS22]. Second [YY19]. Selecting [Héa20, MvOCP13]. Selection [BCMN18, HKZ18b, OR21, SLD19, Tan15, WSB16, YWT20]. Semiautomatic [DNP21]. Semigroups [NP19]. Sensing [ABJN18, BHTG21, HSS^{+}18, YLT18]. Sensitive [ACD15]. Sensitivity [BRPP18, BR22, BMP18, BBCM22, BET^{+}14, DM16, DKPP16, EPS22, FG21, FKL21, GKL18, HBC19, JLK17, LCI14, LLS20, Maz21, Mor18, NHM^{+}16, PRB21, QPO^{+}18, Rah16, SNS16, SLD19, SS15, Tan15, WR16, WSB16, YY19]. Sensor [SK22]. Sequences [BIM^{+}21]. Sequential [BG16, BBR^{+}19, BJL^{+}18, DM15, DNP21, FMS17, HL17, KBJ14, MGBL18, PB21, SN17, SCN21, ST17, Tan15, ZLR19]. Series [BCKW22, Ker23, WB14]. Set [AKB^{+}23, BD23, PSDF14, Soi17]. Set-Valued [BD23]. Sets [ABCG16, BPA20, HGS13, KPR22, vdBSBvB20]. Settings [BKM18]. Several [BCL16]. Shape [BTG14, GP18, Sca22, WB16]. Shapley [BBD20, BBDM21, BBCM22, EDCMD22, Owe14, OP17, PRB21, RB19, SNS16]. Sheet [BPA20]. Shocks [LLS20]. Short [HG17]. sided [KP20]. Signals [CGHM14]. Simple [CS23]. Simulated [CKP23]. Simulation [BLV17, BLS15, GHA19, HBC^{+}17, MWP15, PKW18, SGH^{+}19, SPHJ21, UPMS21, XLWZ21, XLL14]. Simulation-Based [HBC^{+}17]. Simulations [ITG19, PG13]. Simulators [BW16, NHM^{+}16, OC16, OGH22, OCMB17, ZLR19, ZS21]. Single [HSH^{+}22]. Situational [HP14]. Size [SA18]. Sliced [YLT18]. Sliced-Inverse-Regression [YLT18]. Smooth [XS17]. Smoother [FBG^{+}20]. Smoothing [HB18]. Sobol’ [HG19, MD19, Owe13, Owe14, KPR22]. Solute [Cha15]. Solution [BOS15, BGV16, DKLM15, MPL16, MRST16, RS15a, SN17]. Solutions [HSZ19, NT18, SW23, SG15, SRS17]. Solver [CCD16, LES19, SP16]. Solvers [CEGT15, PP23]. Solving [ASG13, CMT18, UP15b, ZKA22]. Source [BCL16, LW22]. Sources [HWM18, SC20]. Space [DKPP16, HP14, MJP13, NT18, PG13, PZ20, RErB^{+}22]. Space-Filling [PZ20]. Space-Time [PG13]. Spaces [BSDGH23, DZ21, FKL21, GR17, ZHYL17]. Sparse [ABJN18, BCWZ14, CLNR15, FMS17, GGS23, HK18, HLY22, HSS^{+}18, KS16, LZ22, LMS21, MS17, MMRT16, ZCK19]. Sparse-Grid [BCWZ14]. Sparsity [HSZ19]. Spatial [ANSS14, OR21, TST18, vWGBS15]. Spatially [CRR16, DPP18]. Spatio [DMS23, SWGE22]. Spatio-Temporal [DMS23, SWGE22]. SPDE [ABJR22, Kue15]. Special [EDCMD22]. Specifying [SO14]. Speckle [CG22]. Spectral [APV17, ASG13, BW23, GP14a, HSZ19, KP20, SE16]. Spline [Rah20, RJ22]. Splines [BW16]. Splitting [BC22, CDGR19]. Spotlight [CG22]. Square [KM15]. Squared [HS21]. Squares [CLNR15, DS19, HNP23, LCE18, NvdGW20, PP23, RKSA15, SNM17]. Stability [CRG^{+}15, DY21, Kue15, PA19, SL22, Ver17]. Stabilized [AAV18, GIMS19, TBR18]. Stable [CLYM20, RErB^{+}22]. State [BCMN18, BKM18, CDMN22, MJP13, PMQ17, RErB^{+}22]. State-Space [MJP13, RErB^{+}22]. State/Parameter [PMQ17]. Stationary [SZ22]. Statistical [ADSG22, BC21, BHR18, CPBP20, DC15, FPGD^{+}16, HBC19, JM13, KBW18, PS17, SAY22, VCNGP16]. Statistically [CMT18]. Statistics [AS22, Gri14]. Stenosis [KBS13]. Stepping [GMN22]. Stiff [AAV18]. Stochastic [APV17, AAV18, AS22, ASG13, AF19, BGM16, BCMZ16, BST16, BST18, BS13a, BIM^{+}21, BOS15, BDKR22, BH15, BPR21, BD23, BU21, CBE18, CQ14, CHYZ13, CPBP20, CCH^{+}20, CKR20, CFGS20, DEH^{+}18, DLZ21, DKLM15, DKPP16, DPP18, EMN16, EM16, EMM20, EEHM14, FHC^{+}18, FG21, GJWZ16, GHT16, GTP14, GW18, GXW22, GIMS19, HUW13, HHS22, HPR18, HLPR18, HG17, HL15, IT17, JKS^{+}21, JS22, JT13, JT15, JLK17, JZZ21, KP18, Kou14, KLS15, KP20, KS16, LS18, LCE18, MS17, MMRT16, Maz21, MS15, MSDO14, MOM18, Mot19, NHM^{+}16, PAS14, PT17, PB23, PSDF14, Pul13, PA19, RSW16, RKSA15, SK22, SP16, SE16, SL22, TJWG15, TST18, UP15b, VCNGP16, WBSC^{+}15, WR16, XLWZ21, XLL14, ZHY21, ZNX14, ZS21]. Stokes [GIMS19, KBJ14, LES19, MUL19]. Stop [HSS^{+}18]. Stop-Sampling [HSS^{+}18]. Stopped [GB18]. Stopping [BLS15, BHR18]. Stopping-Convergence [BLS15]. Strategies [DEH^{+}18, PKNW19]. Strategy [PMM16]. Stratification [DTVW20]. Strong [BC22, Kre23]. Strongly [AS22]. Structure [CLW17, LK22]. Structured [WB14]. Study [BS18, BGR22, CCH^{+}20, KBJ14, OVHW19]. Sub [vAH17]. Sub- [vAH17]. Subnetworks [OHK^{+}20]. Subproblem [BGV16]. Subsampled [SNM17, TI14]. Subset [BLV17]. Subspace [BGA^{+}17, Ede23, GCB18, SE16]. Subspaces [CLS^{+}19, Lee19]. Subsurface [DKST15, DKST19]. Superfloe [CDS22]. Supervised [PO21]. Support [ZGS22]. Surface [BPW15, DEKR18]. Surfaces [BKM18, HL17]. Surrogate [Ade19, AP23, BRGT17, BDW13, DMS23, DFS20, Gri17, Ker23, MS17, OCMB17]. Surrogate-Based [OCMB17]. Surrogates [CSD17, RHH^{+}15, Sca22]. Survey [LMS21]. Symmetric [LS18]. System [Gri17, SPHJ21]. Systems [APSG17, BJP16, BD23, BHTG21, CRR16, DG19, DGVE19, DMS23, DPP18, FG21, FH16, GMC22, LCE18, MS17, MG21, MCTI16, MS15, MRS^{+}20, Pul13, PA19, RA21, RW21, SAS15, SP16, SL22, SRS17, Wal17, WBSC^{+}15, vAH17].

Tailed [Hos17]. Tails [HN17]. Taylor [CG21]. Technique [MMRT16, MRS^{+}20]. Techniques [WSB16]. Teleporting [LWZ22]. Temporal [DMS23, SWGE22]. Tensor [BG15, BRPP18, DKLM15, EMM20, HNP23, KKNT15, KS16, RDGS22]. Tensor-Train [RDGS22]. Term [HG17]. Terms [CPBP20, Soi17]. Test [KPR22]. Their [BGA^{+}17, KBW18]. Theorems [GK20]. Theoretic [BET^{+}14, SG15].

Theoretical [GXW22, TW16]. Theory [Gri16, Gri17, HSH^{+}22, HLR15, SA18, SNS16, TS23]. Theory-based [TS23]. Thin [BW16]. Thin-Plate [BW16]. Time [BCKW22, DG19, GMN22, GIMS19, KP18, Ker23, MS17, OB16, PB23, PG13, SAS15, SS15, WB14]. Time-Dependent [GIMS19, PB23]. Time-Harmonic [KP18]. Time-Series [Ker23]. Time-Stepping [GMN22]. Times [DPP18, GB18, HMR^{+}13, IDL17]. Tolerant [ABJN18]. Tomographic [UBD^{+}22]. Tomography [ADH23, BSS^{+}15, GBD23, HL15]. Topology [AKB^{+}23]. Trade [Maz21]. Trade-Off [Maz21]. Tradeoff [ST15]. Train [DKLM15, EMM20, RDGS22]. Trajectory [Ari19, HSH^{+}22]. Transfer [LK22, PT17]. Transform [CRR16, EB21]. Transformation [WR16]. Transformed [AGZ23, UP15b]. Transient [PAS14]. Transmission [AMV^{+}17]. Transport [CCR19, CCH^{+}20, CKR20, CHKW22, IT17, MDK20, NP19, PMM16, PM18, RML18]. Treatment [CPBP20]. Tree [HNP23]. Trends [PAS14]. Truncation [RPMC21]. Truncation-Related [RPMC21]. Trust [HS21]. Tsunami [BG16, LG17]. TT-Cross [DS19]. Two [BPR21, CCH^{+}20, Hos19, IT17, KP20, Maz21]. Two-Dimensional [CCH^{+}20]. Two-Level [BPR21]. Two-Phase [IT17]. Two-sided [KP20]. Type [BGR22, JM13, TWW20].

Ultrahigh [HKZ18a, HKZ18b]. Unbiased [CFJ^{+}21, HJLT23, SU20, SU21]. Uncertain [EMN16, GP14b, Kou14, LES19, MN21, PM22, SHL14, SRS17, UBD^{+}22, Ver17, WBSC^{+}15]. Uncertainties [AKB^{+}23, ABCG16, BDW13, LLSS17, OVHW19, RV20, RPMC21, RHBH20]. Uncertainty [Ade19, ADH23, APSS21, AJHSZ20, BSS^{+}15, BH15, BLSZ18, BK21, BFKRB22, BRB20, BS21, BU21, BHTG21, BPW15, CFN^{+}23, Cha15, CBE18, CLW17, CQR17, CDS22, CG22, CFGS20, DP16, DLS22, DFS20, DKST15, DKST19, DKPP16, EEHM14, EPS22, FPGD^{+}16, GK20, GK19, GMN22, GCB18, GKK^{+}21, HS21, HK18, HG19, HGS13, HSZ19, HBC^{+}17, HP14, HPW21, HLY22, KP18, KWT^{+}20, KW16, KMW14, LLS22, MMRT16, MPL16, MCTI16, MRS^{+}20, OHK^{+}20, OCMB17, PO21, PKH22, PP23, RJ22, RS15b, SWGE22, Sca22, SC20, SL22, SW23, SS13, TGL23, XLWZ21, YLT18, YWT20, ZCK19]. Unconstrained [Jah23]. Understanding [GMC22]. Unfolding [Lás16]. Uniform [LW17]. Uniqueness [HLPR18]. Universal [BRGT17]. Unknown [BBD20]. Unscented [EB21]. Unstable [BGA^{+}17, GCB18, MS15]. Unsteady [BOS15, PMQ17, RPMC21, SC20]. Updating [EFF^{+}23]. Using [APSG17, AGZ23, AKB^{+}23, BDKR22, BCKW22, BCMN18, BW16, BDW13, CRG^{+}15, CA23, CFGS20, DM16, DFS20, DLZ21, DNP21, FHC^{+}18, GSM22, HUW13, HGS13, HPW21, LLS22, Lee19, LvLP21, MS17, MG21, OGH22, PMM16, RHH^{+}15, RM18, RKSA15, TGG13, TGH17, VV19, VW20, WB16, ZCK19, ZGS22, ZS21, HNP23, MN21].

Validation [BD23, HSS^{+}18, MCWG22]. Value [Gri16, Gri17, HKTW18, HKT20, HGGT20, Owe14, OP17, SO14]. Value-at-Risk [HKT20]. Valued [BD23]. Var [SSJ^{+}13]. Variable [OGH22, SLD19, dRLI19]. Variables [GSM22, KPR22, RKSA15]. Variance [APSG17, BIM^{+}21, BBD20, DTVW20, GMC22, GM16, HPW21, Mor18, Owe13, QPO^{+}18, VCNGP16]. Variance-Based [Mor18]. Variate [PG22]. Variational [AKSAS22, BRB20, LZ22, MSDO14, RS15a]. Variations [MGBL18]. Varying [DG19, WGD22]. Vecchia [KGL22]. Vector [Soi15, Soi17]. Vectors [MS20, Soi17]. Verification [WSB16]. versus [Sta15]. via [ADSG22, ACD15, BRB20, HKTW18, HKZ18a, HKZ18b, KSW14, NP19, SYP19, SO14, Tuo19, dRLI19]. View [UBD^{+}22]. Viewpoint [HHS22]. Viscoelastic [KBS13]. Viscosity [LES19, MUL19]. Volume [AEST15]. Vries [XLL14].

Walkers [LWZ22]. Warped [FSK22, MGBL18]. Warping [MS17]. Wasserstein [DY21, EPS22, FKL21, Min22, WCL22]. Wave [FLL15, GKS21, GMN22, LW22, MMRT16]. Waveform [DY21]. Wavenumber [SW23]. Wavenumber-Explicit [SW23]. Way [BBB^{+}15]. Weak [AAV18, KLS15]. Weighted [BS21, CQ14, TBR18]. Weights [vdBSBvB20]. Well [HN17, Hos17, Lat20, PS20]. Well-Posed [HN17, Hos17]. Well-Posedness [PS20, Lat20]. White [BC22, CGRF18, Jah23, LNW21]. Whittle [KSS^{+}19].

X [ADH23]. X-Ray [ADH23].

Zakai [BCWZ14]. Zero [SLH19, SWTA23].

References

Abdulle:2018:OES

[AAV18] Assyr Abdulle, Ibrahim Almuslimani, and Gilles Vilmart. Optimal explicit stabilized integrator of weak order 1 for stiff and ergodic stochastic differential equations. SIAM/ASA Journal on Uncertainty Quantification, 6(2):937–964, ???? 2018. CODEN SJUQA3. ISSN 2166-

Azzimonti:2016:QUE

[ABCG16] Dario Azzimonti, Julien Bect, Clément Chevalier, and David Ginsbourger. Quantifying uncertainties on excursion sets under a Gaussian random field prior. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):850–874, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Adcock:2018:CSS

[ABJN18] Ben Adcock, Anyi Bao, John D. Jakeman, and Akil Narayan. Compressed sensing with sparse corruptions: Fault-tolerant sparse collocation approximations. SIAM/ASA Journal on Uncertainty Quantification, 6 (4):1424–1453, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Altmeyer:2022:PES

[ABJR22] Randolf Altmeyer, Till Bretschneider, Josef Janák, and Markus Reiß. Parameter estimation in an SPDE model for cell repolarization. SIAM/ASA Journal on Uncertainty Quantification, 10(1):179–199, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1373347.

Agapiou:2014:AGS

[ABPS14] Sergios Agapiou, Johnathan M. Bardsley, Omiros Papaspiliopoulos, and Andrew M. Stuart. Analysis of the Gibbs sampler for hierarchical inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 2 (1):511–544, ???? 2014. CO-DEN SJUQA3. ISSN 2166-2525.

Atar:2015:RBR

[ACD15] Rami Atar, Kenny Chowdhary, and Paul Dupuis. Robust bounds on risk-sensitive functionals via Rényi divergence. SIAM/ASA Journal on Uncertainty Quantification, 3(1):18– 33, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Abdulle:2020:BNH

[AD20] Assyr Abdulle and Andrea Di Blasio. A Bayesian numerical homogenization method for elliptic multiscale inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 8 (1):414–450, ???? 2020. CO-DEN SJUQA3. ISSN 2166-2525.

Adelmann:2019:NUQ

[Ade19] Andreas Adelmann. On nonintrusive uncertainty quantification and surrogate model construction in particle accelerator modeling. SIAM/ASA Journal on Uncertainty Quantification, 7(2):383–416, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Afkham:2023:UQI

[ADH23] Babak Maboudi Afkham, Yiqiu Dong, and Per Christian Hansen. Uncertainty quantification of inclusion boundaries in the context of Xray tomography. SIAM/ASA Journal on Uncertainty Quantification, 11(1):31–61, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1433782.

Akyildiz:2022:SFE

[ADSG22] Ömer Deniz Akyildiz, Connor Duffin, Sotirios Sabanis, and Mark Girolami. Statistical finite elements via Langevin dynamics. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1560–1585, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1463094.

Arbogast:2015:PEE

[AEST15] T. Arbogast, D. Estep, B. Sheehan, and S. Tavener. A posteriori error estimates for mixed finite element and finite volume methods for parabolic problems coupled through a boundary. SIAM/ASA Journal on Uncertainty Quantification, 3(1):169– 198, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Ayi:2019:AAP

[AF19] Nathalie Ayi and Erwan Faou. Analysis of an asymptotic preserving scheme for stochastic linear kinetic equations in the diffusion limit. SIAM/ASA Journal on Uncertainty Quantification, 7(2):760–785, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Allen:2023:EFH

[AGZ23] Sam Allen, David Ginsbourger, and Johanna Ziegel. Evaluating forecasts for high-impact events using transformed kernel scores. SIAM/ASA Journal on Uncertainty Quantification, 11(3):906–940, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M1532184.

Aylwin:2020:DUQ

[AJHSZ20] Ruben Aylwin, Carlos Jerez-Hanckes, Christoph Schwab, and Jakob Zech. Domain uncertainty quantification in computational electromagnetics. SIAM/ASA Journal on Uncertainty Quantification, 8(1): 301–341, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Audouze:2023:RLS

[AKB^{+}23] Christophe Audouze, Aaron Klein, Adrian Butscher, Nigel Morris, Prasanth Nair, and Masayuki Yano. Robust levelset-based topology optimization under uncertainties using anchored ANOVA Petrov– Galerkin method. SIAM/ ASA Journal on Uncertainty Quantification, 11(3):877–905, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M1524722.

Agrawal:2022:VIA

[AKSAS22] Shiv Agrawal, Hwanwoo Kim, Danie Sanz-Alonso, and Alexander Strang. A variational inference approach to inverse problems with gamma hyperpriors. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1533–1559, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M146209X.

Andrianakis:2017:EHM

[AMV^{+}17] Ioannis Andrianakis, Nicky Mc-Creesh, Ian Vernon, Trevelyan J. McKinley, Jeremy E. Oakley, Rebecca N. Nsubuga, Michael Goldstein, and Richard G. White. Efficient history matching of a high dimensional individual-based HIV transmission model. SIAM/ASA Journal on Uncertainty Quantification, 5(1):694–719, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Azzimonti:2014:MFE

[ANSS14] Laura Azzimonti, Fabio Nobile, Laura M. Sangalli, and Piercesare Secchi. Mixed finite elements for spatial regression with PDE penalization. SIAM/ASA Journal on Uncertainty Quantification, 2(1):305– 335, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Alsup:2023:CAS

[AP23] Terrence Alsup and Benjamin Peherstorfer. Contextaware surrogate modeling for balancing approximation and sampling costs in multifidelity importance sampling and Bayesian inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 11(1):285–319, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1445594.

Alexanderian:2017:MVR

[APSG17] Alen Alexanderian, Noemi Petra, Georg Stadler, and Omar Ghattas. Mean-variance riskaverse optimal control of systems governed by PDEs with random parameter fields using quadratic approximations. SIAM/ASA Journal on Uncertainty Quantification, 5(1): 1166–1192, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Alexanderian:2021:ODL

[APSS21] Alen Alexanderian, Noemi Petra, Georg Stadler, and Isaac Sunseri. Optimal design of large-scale Bayesian linear inverse problems under reducible model uncertainty: Good to know what you don’t know. SIAM/ASA Journal on Uncertainty Quantification, 9(1):163– 184, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Abdulle:2017:SMM

[APV17] A. Abdulle, G. A. Pavliotis, and U. Vaes. Spectral methods for multiscale stochastic differential equations. SIAM/ASA Journal on Uncertainty Quantification, 5(1):720–761, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Aristoff:2019:GPR

[Ari19]  David Aristoff.  Generaliz-

ing parallel replica dynamics: Trajectory fragments, asynchronous computing, and PDMPs. SIAM/ASA Journal on Uncertainty Quantification, 7(2):685–719, ???? 2019. CO-DEN SJUQA3. ISSN 2166-2525.

Arbabi:2022:GSM

[AS22] Hassan Arbabi and Themistoklis Sapsis. Generative stochastic modeling of strongly nonlinear flows with non-Gaussian statistics. SIAM/ASA Journal on Uncertainty Quantification, 10(2):555–583, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1359833.

Arnst:2013:HSS

[ASG13] M. Arnst, C. Soize, and R. Ghanem. Hybrid sampling/spectral method for solving stochastic coupled problems. SIAM/ASA Journal on Uncertainty Quantification, 1 (1):218–243, ???? 2013. CO-DEN SJUQA3. ISSN 2166-2525.

Ali:2017:MMC

[AUH17] Ahmad Ahmad Ali, Elisabeth Ullmann, and Michael Hinze. Multilevel Monte Carlo analysis for optimal control of elliptic PDEs with random coefficients. SIAM/ASA Journal on Uncertainty Quantification, 5(1):466–492, ???? 2017. CO-DEN SJUQA3. ISSN 2166-2525.

Berliner:2016:FMM

[BB16] L. Mark Berliner and Jenný Brynjarsdóttir. A framework for multi-model ensembling. SIAM/ASA Journal on Uncertainty Quantification, 4(1): 902–923, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Banerjee:2015:MDE

[BBB^{+}15] Sourabh Banerjee, Ayanendranath Basu, Sourabh Bhattacharya, Smarajit Bose, Dalia Chakrabarty, and Soumendu Sundar Mukherjee. Minimum distance estimation of milky way model parameters and related inference. SIAM/ASA Journal on Uncertainty Quantification, 3(1):91–115, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Broto:2022:BDC

[BBCM22] Baptiste Broto, François Bachoc, Laura Clouvel, and Jean-Marc Martinez. Block-diagonal covariance estimation and application to the Shapley effects in sensitivity analysis. SIAM/ ASA Journal on Uncertainty Quantification, 10(1):379–403, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1358839.

Broto:2020:VRE

[BBD20] Baptiste Broto, François Bachoc, and Marine Depecker.

Variance reduction for estimation of Shapley effects and adaptation to unknown input distribution. SIAM/ASA Journal on Uncertainty Quantification, 8(2):693–716, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Broto:2021:GLA

[BBDM21] Baptiste Broto, François Bachoc, Marine Depecker, and Jean-Marc Martinez. Gaussian linear approximation for the estimation of the Shapley effects. SIAM/ASA Journal on Uncertainty Quantification, 9 (3):1132–1151, ???? 2021. CO-DEN SJUQA3. ISSN 2166-2525.

BenSalem:2019:GPB

[BBR^{+}19] Malek Ben Salem, François Bachoc, Olivier Roustant, Fabrice Gamboa, and Lionel Tomaso. Gaussian process-based dimension reduction for goal-oriented sequential design. SIAM/ASA Journal on Uncertainty Quantification, 7(4):1369–1397, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Bardsley:2021:OBM

[BC21] Johnathan M. Bardsley and Tiangang Cui. Optimizationbased Markov chain Monte Carlo methods for nonlinear hierarchical statistical inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 9(1):29–64, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Brehier:2022:SRC

[BC22] Charles-Edouard Bréhier and David Cohen. Strong rates of convergence of a splitting scheme for Schrödinger equations with nonlocal interaction cubic nonlinearity and white noise dispersion. SIAM/ ASA Journal on Uncertainty Quantification, 10(1):453–480, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1378168.

Binev:2017:DAR

[BCD^{+}17] Peter Binev, Albert Cohen, Wolfgang Dahmen, Ronald De-Vore, Guergana Petrova, and Przemyslaw Wojtaszczyk. Data assimilation in reduced modeling. SIAM/ASA Journal on Uncertainty Quantification, 5 (1):1–29, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Banks:2015:APP

[BCH15] H. T. Banks, Jared Catenacci, and Shuhua Hu. Asymptotic properties of probability measure estimators in a nonparametric model. SIAM/ASA Journal on Uncertainty Quantification, 3(1):417–433, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Bhatnagar:2022:CMC

[BCKW22] Saumya Bhatnagar, Won Chang, Seonjin Kim, and Jiali Wang. Computer model calibration with time series data using deep learning and quantile regression. SIAM/ASA Journal on Uncertainty Quantification, 10(1):1–26, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1382581.

Bao:2016:IRS

[BCL16] Gang Bao, Chuchu Chen, and Peijun Li. Inverse random source scattering problems in several dimensions. SIAM/ASA Journal on Uncertainty Quantification, 4(1):1263–1287, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Binev:2018:GAO

[BCMN18] Peter Binev, Albert Cohen, Olga Mula, and James Nichols. Greedy algorithms for optimal measurements selection in state estimation using reduced models. SIAM/ASA Journal on Uncertainty Quantification, 6(3): 1101–1126, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Bao:2016:FOS

[BCMZ16] Feng Bao, Yanzhao Cao, Amnon Meir, and Weidong Zhao. A first order scheme for backward doubly stochastic differential equations. SIAM/ASA Journal on Uncertainty Quantification, 4(1):413–445, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Bao:2014:HSG

[BCWZ14] Feng Bao, Yanzhao Cao, Clayton Webster, and Guannan Zhang. A hybrid sparse-grid approach for nonlinear filtering problems based on adaptivedomain of the Zakai equation approximations. SIAM/ASA Journal on Uncertainty Quantification, 2(1):784–804, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Bishop:2023:RKB

[BD23] Adrian N. Bishop and Pierre Del Moral. Robust Kalman and Bayesian set-valued filtering and model validation for linear stochastic systems. SIAM/ ASA Journal on Uncertainty Quantification, 11(2):389–425, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M1481270.

Bergou:2022:SLM

[BDKR22] El Houcine Bergou, Youssef Diouane, Vyacheslav Kungurtsev, and Clément W. Royer. A stochastic Levenberg– Marquardt method using random models with complexity results. SIAM/ASA Journal on Uncertainty Quantification, 10(1):507–536, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1366253.

Butler:2013:PUU

[BDW13] T. Butler, C. Dawson, and T. Wildey. Propagation of uncertainties using improved surrogate models. SIAM/ASA Journal on Uncertainty Quantification, 1(1):164–191, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Berger:2013:ME

[BEG13] James Berger, Donald Estep, and Max Gunzburger. Message from the Editors. SIAM/ASA Journal on Uncertainty Quantification, 1(1):1, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Butler:2014:MTC

[BET^{+}14] T. Butler, D. Estep, S. Tavener, C. Dawson, and J. J. Westerink. A measure-theoretic computational method for inverse sensitivity problems III: Multiple quantities of interest. SIAM/ASA Journal on Uncertainty Quantification, 2(1):174– 202, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Birmpa:2022:MUC

[BFKRB22] Panagiota Birmpa, Jinchao Feng, Markos A. Katsoulakis, and Luc Rey-Bellet. Model uncertainty and correctability for directed graphical models. SIAM/ASA Journal on Uncertainty Quantification, 10 (4):1461–1512, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1434453.

Ballani:2015:HTA

[BG15] Jonas Ballani and Lars Grasedyck. Hierarchical tensor approximation of output quantities of parameter-dependent PDEs. SIAM/ASA Journal on Uncertainty Quantification, 3(1): 852–872, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Beck:2016:SDM

[BG16] Joakim Beck and Serge Guillas.

Sequential design with Mutual Information for Computer Experiments (MICE): Emulation of a tsunami model. SIAM/ ASA Journal on Uncertainty Quantification, 4(1):739–766, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Bocquet:2017:DKF

[BGA^{+}17] Marc Bocquet, Karthik S. Gurumoorthy, Amit Apte, Alberto Carrassi, Colin Grudzien, and Christopher K. R. T. Jones. Degenerate Kalman filter error covariances and their convergence onto the unstable subspace. SIAM/ASA Journal on Uncertainty Quantification, 5 (1):304–333, ???? 2017. CO-DEN SJUQA3. ISSN 2166-2525.

Borggaard:2020:BAE

[BGHK20] Jeff Borggaard, Nathan Glatt-Holtz, and Justin Krometis. A Bayesian approach to estimating background flows from a passive scalar. SIAM/ASA Journal on Uncertainty Quantification, 8(3):1036–1060, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Bachouch:2016:ERM

[BGM16] Achref Bachouch, Emmanuel Gobet, and Anis Matoussi. Empirical regression method for backward doubly stochastic differential equations. SIAM/ASA Journal on Uncertainty Quantification, 4(1):358–379, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Bourgey:2022:CSP

[BGR22] Florian Bourgey, Emmanuel Gobet, and Clément Rey. A comparative study of polynomialtype chaos expansions for indicator functions. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1350–1383, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1413146.

Bergou:2016:LMM

[BGV16] E. Bergou, S. Gratton, and L. N. Vicente. Levenberg– Marquardt methods based on probabilistic gradient models and inexact subproblem solution, with application to data assimilation. SIAM/ASA Journal on Uncertainty Quantification, 4(1):924–951, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Berry:2015:NUQ

[BH15] Tyrus Berry and John Harlim. Nonparametric uncertainty quantification for stochastic gradient flows. SIAM/ASA Journal on Uncertainty Quantification, 3(1):484–508, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Blanchard:2018:OAE

[BHR18] Gilles Blanchard, Marc Hoffmann, and Markus Reiß. Optimal adaptation for early stopping in statistical inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 6 (3):1043–1075, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Braverman:2021:PHU

[BHTG21] Amy Braverman, Jonathan Hobbs, Joaquim Teixeira, and Michael Gunson. Post hoc uncertainty quantification for remote sensing observing systems. SIAM/ASA Journal on Uncertainty Quantification, 9 (3):1064–1093, ???? 2021. CO-DEN SJUQA3. ISSN 2166-2525.

Belomestny:2021:VRD

[BIM^{+}21] Denis Belomestny, Leonid Iosipoi, Eric Moulines, Alexey Naumov, and Sergey Samsonov. Variance reduction for dependent sequences with applications to stochastic gradient MCMC. SIAM/ASA Journal on Uncertainty Quantification, 9(2):507–535, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Beskos:2018:MSM

[BJL^{+}18] Alexandros Beskos, Ajay Jasra, Kody Law, Youssef Marzouk, and Yan Zhou. Multilevel sequential Monte Carlo with dimension-independent likelihood-informed proposals. SIAM/ASA Journal on Uncertainty Quantification, 6(2): 762–786, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Berlyand:2016:CRM

[BJP16] Leonid Berlyand, Pierre-Emmanuel Jabin, and Mykhailo Potomkin. Complexity reduction in many particle systems with random initial data. SIAM/ASA Journal on Uncertainty Quantification, 4(1):446–474, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Birmpa:2021:UQM

[BK21] Panagiota Birmpa and Markos A. Katsoulakis. Uncertainty quantification for Markov random fields. SIAM/ASA Journal on Uncertainty Quantification, 9 (4):1457–1498, ???? 2021. CO-DEN SJUQA3. ISSN 2166-2525.

Bousquet:2018:ALS

[BKM18] N. Bousquet, T. Klein, and V. Moutoussamy. Approximation of limit state surfaces in monotonic Monte Carlo settings, with applications to classification. SIAM/ASA Journal on Uncertainty Quantification, 6(1):1–33, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Abdellah:2021:DER

[BLOP21] Amal Ben Abdellah, Pierre L’Ecuyer, Art B. Owen, and Florian Puchhammer. Density estimation by randomized quasi-Monte Carlo. SIAM/ ASA Journal on Uncertainty Quantification, 9(1):280–301, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Belomestny:2015:MSB

[BLS15] Denis Belomestny, Marcel Ladkau, and John Schoenmakers. Multilevel simulation based policy iteration for optimal stopping-convergence and complexity. SIAM/ASA Journal on Uncertainty Quantification, 3(1):460–483, ???? 2015. CO-DEN SJUQA3. ISSN 2166-2525.

Bertozzi:2018:UQG

[BLSZ18] Andrea L. Bertozzi, Xiyang Luo, Andrew M. Stuart, and Konstantinos C. Zygalakis. Uncertainty quantification in graph-based classification of high dimensional data. SIAM/ ASA Journal on Uncertainty Quantification, 6(2):568–595, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Bect:2017:BSS

[BLV17] Julien Bect, Ling Li, and Emmanuel Vazquez. Bayesian subset simulation. SIAM/ASA Journal on Uncertainty Quantification, 5(1):762–786, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Borgonovo:2018:FAM

[BMP18] Emanuele Borgonovo, Max D.

Morris, and Elmar Plischke.

Functional ANOVA with multiple distributions: Implications for the sensitivity analysis of computer experiments. SIAM/ASA Journal on Uncertainty Quantification, 6(1):397– 427, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Bonizzoni:2014:PAD

[BN14] Francesca Bonizzoni and Fabio Nobile. Perturbation analysis for the Darcy problem with log-normal permeability. SIAM/ASA Journal on Uncertainty Quantification, 2(1):223– 244, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Benner:2015:LRS

[BOS15] Peter Benner, Akwum Onwunta, and Martin Stoll. Lowrank solution of unsteady diffusion equations with stochastic coefficients. SIAM/ASA Journal on Uncertainty Quantification, 3(1):622–649, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Bulthuis:2020:MQB

[BPA20] Kevin Bulthuis, Frank Pattyn, and Maarten Arnst. A multifidelity quantile-based approach for confidence sets of random excursion sets with application to ice-sheet dynamics. SIAM/ASA Journal on Uncertainty Quantification, 8(3): 860–890, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Bespalov:2021:TLP

[BPR21] Alex Bespalov, Dirk Praetorius, and Michele Ruggeri. Two-level a posteriori error estimation for adaptive multilevel stochastic Galerkin finite element method. SIAM/ASA Journal on Uncertainty Quantification, 9(3): 1184–1216, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Bryant:2015:EDA

[BPW15] C. M. Bryant, S. Prudhomme, and T. Wildey. Error decomposition and adaptivity for response surface approximations from PDEs with parametric uncertainty. SIAM/ASA Journal on Uncertainty Quantification, 3(1):1020–1045, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Benner:2018:LRE

[BQS18] Peter Benner, Yue Qiu, and Martin Stoll. Low-rank eigenvector compression of posterior covariance matrices for linear Gaussian inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 6(2):965– 989, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Bespalov:2018:EAA

[BR18] Alex Bespalov and Leonardo Rocchi. Efficient adaptive algorithms for elliptic PDEs with random data. SIAM/ASA Journal on Uncertainty Quantification, 6(1):243–272, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Barr:2022:GKM

[BR22] John Barr and Herschel Rabitz. A generalized kernel method for global sensitivity analysis. SIAM/ASA Journal on Uncertainty Quantification, 10(1):27–54, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1354829.

Birrell:2020:UQM

[BRB20] Jeremiah Birrell and Luc Rey-Bellet. Uncertainty quantification for Markov processes via variational principles and functional inequalities. SIAM/ASA Journal on Uncertainty Quantification, 8(2):539–572, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

BenSalem:2017:UPD

[BRGT17] Malek Ben Salem, Olivier Roustant, Fabrice Gamboa, and Lionel Tomaso. Universal prediction distribution for surrogate models. SIAM/ASA Journal on Uncertainty Quantification, 5(1):1086–1109, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Ballester-Ripoll:2018:TAA

[BRPP18] Rafael Ballester-Ripoll, Enrique G. Paredes, and Renato Pajarola. Tensor algorithms for advanced sensitivity metrics. SIAM/ASA Journal on Uncertainty Quantification, 6 (3):1172–1197, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Batou:2013:CLM

[BS13a] A. Batou and C. Soize. Calculation of Lagrange multipliers in the construction of maximum entropy distributions in high stochastic dimension. SIAM/ASA Journal on Uncertainty Quantification, 1(1):431– 451, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Bender:2013:PEB

[BS13b] Christian Bender and Jessica Steiner. A posteriori estimates for backward SDEs. SIAM/ASA Journal on Uncertainty Quantification, 1(1):139– 163, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Barth:2018:SEP

[BS18] Andrea Barth and Andreas Stein. A study of elliptic partial differential equations with jump diffusion coefficients. SIAM/ASA Journal on Uncertainty Quantification, 6(4): 1707–1743, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Blanchard:2021:OWO

[BS21] Antoine Blanchard and Themistoklis Sapsis. Output-weighted optimal sampling for Bayesian experimental design and uncertainty quantification. SIAM/ ASA Journal on Uncertainty Quantification, 9(2):564–592, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Blanchet-Scalliet:2023:GPR

[BSDGH23] Christophette Blanchet-Scalliet, Bruno Demory, Thierry Gonon, and Céline Helbert. Gaussian process regression on nested spaces. SIAM/ASA Journal on Uncertainty Quantification, 11(2):426–451, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1445053.

Bardsley:2015:RTO

[BSS^{+}15] Johnathan M. Bardsley, Aku Seppänen, Antti Solonen, Heikki Haario, and Jari Kaipio. Randomize-then-optimize for sampling and uncertainty quantification in electrical impedance tomography. SIAM/ ASA Journal on Uncertainty Quantification, 3(1):1136–1158, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Barajas-Solano:2016:SCM

[BST16] David A. Barajas-Solano and Daniel M. Tartakovsky. Stochastic collocation methods for nonlinear parabolic equations with random coefficients. SIAM/ ASA Journal on Uncertainty Quantification, 4(1):475–494, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Barajas-Solano:2018:PCD

[BST18] David A. Barajas-Solano and Alexandre M. Tartakovsky. Probability and cumulative density function methods for the stochastic advection–reaction equation. SIAM/ASA Journal on Uncertainty Quantification, 6(1):180–212, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Brown:2018:LRI

[BSV18] D. Andrew Brown, Arvind Saibaba, and Sarah Vallélian. Low-rank independence samplers in hierarchical Bayesian inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 6(3):1076–1100, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Bui-Thanh:2014:AID

[BTG14] Tan Bui-Thanh and Omar Ghattas. An analysis of infinite dimensional Bayesian inverse shape acoustic scattering and its numerical approximation. SIAM/ASA Journal on Uncertainty Quantification, 2 (1):203–222, ???? 2014. CO-DEN SJUQA3. ISSN 2166-2525.

Branicki:2021:LUQ

[BU21] Michal Branicki and Kenneth Uda. Lagrangian uncertainty quantification and information inequalities for stochastic flows. SIAM/ASA Journal on Uncertainty Quantification, 9(3): 1242–1313, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Bowman:2016:EMS

[BW16] Veronica E. Bowman and David C. Woods. Emulation of multivariate simulators using thin-plate splines with application to atmospheric dispersion. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):1323–1344, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Bungert:2023:CDD

[BW23] Leon Bungert and Philipp Wacker. Complete deterministic dynamics and spectral decomposition of the linear ensemble Kalman inversion. SIAM/ASA Journal on Uncertainty Quantification, 11(1):320–357, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1429461.

Chen:2023:SPB

[CA23] Yian Chen and Mihai Anitescu. Scalable physics-based maximum likelihood estimation using hierarchical matrices. SIAM/ASA Journal on Uncertainty Quantification, 11(2):682–725, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1458880.

Chaudhry:2018:EDE

[CBE18] Jehanzeb H. Chaudhry, Nathanial Burch, and Donald Estep. Efficient distribution estimation and uncertainty quantification for elliptic problems on domains with stochastic boundaries. SIAM/ASA Journal on Uncertainty Quantification, 6 (3):1127–1150, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Capistran:2016:BAO

[CCD16] Marcos A. Capistrán, J. Andrés Christen, and Sophie Donnet. Bayesian analysis of ODEs: Solver optimal accuracy and Bayes factors. SIAM/ASA Journal on Uncertainty Quantification, 4(1):829–849, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Cotter:2020:PFS

[CCH^{+}20] Colin Cotter, Dan Crisan, Darryl D. Holm, Wei Pan, and Igor Shevchenko. A particle filter for stochastic advection by Lie transport: a case study for the damped and forced incompressible two-dimensional Euler equation. SIAM/ASA Journal on Uncertainty Quantification, 8(4):1446–1492, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Cotter:2019:ETA

[CCR19] Colin Cotter, Simon Cotter, and Paul Russell. Ensemble transport adaptive importance sampling. SIAM/ASA Journal on Uncertainty Quantification, 7(2):444–471, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Callahan:2017:BML

[CCS17] Margaret Callahan, Daniela Calvetti, and Erkki Somersalo. Beyond the model limit: Parameter inference across scales.

SIAM/ASA Journal on Uncertainty Quantification, 5(1):665– 693, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Chen:2022:AAP

[CCZW22] Jialei Chen, Zhehui Chen, Chuck Zhang, and C. F. Jeff Wu. APIK: Active physicsinformed kriging model with partial differential equations. SIAM/ASA Journal on Uncertainty Quantification, 10 (1):481–506, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1389285.

Cockayne:2021:PGF

[CD21] Jonathan Cockayne and Andrew Duncan. Probabilistic gradients for fast calibration of differential equation models. SIAM/ASA Journal on Uncertainty Quantification, 9(4): 1643–1672, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Cerou:2019:ANA

[CDGR19] Frédéric Cérou, Bernard Delyon, Arnaud Guyader, and Mathias Rousset. On the asymptotic normality of adaptive multilevel splitting. SIAM/ ASA Journal on Uncertainty Quantification, 7(1):1–30, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Conrad:2018:PLA

[CDM^{+}18] Patrick R. Conrad, Andrew D. Davis, Youssef M. Marzouk, Natesh S. Pillai, and Aaron Smith. Parallel local approximation MCMC for expensive models. SIAM/ASA Journal on Uncertainty Quantification, 6(1):339–373, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Cohen:2022:NRM

[CDMN22] Albert Cohen, Wolfgang Dahmen, Olga Mula, and James Nichols. Nonlinear reduced models for state and parameter estimation. SIAM/ASA Journal on Uncertainty Quantification, 10(1):227–267, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1380818.

Chen:2022:SPP

[CDS22] Nan Chen, Quanling Deng, and Samuel Stechmann. Superfloe parameterization with physics constraints for uncertainty quantification of sea ice floes. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1384–1409, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1428777.

Chaudhry:2015:PAI

[CEGT15] J. H. Chaudhry, D. Estep, V. Ginting, and S. Tavener. A posteriori analysis for iterative solvers for nonautonomous evolution problems. SIAM/ASA Journal on Uncertainty Quantification, 3(1):434–459, ????

2015. CODEN SJUQA3. ISSN 2166-2525.

Crepey:2020:UQS

[CFGS20] S. Crépey, G. Fort, E. Gobet, and U. Stazhynski. Uncertainty quantification for stochastic approximation limits using chaos expansion. SIAM/ASA Journal on Uncertainty Quantification, 8(3):1061–1089, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Chada:2021:UID

[CFJ^{+}21] Neil K. Chada, Jordan Franks, Ajay Jasra, Kody J. Law, and Matti Vihola. Unbiased inference for discretely observed hidden Markov model diffusions. SIAM/ASA Journal on Uncertainty Quantification, 9(2):763– 787, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Cartee:2023:QMU

[CFN^{+}23] Elliot Cartee, Antonio Farah, April Nellis, Jacob Van Hook, and Alexander Vladimirsky. Quantifying and managing uncertainty in piecewise-deterministic Markov processes. SIAM/ ASA Journal on Uncertainty Quantification, 11(3):814–847, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1357275.

Chen:2021:TAC

[CG21] Peng Chen and Omar Ghattas. Taylor approximation for chance constrained optimization problems governed by partial differential equations with high-dimensional random parameters. SIAM/ASA Journal on Uncertainty Quantification, 9(4):1381–1410, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Churchill:2022:SBS

[CG22] Victor Churchill and Anne Gelb. Sampling-based spotlight SAR image reconstruction from phase history data for speckle reduction and uncertainty quantification. SIAM/ASA Journal on Uncertainty Quantification, 10(3):1225–1249, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1379721.

Chen:2014:MAP

[CGHM14] Nan Chen, Dimitrios Giannakis, Radu Herbei, and Andrew J. Majda. An MCMC algorithm for parameter estimation in signals with hidden intermittent instability. SIAM/ ASA Journal on Uncertainty Quantification, 2(1):647–669, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Croci:2018:EWN

[CGRF18] M. Croci, M. B. Giles, M. E.

Rognes, and P. E. Farrell. Efficient white noise sampling and coupling for multilevel Monte Carlo with nonnested meshes. SIAM/ASA Journal on Uncertainty Quantification, 6(4): 1630–1655, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Charrier:2015:NAA

[Cha15] Julia Charrier. Numerical analysis of the advection-diffusion of a solute in porous media with uncertainty. SIAM/ASA Journal on Uncertainty Quantification, 3(1):650–685, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Chen:2020:ICC

[Che20] Nan Chen. An information criterion for choosing observation locations in data assimilation and prediction. SIAM/ASA Journal on Uncertainty Quantification, 8(4):1548–1573, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Clason:2019:GMB

[CHKP19] Christian Clason, Tapio Helin, Remo Kretschmann, and Petteri Piiroinen. Generalized modes in Bayesian inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 7 (2):652–684, ???? 2019. CO-DEN SJUQA3. ISSN 2166-2525.

Cox:2022:MCM

[CHKW22] Alexander M. G. Cox, Simon C. Harris, Andreas E. Kyprianou, and Minmin Wang. Monte Carlo methods for the neutron transport equation. SIAM/ ASA Journal on Uncertainty Quantification, 10(2):775–825, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1390578.

Cheng:2013:DDS

[CHYZ13] Mulin Cheng, Thomas Y.

Hou, Mike Yan, and Zhiwen Zhang. A data-driven stochastic method for elliptic PDEs with random coefficients. SIAM/ASA Journal on Uncertainty Quantification, 1(1): 452–493, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Chada:2022:MEK

[CJY22] Neil K. Chada, Ajay Jasra, and Fangyuan Yu. Multilevel ensemble Kalman–Bucy filters. SIAM/ASA Journal on Uncertainty Quantification, 10(2):584–618, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1423762.

Chak:2023:GLE

[CKP23] Martin Chak, Nikolas Kantas, and Grigorios A. Pavliotis. On the generalized Langevin equation for simulated annealing. SIAM/ ASA Journal on Uncertainty Quantification, 11(1):139–167, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1462970.

Cotter:2020:TMA

[CKR20] Simon L. Cotter, Ioannis G.

Kevrekidis, and Paul T. Russell. Transport map accelerated adaptive importance sampling, and application to inverse problems arising from multiscale stochastic reaction networks. SIAM/ASA Journal on Uncertainty Quantification, 8(4):1383–1413, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Chevreuil:2015:LSM

[CLNR15] M. Chevreuil, R. Lebrun, A. Nouy, and P. Rai. A leastsquares method for sparse low rank approximation of multivariate functions. SIAM/ASA Journal on Uncertainty Quantification, 3(1):897–921, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Coleman:2019:GFC

[CLS^{+}19] Kayla D. Coleman, Allison Lewis, Ralph C. Smith, Brian Williams, Max Morris, and Bassam Khuwaileh. Gradientfree construction of active subspaces for dimension reduction in complex models with applications to neutronics. SIAM/ ASA Journal on Uncertainty Quantification, 7(1):117–142, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Chen:2017:FCS

[CLW17] Hao Chen, Jason L. Loeppky, and William J. Welch. Flexible correlation structure for accurate prediction and uncertainty quantification in Bayesian Gaussian process emulation of a computer model. SIAM/ASA Journal on Uncertainty Quantification, 5(1):598–620, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Crisan:2020:SAS

[CLYM20] Dan Crisan, Alberto López-Yela, and Joaquin Miguez. Stable approximation schemes for optimal filters. SIAM/ASA Journal on Uncertainty Quantification, 8(1):483–509, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Chen:2018:RAE

[CMT18] Nan Chen, Andrew J. Majda, and Xin T. Tong. Rigorous analysis for efficient statistically accurate algorithms for solving Fokker–Planck equations in large dimensions. SIAM/ASA Journal on Uncertainty Quantification, 6(3): 1198–1223, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Constantinescu:2020:STI

[CPBP20] Emil M. Constantinescu, Noémi Petra, Julie Bessac, and Cosmin G. Petra. Statistical treatment of inverse problems constrained by differential equations-based models with stochastic terms. SIAM/ASA Journal on Uncertainty Quantification, 8(1):170–197, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Chen:2014:WRB

[CQ14] Peng Chen and Alfio Quarteroni. Weighted reduced basis method for stochastic optimal control problems with elliptic PDE constraint. SIAM/ASA Journal on Uncertainty Quantification, 2(1):364–396, ????

2014. CODEN SJUQA3. ISSN 2166-2525.

Chen:2017:RBM

[CQR17] Peng Chen, Alfio Quarteroni, and Gianluigi Rozza. Reduced basis methods for uncertainty quantification. SIAM/ASA Journal on Uncertainty Quantification, 5(1):813–869, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Chen:2015:OLE

[CRG^{+}15] Si Chen, Kristofer-Roy G. Reyes, Maneesh K. Gupta, Michael C. McAlpine, and Warren B. Powell. Optimal learning in experimental design using the knowledge gradient policy with application to characterizing nanoemulsion stability. SIAM/ASA Journal on Uncertainty Quantification, 3(1): 320–345, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Chustagulprom:2016:HET

[CRR16] Nawinda Chustagulprom, Sebastian Reich, and Maria Reinhardt. A hybrid ensemble transform particle filter for nonlinear and spatially extended dynamical systems. SIAM/ ASA Journal on Uncertainty Quantification, 4(1):592–608, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Chernov:2023:SBF

[CS23] Alexey Chernov and Erik Marc Schetzke. A simple, biasfree approximation of covariance functions by the multilevel Monte Carlo method having nearly optimal complexity. SIAM/ASA Journal on Uncertainty Quantification, 11(3):941–969, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M1506845.

Cesmelioglu:2017:PBK

[CSD17] A. Cesmelioglu, M. Song, and D. Drignei. Physics-based kriging surrogates for a class of finite element codes. SIAM/ASA Journal on Uncertainty Quantification, 5(1):870–889, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Damblin:2018:AND

[DBK^{+}18] Guillaume Damblin, Pierre Barbillon, Merlin Keller, Alberto Pasanisi, and Éric Parent. Adaptive numerical designs for the calibration of computer codes. SIAM/ASA Journal on Uncertainty Quantification, 6(1):151–179, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Drohmann:2015:RMS

[DC15] Martin Drohmann and Kevin Carlberg. The ROMES method for statistical modeling of reduced-order-model error. SIAM/ASA Journal on Uncertainty Quantification, 3 (1):116–145, ???? 2015. CO-DEN SJUQA3. ISSN 2166-

Das:2016:PAD

[DC16] Sonjoy Das and Sourish Chakravarty. Predictive algorithm for detection of microcracks from macroscale observables. SIAM/ ASA Journal on Uncertainty Quantification, 4(1):660–707, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Detommaso:2019:CLM

[DDS19] Gianluca Detommaso, Tim Dodwell, and Rob Scheichl. Continuous level Monte Carlo and sample-adaptive model hierarchies. SIAM/ASA Journal on Uncertainty Quantification, 7(1):93–116, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

DElia:2018:EGS

[DEH^{+}18] M. D’Elia, H. C. Edwards, J. Hu, E. Phipps, and S. Rajamanickam. Ensemble grouping strategies for embedded stochastic collocation methods applied to anisotropic diffusion problems. SIAM/ASA Journal on Uncertainty Quantification, 6(1):87–117, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Djurdjevac:2018:ESF

[DEKR18] Ana Djurdjevac, Charles M. Elliott, Ralf Kornhuber, and Thomas Ranner. Evolving surface finite element methods for random advection–diffusion equations. SIAM/ASA Journal on Uncertainty Quantification, 6(4):1656–1684, ???? 2018.

CODEN SJUQA3. ISSN 2166-2525.

Ditkowski:2020:DEU

[DFS20] Adi Ditkowski, Gadi Fibich, and Amir Sagiv. Density estimation in uncertainty propagation problems using a surrogate model. SIAM/ASA Journal on Uncertainty Quantification, 8(1):261–300, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

DElia:2013:CGS

[DG13] Marta D’Elia and Max Gunzburger. Coarse-grid sampling interpolatory methods for approximating Gaussian random fields. SIAM/ASA Journal on Uncertainty Quantification, 1 (1):270–296, ???? 2013. CO-DEN SJUQA3. ISSN 2166-2525.

Dasgupta:2019:FPE

[DG19] Agnimitra Dasgupta and Debraj Ghosh. Failure probability estimation of linear time varying systems by progressive refinement of reduced order models. SIAM/ASA Journal on Uncertainty Quantification, 7(3): 1007–1028, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Dematteis:2019:EEQ

[DGVE19] Giovanni Dematteis, Tobias Grafke, and Eric Vanden-Eijnden. Extreme event quantification in dynamical systems with random components. SIAM/ASA Journal on Uncertainty Quantification, 7(3):

1029–1059, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

deHoop:2023:CRL

[dHKNS23] Maarten V. de Hoop, Nikola B. Kovachki, Nicholas H. Nelsen, and Andrew M. Stuart. Convergence rates for learning linear operators from noisy data. SIAM/ASA Journal on Uncertainty Quantification, 11(2):480–513, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1442942.

Dia:2023:CMB

[Dia23] Ben Mansour Dia. A continuation method in Bayesian inference. SIAM/ASA Journal on Uncertainty Quantification, 11(2):646–681, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/19M130251X.

Djurdjevac:2021:LPP

[Dju21]  Ana Djurdjevac.  Linear

parabolic problems in random moving domains. SIAM/ASA Journal on Uncertainty Quantification, 9(2):848–879, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Dolgov:2015:PCE

[DKLM15] Sergey Dolgov, Boris N. Khoromskij, Alexander Litvinenko, and Hermann G. Matthies. Polynomial chaos expansion of random coefficients and the solution of stochastic partial differential equations in the tensor train format. SIAM/ ASA Journal on Uncertainty Quantification, 3(1):1109–1135, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Dupuis:2016:PSI

[DKPP16] Paul Dupuis, Markos A. Katsoulakis, Yannis Pantazis, and Petr Plechác. Path-space information bounds for uncertainty quantification and sensitivity analysis of stochastic dynamics. SIAM/ASA Journal on Uncertainty Quantification, 4(1):80–111, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Dodwell:2015:HMM

[DKST15] T. J. Dodwell, C. Ketelsen, R. Scheichl, and A. L. Teckentrup. A hierarchical multilevel Markov chain Monte Carlo algorithm with applications to uncertainty quantification in subsurface flow. SIAM/ ASA Journal on Uncertainty Quantification, 3(1):1075–1108, ???? 2015. CODEN SJUQA3. ISSN 2166-2525. See erratum [DKST19].

Dodwell:2019:EHM

[DKST19] T. J. Dodwell, C. Ketelsen, R. Scheichl, and A. L. Teckentrup. ERRATUM: A Hierarchical Multilevel Markov Chain Monte Carlo Algorithm with Applications to Uncertainty Quantification in Subsurface Flow. SIAM/ASA Journal on Uncertainty Quantification, 7(4):1398–1399, ???? 2019. CODEN SJUQA3. ISSN 2166-2525. See [DKST15].

Dick:2016:HOQ

[DLS16] Josef Dick, Quoc T. Le Gia, and Christoph Schwab. Higher order quasi-Monte Carlo integration for holomorphic, parametric operator equations. SIAM/ASA Journal on Uncertainty Quantification, 4(1): 48–79, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Dick:2022:EPL

[DLS22] Josef Dick, Marcello Longo, and Christoph Schwab. Extrapolated polynomial lattice rule integration in computational uncertainty quantification. SIAM/ASA Journal on Uncertainty Quantification, 10(2):651–686, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1338137.

Dobson:2021:UCM

[DLZ21] Matthew Dobson, Yao Li, and Jiayu Zhai. Using coupling methods to estimate sample quality of stochastic differential equations. SIAM/ASA Journal on Uncertainty Quantification, 9(1):135–162, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

DelMoral:2015:SMC

[DM15] Pierre Del Moral and Lawrence M. Murray. Sequential Monte Carlo with highly informative observations. SIAM/ASA Journal on Uncertainty Quantification, 3(1):969–997, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

DeLozzo:2016:EDB

[DM16] Matthias De Lozzo and Amandine Marrel. Estimation of the derivative-based global sensitivity measures using a Gaussian process metamodel. SIAM/ASA Journal on Uncertainty Quantification, 4(1): 708–738, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Deveney:2023:DSA

[DMS23] Teo Deveney, Eike H. Mueller, and Tony Shardlow. Deep surrogate accelerated delayedacceptance Hamiltonian Monte Carlo for Bayesian inference of spatio-temporal heat fluxes in rotating disc systems. SIAM/ ASA Journal on Uncertainty Quantification, 11(3):970–995, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M1513113.

Drovandi:2021:SMH

[DNP21] Christopher Drovandi, David J.

Nott, and Daniel E. Pagendam. A semiautomatic method for history matching using sequential Monte Carlo. SIAM/ASA Journal on Uncertainty Quantification, 9(3):1034–1063, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Dolz:2020:HOP

[Döl20] Jürgen Dölz. A higher order perturbation approach for electromagnetic scattering problems on random domains. SIAM/ASA Journal on Uncertainty Quantification, 8(2): 748–774, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Despres:2016:UPI

[DP16] Bruno Despres and Benoit Perthame. Uncertainty propagation; intrusive kinetic formulations of scalar conservation laws. SIAM/ASA Journal on Uncertainty Quantification, 4(1):980–1013, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Durante:2018:SGM

[DPP18] Joseph Durante, Raj Patel, and Warren B. Powell. Scenario generation methods that replicate crossing times in spatially distributed stochastic systems. SIAM/ASA Journal on Uncertainty Quantification, 6(2):596– 626, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

delRosario:2019:LVD

[dRLI19] Zachary del Rosario, Minyong Lee, and Gianluca Iaccarino. Lurking variable detection via dimensional analysis. SIAM/ASA Journal on Uncertainty Quantification, 7(1): 232–259, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Du:2017:RAC

[DS17] Hailiang Du and Leonard A.

Smith. Rising above chaotic likelihoods. SIAM/ASA Journal on Uncertainty Quantification, 5(1):246–258, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Dolgov:2019:HAL

[DS19] Sergey Dolgov and Robert Scheichl. A hybrid alternating least squares–TT-cross algorithm for parametric PDEs. SIAM/ASA Journal on Uncertainty Quantification, 7(1): 260–291, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Dinner:2020:SGV

[DTVW20] Aaron R. Dinner, Erik H.

Thiede, Brian Van Koten, and Jonathan Weare. Stratification as a general variance reduction method for Markov Chain Monte Carlo. SIAM/ ASA Journal on Uncertainty Quantification, 8(3):1139–1188, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Dunlop:2021:SGP

[DY21] Matthew M. Dunlop and Yunan Yang. Stability of Gibbs posteriors from the Wasserstein loss for Bayesian full waveform inversion. SIAM/ASA Journal on Uncertainty Quantification, 9(4):1499–1526, ???? 2021. CODEN SJUQA3. ISSN 2166-

Dette:2021:RKH

[DZ21] Holger Dette and Anatoly A. Zhigljavsky. Reproducing kernel Hilbert spaces, polynomials, and the classical moment problem. SIAM/ASA Journal on Uncertainty Quantification, 9 (4):1589–1614, ???? 2021. CO-DEN SJUQA3. ISSN 2166-2525.

Easley:2021:HOU

[EB21] Deanna C. Easley and Tyrus Berry. A higher order unscented transform. SIAM/ASA Journal on Uncertainty Quantification, 9(3):1094–1131, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Everink:2023:BIP

[EDA23] Jasper M. Everink, Yiqiu Dong, and Martin S. Andersen. Bayesian inference with projected densities. SIAM/ASA Journal on Uncertainty Quantification, 11(3):1025–1043, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M150695X.

Elie-Dit-Cosaque:2022:GOS

[EDCMD22] Kevin Elie-Dit-Cosaque and Veronique Maume-Deschamps. Goal-oriented Shapley effects with special attention to the quantile-oriented case. SIAM/ASA Journal on Uncertainty Quantification, 10 (3):1037–1069, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1395247.

Edeling:2023:DAS

[Ede23]  Wouter Edeling.  On the

deep active-subspace method. SIAM/ASA Journal on Uncertainty Quantification, 11(1): 62–90, ???? 2023. CODEN SJUQA3. ISSN 2166-2525.

URL https://epubs.siam.
org/doi/10.1137/21M1463240.

Elfverson:2014:UQA

[EEHM14] Daniel Elfverson, Donald J. Estep, Fredrik Hellman, and Axel Målqvist. Uncertainty quantification for approximate pquantiles for physical models with stochastic inputs. SIAM/ ASA Journal on Uncertainty Quantification, 2(1):826–850, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Ettehad:2021:ICO

[EF21] Mahmood Ettehad and Simon Foucart. Instances of computational optimal recovery: Dealing with observation errors. SIAM/ASA Journal on Uncertainty Quantification, 9(4): 1438–1456, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Ehre:2023:CDR

[EFF^{+}23] Max Ehre, Rafael Flock, Martin Fußeder, Iason Papaioannou, and Daniel Straub. Certified dimension reduction for Bayesian updating with the cross-entropy method. SIAM/ ASA Journal on Uncertainty Quantification, 11(1):358–388, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M1484031.

Elfverson:2016:MMC

[EHM16] Daniel Elfverson, Fredrik Hellman, and Axel Målqvist. A multilevel Monte Carlo method for computing failure probabilities. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):312–330, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Elman:2013:RBC

[EL13] Howard C. Elman and Qifeng Liao. Reduced basis collocation methods for partial differential equations with random coefficients. SIAM/ASA Journal on Uncertainty Quantification, 1(1):192–217, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Eigel:2016:LEE

[EM16] Martin Eigel and Christian Merdon. Local equilibration error estimators for guaranteed error control in adaptive stochastic higher-order Galerkin finite element methods. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):1372–1397, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Eigel:2020:ASG

[EMM20] Martin Eigel, Manuel Marschall, and Michael Multerer. An adaptive stochastic Galerkin tensor train discretization for randomly perturbed domains. SIAM/ASA Journal on Uncertainty Quantification, 8(3): 1189–1214, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Eigel:2016:AMM

[EMN16] Martin Eigel, Christian Merdon, and Johannes Neumann. An adaptive multilevel Monte Carlo method with stochastic bounds for quantities of interest with uncertain data. SIAM/ ASA Journal on Uncertainty Quantification, 4(1):1219–1245, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Ernst:2022:WSR

[EPS22] Oliver G. Ernst, Alois Pichler, and Björn Sprungk. Wasserstein sensitivity of risk and uncertainty propagation. SIAM/ ASA Journal on Uncertainty Quantification, 10(3):915–948, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1325459.

Ernst:2015:AEP

[ESS15] Oliver G. Ernst, Björn Sprungk, and Hans-Jörg Starkloff. Analysis of the ensemble and polynomial chaos Kalman filters in Bayesian inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 3(1): 823–851, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Fillion:2020:IEK

[FBG^{+}20] Anthony Fillion, Marc Bocquet, Serge Gratton, Selime Gürol, and Pavel Sakov. An iterative ensemble Kalman smoother in presence of additive model error. SIAM/ASA Journal on Uncertainty Quantification, 8(1):198–228, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Fang:2021:ISP

[FG21] Wei Fang and Mike B. Giles. Importance sampling for pathwise sensitivity of stochastic chaotic systems. SIAM/ASA Journal on Uncertainty Quantification, 9(3):1217–1241, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Fohring:2016:AOE

[FH16] J. Fohring and E. Haber. Adaptive A-optimal experimental design for linear dynamical systems. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):1138–1159, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Fadikar:2018:CSA

[FHC^{+}18] Arindam Fadikar, Dave Higdon, Jiangzhuo Chen, Bryan Lewis, Srinivasan Venkatramanan, and Madhav Marathe. Calibrating a stochastic, agentbased model using quantilebased emulation. SIAM/ASA Journal on Uncertainty Quantification, 6(4):1685–1706, ????

2018. CODEN SJUQA3. ISSN 2166-2525.

Fort:2021:GSA

[FKL21] Jean-Claude Fort, Thierry Klein, and Agnès Lagnoux. Global sensitivity analysis and Wasserstein spaces. SIAM/ ASA Journal on Uncertainty Quantification, 9(2):880–921, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Feng:2015:ENM

[FLL15] Xiaobing Feng, Junshan Lin, and Cody Lorton. An efficient numerical method for acoustic wave scattering in random media. SIAM/ASA Journal on Uncertainty Quantification, 3 (1):790–822, ???? 2015. CO-DEN SJUQA3. ISSN 2166-2525.

Fajraoui:2017:SDE

[FMS17] Noura Fajraoui, Stefano Marelli, and Bruno Sudret. Sequential design of experiment for sparse polynomial chaos expansions. SIAM/ASA Journal on Uncertainty Quantification, 5 (1):1061–1085, ???? 2017. CO-DEN SJUQA3. ISSN 2166-2525.

Fox:2016:FSL

[FN16] Colin Fox and Richard A. Norton. Fast sampling in a linear-Gaussian inverse problem. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):1191–1218, ???? 2016. CO-DEN SJUQA3. ISSN 2166-

Frenklach:2016:CSD

[FPGD^{+}16] Michael Frenklach, Andrew Packard, Gonzalo Garcia-Donato, Rui Paulo, and Jerome Sacks. Comparison of statistical and deterministic frameworks of uncertainty quantification. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):875–901, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Francom:2022:LWE

[FSK22] Devin Francom, Bruno Sansó, and Ana Kupresanin. Landmarkwarped emulators for models with misaligned functional response. SIAM/ASA Journal on Uncertainty Quantification, 10(1):125–150, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M135279X.

Giles:2018:MEE

[GB18] Michael B. Giles and Francisco Bernal. Multilevel estimation of expected exit times and other functionals of stopped diffusions. SIAM/ASA Journal on Uncertainty Quantification, 6 (4):1454–1474, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Ginsbourger:2014:BAR

[GBC^{+}14] David Ginsbourger, Jean Baccou, Clément Chevalier, Frédéric Perales, Nicolas Garland, and Yann Monerie. Bayesian adaptive reconstruction of profile optima and optimizers. SIAM/ ASA Journal on Uncertainty Quantification, 2(1):490–510, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Golchi:2015:MEC

[GBCC15] S. Golchi, D. R. Bingham, H. Chipman, and D. A. Campbell. Monotone emulation of computer experiments. SIAM/ ASA Journal on Uncertainty Quantification, 3(1):370–392, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Goncharov:2023:NPL

[GBD23] Fedor Goncharov, Éric Barat, and Thomas Dautremer. Nonparametric posterior learning for emission tomography. SIAM/ASA Journal on Uncertainty Quantification, 11(2):452–479, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1463367.

Grudzien:2018:AFU

[GCB18] Colin Grudzien, Alberto Carrassi, and Marc Bocquet. Asymptotic forecast uncertainty and the unstable subspace in the presence of additive model error. SIAM/ASA Journal on Uncertainty Quantification, 6(4):1335–1363, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Glaubitz:2023:GSB

[GGS23] Jan Glaubitz, Anne Gelb, and Guohui Song. Generalized sparse Bayesian learning and application to image reconstruction. SIAM/ ASA Journal on Uncertainty Quantification, 11(1):262–284, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M147236X.

Giles:2019:MNS

[GHA19] Michael B. Giles and Abdul-Lateef Haji-Ali. Multilevel nested simulation for efficient risk estimation. SIAM/ASA Journal on Uncertainty Quantification, 7(2):497–525, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Geiersbach:2016:OAF

[GHT16] Caroline Geiersbach, Clemens Heitzinger, and Gerhard Tulzer. Optimal approximation of the first-order corrector in multiscale stochastic elliptic PDE. SIAM/ASA Journal on Uncertainty Quantification, 4(1): 1246–1262, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Gunzburger:2019:EFR

[GIMS19] M. Gunzburger, T. Iliescu, M. Mohebujjaman, and M. Schneier. An evolve-filter-relax stabilized reduced order stochastic collocation method for the timedependent Navier–Stokes equations. SIAM/ASA Journal on Uncertainty Quantification, 7 (4):1162–1184, ???? 2019. CO-DEN SJUQA3. ISSN 2166-2525.

Galindo:2016:ASC

[GJWZ16] D. Galindo, P. Jantsch, C. G.

Webster, and G. Zhang. Accelerating stochastic collocation methods for partial differential equations with random input data. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):1111–1137, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Grigo:2019:BMD

[GK19] Constantin Grigo and Phaedon-Stelios Koutsourelakis. Bayesian model and dimension reduction for uncertainty propagation: Applications in random media. SIAM/ASA Journal on Uncertainty Quantification, 7 (1):292–323, ???? 2019. CO-DEN SJUQA3. ISSN 2166-2525.

Giordano:2020:BMT

[GK20] Matteo Giordano and Hanne Kekkonen. Bernstein–von Mises theorems and uncertainty quantification for linear inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 8(1):342–373, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Guth:2021:QMC

[GKK^{+}21] Philipp A. Guth, Vesa Kaarnioja, Frances Y. Kuo, Claudia Schillings, and Ian H. Sloan. A quasi-Monte Carlo method for optimal control under uncertainty. SIAM/ASA Journal on Uncertainty Quantification, 9(2):354–383, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Gamboa:2018:SAB

[GKL18] Fabrice Gamboa, Thierry Klein, and Agnès Lagnoux. Sensitivity analysis based on Cramér–von Mises distance. SIAM/ASA Journal on Uncertainty Quantification, 6(2): 522–548, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Ganesh:2021:QMC

[GKS21] M. Ganesh, Frances Y. Kuo, and Ian H. Sloan. Quasi-Monte Carlo finite element analysis for wave propagation in heterogeneous random media. SIAM/ASA Journal on Uncertainty Quantification, 9 (1):106–134, ???? 2021. CO-DEN SJUQA3. ISSN 2166-2525.

Gorodetsky:2016:MKI

[GM16] Alex Gorodetsky and Youssef Marzouk. Mercer kernels and integrated variance experimental design: Connections between Gaussian process regression and polynomial approximation. SIAM/ASA Journal on Uncertainty Quantification, 4(1):796–828, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Gilpin:2022:CCP

[GMC22] Shay Gilpin, Tomoko Matsuo, and Stephen E. Cohn. Continuum covariance propagation for understanding variance loss in advective systems. SIAM/ ASA Journal on Uncertainty Quantification, 10(3):886–914, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1442449.

Grote:2022:UQM

[GMN22] Marcus J. Grote, Simon Michel, and Fabio Nobile. Uncertainty quantification by multilevel Monte Carlo and local time-stepping for wave propagation. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1601–1628, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1429047.

Giles:2015:MMC

[GNR15] Michael B. Giles, Tigran Nagapetyan, and Klaus Ritter. Multilevel Monte Carlo approximation of distribution functions and densities. SIAM/ASA Journal on Uncertainty Quantification, 3(1):267–295, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Gramacy:2014:MPA

[GNW14] Robert B. Gramacy, Jarad Niemi, and Robin M. Weiss. Massively parallel approximate Gaussian process regression. SIAM/ASA Journal on Uncertainty Quantification, 2(1):564– 584, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Gauthier:2014:SAI

[GP14a] Bertrand Gauthier and Luc Pronzato. Spectral approximation of the IMSE criterion for optimal designs in kernel-based interpolation models. SIAM/ ASA Journal on Uncertainty Quantification, 2(1):805–825, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Gordon:2014:PFD

[GP14b] Andrew Gordon and Catherine E. Powell. A preconditioner for fictitious domain formulations of elliptic PDEs on uncertain parameterized domains. SIAM/ASA Journal on Uncertainty Quantification, 2(1):622– 646, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Gantner:2018:HOQ

[GP18] R. N. Gantner and M. D. Peters. Higher-order quasi-Monte Carlo for Bayesian shape inversion. SIAM/ASA Journal on Uncertainty Quantification, 6(2):707–736, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Griebel:2017:RKH

[GR17] Michael Griebel and Christian Rieger. Reproducing kernel Hilbert spaces for parametric partial differential equations. SIAM/ASA Journal on Uncertainty Quantification, 5(1):111– 137, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Grigoriu:2014:RSR

[Gri14] M. Grigoriu. Response statistics for random heterogeneous microstructures. SIAM/ASA Journal on Uncertainty Quantification, 2(1):252–275, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Grigoriu:2016:MMM

[Gri16]  M. Grigoriu.  Microstruc-

ture models and material response by extreme value theory. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):190–217, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Grigoriu:2017:ESR

[Gri17] M. Grigoriu. Estimates of system response maxima by extreme value theory and surrogate models. SIAM/ASA Journal on Uncertainty Quantification, 5(1):922–955, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Giles:2022:ANM

[GSM22] Michael Giles and Oliver Sheridan-Methven. Analysis of nested multilevel Monte Carlo using approximate normal random variables. SIAM/ ASA Journal on Uncertainty Quantification, 10(1):200–226, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1399385.

Gratton:2014:SCM

[GTP14] Serge Gratton and David Titley-Peloquin. Stochastic conditioning of matrix functions. SIAM/ASA Journal on Uncertainty Quantification, 2 (1):763–783, ???? 2014. CO-DEN SJUQA3. ISSN 2166-2525.

Gu:2018:SGS

[GW18] Mengyang Gu and Long Wang. Scaled Gaussian stochastic process for computer model calibration and prediction. SIAM/ ASA Journal on Uncertainty Quantification, 6(4):1555–1583, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Gu:2022:TFS

[GXW22] Mengyang Gu, Fangzheng Xie, and Long Wang. A theoretical framework of the scaled Gaussian stochastic process in prediction and calibration. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1435–1460, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1409949.

Houssineau:2018:SFC

[HB18] Jeremie Houssineau and Adrian N. Bishop. Smoothing and filtering with a class of outer measures. SIAM/ASA Journal on Uncertainty Quantification, 6 (2):845–866, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Hobbs:2017:SBU

[HBC^{+}17] Jonathan Hobbs, Amy Braverman, Noel Cressie, Robert Granat, and Michael Gunson. Simulation-based uncertainty quantification for estimating atmospheric CO_{2} from satellite data. SIAM/ASA Journal on Uncertainty Quantification, 5(1):956–985, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Hart:2019:GSA

[HBC19] Joseph L. Hart, Julie Bessac, and Emil M. Constantinescu. Global sensitivity analysis for statistical model parameters. SIAM/ASA Journal on Uncertainty Quantification, 7(1):67– 92, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

He:2018:IIG

[HC18] Xu He and Peter Chien. On the instability issue of gradientenhanced Gaussian process emulators for computer experiments. SIAM/ASA Journal on Uncertainty Quantification, 6(2):627–644, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Heas:2020:SRM

[Héa20] P. Héas. Selecting reduced models in the cross-entropy method. SIAM/ASA Journal on Uncertainty Quantification, 8(2):511–538, ???? 2020. CO-DEN SJUQA3. ISSN 2166-

Hessling:2013:DSP

[Hes13] Jan Peter Hessling. Deterministic sampling for propagating model covariance. SIAM/ASA Journal on Uncertainty Quantification, 1(1):297–318, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Hessling:2014:ICM

[Hes14] Jan Peter Hessling. Identification of complex models. SIAM/ASA Journal on Uncertainty Quantification, 2(1):717– 744, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Huang:2017:NMS

[HG17] Ya-Ting Huang and James Glimm. A novel methodology of stochastic short term forecasting of cloud boundaries. SIAM/ ASA Journal on Uncertainty Quantification, 5(1):1279–1294, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Hart:2019:RSI

[HG19] Joseph L. Hart and Pierre A. Gremaud. Robustness of the Sobol’ indices to marginal distribution uncertainty. SIAM/ ASA Journal on Uncertainty Quantification, 7(4):1224–1244, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Hironaka:2020:MMC

[HGGT20] Tomohiko Hironaka, Michael B. Giles, Takashi Goda, and Howard Thom. Multilevel Monte Carlo estimation of the expected value of sample information. SIAM/ASA Journal on Uncertainty Quantification, 8 (3):1236–1259, ???? 2020. CO-DEN SJUQA3. ISSN 2166-2525.

Heaton:2013:MUC

[HGS13] Matthew J. Heaton, Tamara A.

Greasby, and Stephan R. Sain. Modeling uncertainty in climate using ensembles of regional and global climate models and multiple observationbased data sets. SIAM/ASA Journal on Uncertainty Quantification, 1(1):535–559, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Hagemann:2022:SNF

[HHS22] Paul Hagemann, Johannes Hertrich, and Gabriele Steidl. Stochastic normalizing flows for inverse problems: a Markov chains viewpoint. SIAM/ASA Journal on Uncertainty Quantification, 10(3):1162–1190, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1450604.

Heng:2023:UED

[HJLT23] Jeremy Heng, Ajay Jasra, Kody J. H. Law, and Alexander Tarakanov. On unbiased estimation for discretized models. SIAM/ASA Journal on Uncertainty Quantification, 11(2):616–645, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1460788.

Hall:2018:RID

[HK18] Eric Joseph Hall and Markos A. Katsoulakis. Robust information divergences for model-form uncertainty arising from sparse data in random PDE. SIAM/ ASA Journal on Uncertainty Quantification, 6(4):1364–1394, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Heinkenschloss:2020:ARO

[HKT20] Matthias Heinkenschloss, Boris Kramer, and Timur Takhtaganov. Adaptive reduced-order model construction for conditional value-at-risk estimation. SIAM/ ASA Journal on Uncertainty Quantification, 8(2):668–692, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Heinkenschloss:2018:CVR

[HKTW18] Matthias Heinkenschloss, Boris Kramer, Timur Takhtaganov, and Karen Willcox. Conditionalvalue-at-risk estimation via reduced-order models. SIAM/ ASA Journal on Uncertainty Quantification, 6(4):1395–1423, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Hu:2018:BIFa

[HKZ18a] Grace X. Hu, David R. Kuipers, and Yong Zeng. Bayesian inference via filtering equations for ultrahigh frequency data (I): Model and estimation. SIAM/ASA Journal on Uncertainty Quantification, 6 (1):34–60, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Hu:2018:BIFb

[HKZ18b] Grace X. Hu, David R. Kuipers, and Yong Zeng. Bayesian inference via filtering equations for ultrahigh frequency data (II): Model selection. SIAM/ASA Journal on Uncertainty Quantification, 6(1): 61–86, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Hyvonen:2015:SGF

[HL15] N. Hyvönen and M. Leinonen.

Stochastic Galerkin finite element method with local conductivity basis for electrical impedance tomography. SIAM/ ASA Journal on Uncertainty Quantification, 3(1):998–1019, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Hu:2017:SDR

[HL17] Ruimeng Hu and Mike Ludkovski. Sequential design for ranking response surfaces. SIAM/ASA Journal on Uncertainty Quantification, 5(1): 212–239, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Hegde:2018:CAM

[HLO^{+}18] Arun Hegde, Wenyu Li, James Oreluk, Andrew Packard, and Michael Frenklach. Consistency analysis for massively inconsistent datasets in boundto-bound data collaboration. SIAM/ASA Journal on Uncertainty Quantification, 6(2):429– 456, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Heitzinger:2018:EUC

[HLPR18] Clemens Heitzinger, Michael Leumüller, Gudmund Pammer, and Stefan Rigger. Existence, uniqueness, and a comparison of nonintrusive methods for the stochastic nonlinear Poisson– Boltzmann equation. SIAM/ ASA Journal on Uncertainty Quantification, 6(3):1019–1042, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Hooker:2015:CTE

[HLR15] Giles Hooker, Kevin K. Lin, and Bruce Rogers. Control theory and experimental design in diffusion processes. SIAM/ASA Journal on Uncertainty Quantification, 3(1):234–264, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Hu:2022:GFR

[HLY22] Mengqi Hu, Yifei Lou, and Xiu Yang. A general framework of rotational sparse approximation in uncertainty quantification. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1410–1434, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1391602.

Hung:2020:LID

[HML20] Ying-Chao Hung, George Michailidis, and Horace PakHai Lok. Locating infinite discontinuities in computer experiments. SIAM/ASA Journal on Uncertainty Quantification, 8 (2):717–747, ???? 2020. CO-DEN SJUQA3. ISSN 2166-2525.

Higham:2013:MET

[HMR^{+}13] Desmond J. Higham, Xuerong Mao, Mikolaj Roj, Qingshuo Song, and George Yin. Mean exit times and the multilevel Monte Carlo method. SIAM/ ASA Journal on Uncertainty Quantification, 1(1):2–18, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Hosseini:2017:WPBa

[HN17] Bamdad Hosseini and Nilima Nigam. Well-posed Bayesian inverse problems: Priors with exponential tails. SIAM/ASA Journal on Uncertainty Quantification, 5(1):436–465, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Haberstich:2023:ALT

[HNP23] Cécile Haberstich, A. Nouy, and G. Perrin. Active learning of tree tensor networks using optimal least squares. SIAM/ ASA Journal on Uncertainty Quantification, 11(3):848–876, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1415911.

Hosseini:2017:WPBb

[Hos17] Bamdad Hosseini. Well-posed Bayesian inverse problems with infinitely divisible and heavytailed prior measures. SIAM/ ASA Journal on Uncertainty Quantification, 5(1):1024–1060, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Hosseini:2019:TMH

[Hos19]  Bamdad Hosseini.  Two

Metropolis–Hastings algorithms for posterior measures with non-Gaussian priors in infinite dimensions. SIAM/ASA Journal on Uncertainty Quantification, 7(4):1185–1223, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Horwood:2014:GMD

[HP14] Joshua T. Horwood and Aubrey B. Poore. Gauss von Mises distribution for improved uncertainty realism in space situational awareness. SIAM/ ASA Journal on Uncertainty Quantification, 2(1):276–304, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Heitzinger:2018:CFM

[HPR18] Clemens Heitzinger, Gudmund Pammer, and Stefan Rigger. Cubature formulas for multisymmetric functions and applications to stochastic partial differential equations. SIAM/ASA Journal on Uncertainty Quantification, 6(1):213–242, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Harbrecht:2016:MAQ

[HPS16] Helmut Harbrecht, Michael Peters, and Markus Siebenmorgen. Multilevel accelerated quadrature for PDEs with lognormally distributed diffusion coefficient. SIAM/ASA Journal on Uncertainty Quantification, 4(1):520–551, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Hu:2021:UQB

[HPW21] Jingwei Hu, Lorenzo Pareschi, and Yubo Wang. Uncertainty quantification for the BGK model of the Boltzmann equation using multilevel variance reduced Monte Carlo methods. SIAM/ASA Journal on Uncertainty Quantification, 9(2):650– 680, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Hoang:2021:MMC

[HQS21] Viet Ha Hoang, Jia Hao Quek, and Christoph Schwab. Multilevel Markov chain Monte Carlo for Bayesian inversion of parabolic partial differential equations under Gaussian prior. SIAM/ASA Journal on Uncertainty Quantification, 9 (2):384–419, ???? 2021. CO-DEN SJUQA3. ISSN 2166-2525.

Hadji:2021:CWT

[HS21] Amine Hadji and Botond Szabó. Can we trust Bayesian uncertainty quantification from Gaussian process priors with squared exponential covariance kernel? SIAM/ASA Journal on Uncertainty Quantification, 9(1):185–230, ???? 2021. CO-DEN SJUQA3. ISSN 2166-

Harlim:2020:KMB

[HSAY20] John Harlim, Daniel Sanz-Alonso, and Ruiyi Yang. Kernel methods for Bayesian elliptic inverse problems on manifolds. SIAM/ASA Journal on Uncertainty Quantification, 8 (4):1414–1445, ???? 2020. CO-DEN SJUQA3. ISSN 2166-2525.

He:2022:ATR

[HSH^{+}22] Yuchen He, Namjoon Suh, Xiaoming Huo, Sung Ha Kang, and Yajun Mei. Asymptotic theory of ℓ_{1}-regularized PDE identification from a single noisy trajectory. SIAM/ASA Journal on Uncertainty Quantification, 10(3):1012–1036, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1398884.

Huan:2018:CSC

[HSS^{+}18] Xun Huan, Cosmin Safta, Khachik Sargsyan, Zachary P. Vane, Guilhem Lacaze, Joseph C. Oefelein, and Habib N. Najm. Compressive sensing with crossvalidation and stop-sampling for sparse polynomial chaos expansions. SIAM/ASA Journal on Uncertainty Quantification, 6(2):907–936, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Herrmann:2019:UQS

[HSZ19] Lukas Herrmann, Christoph Schwab, and Jakob Zech. Uncertainty quantification for spectral fractional diffusion: Sparsity analysis of parametric solutions. SIAM/ASA Journal on Uncertainty Quantification, 7(3):913–947, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Haasdonk:2013:RBM

[HUW13] Bernard Haasdonk, Karsten Urban, and Bernhard Wieland. Reduced basis methods for parameterized partial differential equations with stochastic influences using the Karhunen– Loève expansion. SIAM/ASA Journal on Uncertainty Quantification, 1(1):79–105, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Haaland:2018:FCS

[HWM18] Benjamin Haaland, Wenjia Wang, and Vaibhav Maheshwari. A framework for controlling sources of inaccuracy in Gaussian process emulation of deterministic computer experiments. SIAM/ASA Journal on Uncertainty Quantification, 6(2):497–521, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Hao:2021:NAO

[HZ21] Zhaopeng Hao and Zhongqiang Zhang. Numerical approximation of optimal convergence for fractional elliptic equations with additive fractional Gaussian noise. SIAM/ASA Journal on Uncertainty Quantification, 9(3):1013–1033, ???? 2021.

CODEN SJUQA3. ISSN 2166-2525.

Iolov:2017:ODE

[IDL17] Alexandre Iolov, Susanne Ditlevsen, and André Longtin. Optimal design for estimation in diffusion processes from first hitting times. SIAM/ASA Journal on Uncertainty Quantification, 5(1):88–110, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Ibrahima:2017:MDS

[IT17] Fayadhoi Ibrahima and Hamdi A. Tchelepi. Multipoint distribution of saturation for stochastic nonlinear two-phase transport. SIAM/ASA Journal on Uncertainty Quantification, 5 (1):353–377, ???? 2017. CO-DEN SJUQA3. ISSN 2166-2525.

Iza-Teran:2019:GML

[ITG19] Rodrigo Iza-Teran and Jochen Garcke. A geometrical method for low-dimensional representations of simulations. SIAM/ ASA Journal on Uncertainty Quantification, 7(2):472–496, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Jahn:2023:NLF

[Jah23] Tim Jahn. Noise level free regularization of general linear inverse problems under unconstrained white noise. SIAM/ ASA Journal on Uncertainty Quantification, 11(2):591–615, ???? 2023. CODEN SJUQA3.

ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M1506675.

Jiang:2016:GOR

[JCN16] Jiahua Jiang, Yanlai Chen, and Akil Narayan. A goal-oriented reduced basis methods-accelerated generalized polynomial chaos algorithm. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):1398–1420, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Jadamba:2021:COF

[JKS^{+}21] Baasansuren Jadamba, Akhtar A.

Khan, Miguel Sama, Hans-Jorg Starkloff, and Christiane Tammer. A convex optimization framework for the inverse problem of identifying a random parameter in a stochastic partial differential equation. SIAM/ ASA Journal on Uncertainty Quantification, 9(2):922–952, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Jimenez:2017:NPC

[JLK17] M. Navarro Jimenez, O. P. Le Maître, and O. M. Knio. Nonintrusive polynomial chaos expansions for sensitivity analysis in stochastic differential equations. SIAM/ASA Journal on Uncertainty Quantification, 5 (1):378–402, ???? 2017. CO-DEN SJUQA3. ISSN 2166-

Jin:2013:OIS

[JM13] Qinian Jin and Peter Mathé. Oracle inequality for a statistical Raus–Gfrerer-type rule. SIAM/ASA Journal on Uncertainty Quantification, 1(1):386– 407, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Jahnke:2022:MSC

[JS22] Tobias Jahnke and Benny Stein. A multilevel stochastic collocation method for Schrödinger equations with a random potential. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1753–1780, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1440517.

Jarman:2013:CCS

[JT13] K. D. Jarman and A. M. Tartakovsky. A comparison of closures for stochastic advectiondiffusion equations. SIAM/ ASA Journal on Uncertainty Quantification, 1(1):319–347, ???? 2013. CODEN SJUQA3. ISSN 2166-2525. See erratum [JT15].

Jarman:2015:ECC

[JT15] K. D. Jarman and A. M. Tartakovsky. Erratum: A Comparison of Closures for Stochastic Advection-Diffusion Equations. SIAM/ASA Journal on Uncertainty Quantification, 3(1):265– 266, ???? 2015. CODEN SJUQA3. ISSN 2166-2525. See [JT13].

Jin:2021:SPS

[JZZ21] Bangti Jin, Zehui Zhou, and Jun Zou. On the saturation phenomenon of stochastic gradient descent for linear inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 9(4):1553–1588, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Kantas:2014:SMC

[KBJ14] Nikolas Kantas, Alexandros Beskos, and Ajay Jasra. Sequential Monte Carlo methods for high-dimensional inverse problems: a case study for the Navier–Stokes equations. SIAM/ASA Journal on Uncertainty Quantification, 2(1): 464–489, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Kenz:2013:CFB

[KBS13] Zackary R. Kenz, H. T. Banks, and Ralph C. Smith. Comparison of frequentist and Bayesian confidence analysis methods on a viscoelastic stenosis model. SIAM/ASA Journal on Uncertainty Quantification, 1(1):348– 369, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Kyzyurova:2018:CCM

[KBW18] Ksenia N. Kyzyurova, James O.

Berger, and Robert L. Wolpert. Coupling computer models through linking their statistical emulators. SIAM/ASA Journal on Uncertainty Quantification, 6(3):1151–1171, ???? 2018.

CODEN SJUQA3. ISSN 2166-2525.

Kerleguer:2023:MSM

[Ker23] Baptiste Kerleguer. Multifidelity surrogate modeling for time-series outputs. SIAM/ ASA Journal on Uncertainty Quantification, 11(2):514–539, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1386694.

Katzfuss:2022:SVA

[KGL22] Matthias Katzfuss, Joseph Guinness, and Earl Lawrence. Scaled Vecchia approximation for fast computer-model emulation. SIAM/ASA Journal on Uncertainty Quantification, 10(2):537–554, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1352156.

Kang:2016:KAR

[KJ16] Lulu Kang and V. Roshan Joseph. Kernel approximation: From regression to interpolation. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):112–129, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Kouri:2022:RAO

[KJH22] Drew P. Kouri, John D. Jakeman, and J. Gabriel Huerta. Risk-adapted optimal experimental design. SIAM/ASA Journal on Uncertainty Quantification, 10(2):687–716, ????

2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1357615.

Kressner:2015:LRT

[KKNT15] Daniel Kressner, Rajesh Kumar, Fabio Nobile, and Christine Tobler. Low-rank tensor approximation for high-order correlation functions of Gaussian random fields. SIAM/ASA Journal on Uncertainty Quantification, 3(1):393–416, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Kahle:2019:BPI

[KLLU19] Christian Kahle, Kei Fong Lam, Jonas Latz, and Elisabeth Ullmann. Bayesian parameter identification in Cahn– Hilliard models for biological growth. SIAM/ASA Journal on Uncertainty Quantification, 7(2):526–552, ???? 2019. CO-DEN SJUQA3. ISSN 2166-2525.

Kovacs:2015:WCF

[KLS15] Mihály Kovács, Felix Lindner, and René L. Schilling. Weak convergence of finite element approximations of linear stochastic evolution equations with additive Lévy noise. SIAM/ASA Journal on Uncertainty Quantification, 3(1): 1159–1199, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Kwiatkowski:2015:CSR

[KM15] Evan Kwiatkowski and Jan Mandel. Convergence of the square root ensemble Kalman filter in the large ensemble limit. SIAM/ASA Journal on Uncertainty Quantification, 3 (1):1–17, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Kohler:2014:CPA

[KMW14] Dominic Kohler, Johannes Müller, and Utz Wever. Cellular probabilistic automata — a novel method for uncertainty propagation. SIAM/ASA Journal on Uncertainty Quantification, 2(1):29–54, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Krumscheid:2018:MMC

[KN18] S. Krumscheid and F. Nobile. Multilevel Monte Carlo approximation of functions. SIAM/ ASA Journal on Uncertainty Quantification, 6(3):1256–1293, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Kouri:2014:MSC

[Kou14]  D. P. Kouri.  A multi-

level stochastic collocation algorithm for optimization of PDEs with uncertain coefficients. SIAM/ASA Journal on Uncertainty Quantification, 2 (1):55–81, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Kamilis:2018:UQL

[KP18] Dimitris Kamilis and Nick Polydorides. Uncertainty quantification for low-frequency, time-harmonic Maxwell equations with stochastic conductivity models. SIAM/ASA Journal on Uncertainty Quantification, 6(4):1295–1334, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Kubinova:2020:BPS

[KP20] Marie Kubínová and Ivana Pultarová. Block preconditioning of stochastic Galerkin problems: New two-sided guaranteed spectral bounds. SIAM/ ASA Journal on Uncertainty Quantification, 8(1):88–113, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Klein:2022:TCS

[KPR22] Thierry Klein, Nicolas Peteilh, and Paul Rochet. Test comparison for Sobol indices over nested sets of variables. SIAM/ASA Journal on Uncertainty Quantification, 10 (4):1586–1600, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1457370.

Kretschmann:2023:MOM

[Kre23] Remo Kretschmann. Are minimizers of the Onsager–Machlup functional strong posterior modes? SIAM/ASA Journal on Uncertainty Quantification, 11(4):1105–1138, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/23M1546579.

Koley:2021:MMC

[KRS21] Ujjwal Koley, Deep Ray, and Tanmay Sarkar. Multilevel Monte Carlo finite difference methods for fractional conservation laws with random data. SIAM/ASA Journal on Uncertainty Quantification, 9(1):65– 105, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Kunoth:2016:SAT

[KS16] Angela Kunoth and Christoph Schwab. Sparse adaptive tensor Galerkin approximations of stochastic PDE-constrained control problems. SIAM/ASA Journal on Uncertainty Quantification, 4(1):1034–1059, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Kouri:2018:EOC

[KS18] D. P. Kouri and T. M. Surowiec. Existence and optimality conditions for riskaverse PDE-constrained optimization. SIAM/ASA Journal on Uncertainty Quantification, 6(2):787–815, ???? 2018. CODEN SJUQA3. ISSN 2166-2525. See corrigendum [KS22].

Kouri:2022:CEO

[KS22] Drew P. Kouri and Thomas M. Surowiec. Corrigendum: Existence and Optimality Conditions for Risk-Averse PDE-Constrained Optimization”. SIAM/ASA Journal on Uncertainty Quantification, 10 (3):1321–1322, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M143251X. See [KS18].

Khristenko:2019:ABE

[KSS^{+}19] U. Khristenko, L. Scarabosio, P. Swierczynski, E. Ullmann, and B. Wohlmuth. Analysis of boundary effects on PDE-Based sampling of Whittle– Matérn random fields. SIAM/ ASA Journal on Uncertainty Quantification, 7(3):948–974, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Kleiber:2014:MCD

[KSW14] William Kleiber, Stephan R.

Sain, and Michael J. Wiltberger. Model calibration via deformation. SIAM/ASA Journal on Uncertainty Quantification, 2(1):545–563, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Kuehn:2015:NCS

[Kue15] Christian Kuehn. Numerical continuation and SPDE stability for the 2D cubicquintic Allen–Cahn equation. SIAM/ASA Journal on Uncertainty Quantification, 3(1):762– 789, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Ko:2016:AMQ

[KW16] Jordan Ko and Henry P.

Wynn. The algebraic method in quadrature for uncertainty quantification. SIAM/ASA Journal on Uncertainty Quantification, 4(1):331–357, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Karvonen:2020:MLE

[KWT^{+}20] Toni Karvonen, George Wynne, Filip Tronarp, Chris Oates, and Simo Särkkä. Maximum likelihood estimation and uncertainty quantification for Gaussian process approximation of deterministic functions. SIAM/ASA Journal on Uncertainty Quantification, 8(3):926– 958, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Laszlo:2016:CEP

[Lás16] András László. Convergence and error propagation results on a linear iterative unfolding method. SIAM/ASA Journal on Uncertainty Quantification, 4(1):1345–1371, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Latz:2020:WPB

[Lat20] Jonas Latz. On the wellposedness of Bayesian inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 8(1):451–482, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Lee:2018:SLS

[LCE18] Kookjin Lee, Kevin Carlberg, and Howard C. Elman. Stochastic least-squares Petrov–Galerkin method for parameterized linear systems. SIAM/ASA Journal on Uncertainty Quantification, 6(1): 374–396, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

LeGratiet:2014:BAG

[LCI14] Loic Le Gratiet, Claire Cannamela, and Bertrand Iooss. A Bayesian approach for global sensitivity analysis of (multifidelity) computer codes. SIAM/ ASA Journal on Uncertainty Quantification, 2(1):336–363, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Lykkegaard:2023:MDA

[LDF^{+}23] M. B. Lykkegaard, T. J.

Dodwell, C. Fox, G. Mingas, and R. Scheichl. Multilevel delayed acceptance MCMC. SIAM/ASA Journal on Uncertainty Quantification, 11(1):1–30, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M1476770.

LeGratiet:2013:BAH

[Le 13] Loic Le Gratiet. Bayesian analysis of hierarchical multifidelity codes. SIAM/ASA Journal on Uncertainty Quantification, 1 (1):244–269, ???? 2013. CO-DEN SJUQA3. ISSN 2166-2525.

Lee:2019:MAS

[Lee19] Minyong R. Lee. Modified active subspaces using the average of gradients. SIAM/ASA Journal on Uncertainty Quantification, 7(1):53–66, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Lee:2019:LRS

[LES19] Kookjin Lee, Howard C. Elman, and Bedrich Sousedík. A low-rank solver for the Navier– Stokes equations with uncertain viscosity. SIAM/ASA Journal on Uncertainty Quantification, 7(4):1275–1300, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Labovsky:2014:EAM

[LG14] A. Labovsky and Max Gunzburger. An efficient and accurate method for the identification of the most influential random parameters appearing in the input data for PDEs. SIAM/ASA Journal on Uncertainty Quantification, 2(1):82– 105, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Liu:2017:DRG

[LG17] Xiaoyu Liu and Serge Guillas. Dimension reduction for Gaussian process emulation: an application to the influence of bathymetry on tsunami heights. SIAM/ASA Journal on Uncertainty Quantification, 5(1):787–812, ???? 2017. CO-DEN SJUQA3. ISSN 2166-2525.

Li:2021:RMD

[LHO^{+}21] Wenyu Li, Arun Hegde, James Oreluk, Andrew Packard, and Michael Frenklach. Representing model discrepancy in bound-to-bound data collaboration. SIAM/ASA Journal on Uncertainty Quantification, 9(1):231–259, ???? 2021. CO-DEN SJUQA3. ISSN 2166-2525.

Li:2022:NTS

[LK22] Moyan Li and Raed Kontar. On negative transfer and structure of latent functions in multioutput Gaussian processes. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1714–1732, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1436816.

Lopez-Lopera:2018:FDG

[LLBDR18] Andrés F. López-Lopera, François Bachoc, Nicolas Durrande, and Olivier Roustant. Finitedimensional Gaussian approximation with linear inequality constraints. SIAM/ASA Journal on Uncertainty Quantification, 6(3):1224–1255, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Li:2020:SAB

[LLS20] Qin Li, Jian-Guo Liu, and Ruiwen Shu. Sensitivity analysis of Burgers’ equation with shocks. SIAM/ASA Journal on Uncertainty Quantification, 8 (4):1493–1521, ???? 2020. CO-DEN SJUQA3. ISSN 2166-2525.

Lan:2022:SBU

[LLS22] Shiwei Lan, Shuyi Li, and Babak Shahbaba. Scaling up Bayesian uncertainty quantification for inverse problems using deep neural networks. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1684–1713, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1439456.

Liu:2017:QAG

[LLSS17] Dishi Liu, Alexander Litvinenko, Claudia Schillings, and Volker Schulz. Quantification of airfoil geometry-induced aerodynamic uncertainties — comparison of approaches. SIAM/ ASA Journal on Uncertainty Quantification, 5(1):334–352, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Luthen:2021:SPC

[LMS21] Nora Lüthen, Stefano Marelli, and Bruno Sudret. Sparse polynomial chaos expansions: Literature survey and benchmark. SIAM/ASA Journal on Uncertainty Quantification, 9(2):593– 649, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Lu:2021:ARL

[LNW21] Shuai Lu, Pingping Niu, and Frank Werner. On the asymptotical regularization for linear inverse problems in presence of white noise. SIAM/ASA Journal on Uncertainty Quantification, 9(1):1–28, ???? 2021. CO-DEN SJUQA3. ISSN 2166-2525.

Lee:2018:IMS

[LS18] Kookjin Lee and Bedrich Sousedík. Inexact methods for symmetric stochastic eigenvalue problems. SIAM/ASA Journal on Uncertainty Quantification, 6(4):1744–1776, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Lie:2018:RFM

[LST18] H. C. Lie, T. J. Sullivan, and A. L. Teckentrup. Random forward models and loglikelihoods in Bayesian inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 6(4):1600–1629, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Lu:2017:GAP

[LSW17] Yulong Lu, Andrew Stuart, and Hendrik Weber. Gaussian approximations for probability measures on R^{d}. SIAM/ASA Journal on Uncertainty Quantification, 5(1):1136–1165, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Lucini:2021:MEE

[LvLP21] María Magdalena Lucini, Peter Jan van Leeuwen, and Manuel Pulido. Model error estimation using the expectation maximization algorithm and a particle flow filter. SIAM/ASA Journal on Uncertainty Quantification, 9(2):681–707, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Li:2017:URL

[LW17] Qin Li and Li Wang. Uniform regularity for linear kinetic equations with random input based on hypocoercivity. SIAM/ASA Journal on Uncertainty Quantification, 5(1): 1193–1219, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Li:2022:IRS

[LW22] Peijun Li and Xu Wang. An inverse random source problem for the biharmonic wave equation. SIAM/ASA Journal on Uncertainty Quantification, 10(3):949–974, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1429138.

Lindsey:2022:EMC

[LWZ22] Michael Lindsey, Jonathan Weare, and Anna Zhang. Ensemble Markov chain Monte Carlo with teleporting walkers. SIAM/ASA Journal on Uncertainty Quantification, 10(3):860–885, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1425062.

Lv:2023:FCC

[LYWD23] Shurui Lv, Jun Yu, Yan Wang, and Jiang Du. Fast calibration for computer models with massive physical observations. SIAM/ASA Journal on Uncertainty Quantification, 11(3):1069–1104, ????

2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M153673X.

Law:2022:SOV

[LZ22] Kody J. H. Law and Vitaly Zankin. Sparse online variational Bayesian regression. SIAM/ASA Journal on Uncertainty Quantification, 10(3):1070–1100, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1401188.

Ma:2020:OBA

[Ma20] Pulong Ma. Objective Bayesian analysis of a cokriging model for hierarchical multifidelity codes. SIAM/ASA Journal on Uncertainty Quantification, 8 (4):1358–1382, ???? 2020. CO-DEN SJUQA3. ISSN 2166-2525.

Mazo:2021:TBE

[Maz21] Gildas Mazo. A trade-off between explorations and repetitions for estimators of two global sensitivity indices in stochastic models induced by probability measures. SIAM/ ASA Journal on Uncertainty Quantification, 9(4):1673–1713, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Madrigal-Cianci:2023:ACM

[MCNT23] Juan P. Madrigal-Cianci, Fabio Nobile, and Raúl Tempone. Analysis of a class of multilevel Markov chain Monte Carlo algorithms based on independent Metropolis–Hastings. SIAM/ ASA Journal on Uncertainty Quantification, 11(1):91–138, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1420927.

Mittal:2016:FUP

[MCTI16] A. Mittal, X. Chen, C. H. Tong, and G. Iaccarino. A flexible uncertainty propagation framework for general multiphysics systems. SIAM/ASA Journal on Uncertainty Quantification, 4(1):218–243, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Mohammadi:2022:CVB

[MCWG22] Hossein Mohammadi, Peter Challenor, Daniel Williamson, and Marc Goodfellow. Cross-Validation–based adaptive sampling for Gaussian process models. SIAM/ASA Journal on Uncertainty Quantification, 10(1):294–316, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1404260.

Mycek:2019:MMC

[MD19] Paul Mycek and Matthias De Lozzo. Multilevel Monte Carlo covariance estimation for the computation of Sobol’ indices. SIAM/ASA Journal on Uncertainty Quantification, 7(4): 1323–1348, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Moosmuller:2020:GAT

[MDK20] Caroline Moosmüller, Felix Dietrich, and Ioannis G. Kevrekidis. A geometric approach to the transport of discontinuous densities. SIAM/ ASA Journal on Uncertainty Quantification, 8(3):1012–1035, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Ming:2021:LGP

[MG21] Deyu Ming and Serge Guillas.

Linked Gaussian process emulation for systems of computer models using Matérn kernels and adaptive design. SIAM/ ASA Journal on Uncertainty Quantification, 9(4):1615–1642, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Marmin:2018:WGP

[MGBL18] Sébastien Marmin, David Ginsbourger, Jean Baccou, and Jacques Liandrat. Warped Gaussian processes and derivativebased sequential designs for functions with heterogeneous variations. SIAM/ASA Journal on Uncertainty Quantification, 6(3):991–1018, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Minh:2022:FSA

[Min22] Hà Quang Minh. Finite sample approximations of exact and entropic Wasserstein distances between covariance operators and Gaussian processes. SIAM/ ASA Journal on Uncertainty Quantification, 10(1):96–124, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1410488.

Murray:2013:DSS

[MJP13] Lawrence M. Murray, Emlyn M. Jones, and John Parslow. On disturbance statespace models and the particle marginal Metropolis–Hastings sampler. SIAM/ASA Journal on Uncertainty Quantification, 1(1):494–521, ???? 2013. CO-DEN SJUQA3. ISSN 2166-2525.

Matsuda:2021:EOD

[MM21] Takeru Matsuda and Yuto Miyatake. Estimation of ordinary differential equation models with discretization error quantification. SIAM/ASA Journal on Uncertainty Quantification, 9(1):302–331, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Malenova:2016:SSC

[MMRT16] G. Malenova, M. Motamed, O. Runborg, and R. Tempone. A sparse stochastic collocation technique for high-frequency wave propagation with uncertainty. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):1084–1110, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Martin:2021:PCO

[MN21] Matthieu Martin and Fabio Nobile. PDE-constrained optimal control problems with uncertain parameters using SAGA. SIAM/ASA Journal on Uncertainty Quantification, 9(3):979– 1012, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Morrison:2018:RMI

[MOM18] Rebecca E. Morrison, Todd A.

Oliver, and Robert D. Moser. Representing model inadequacy: a stochastic operator approach. SIAM/ASA Journal on Uncertainty Quantification, 6(2):457–496, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Morris:2018:DFM

[Mor18] Max D. Morris. Decomposing functional model inputs for variance-based sensitivity analysis. SIAM/ASA Journal on Uncertainty Quantification, 6 (4):1584–1599, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Motamed:2019:FSP

[Mot19] Mohammad Motamed. Fuzzystochastic partial differential equations. SIAM/ASA Journal on Uncertainty Quantification, 7(3):1076–1104, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Manzoni:2016:ASB

[MPL16] A. Manzoni, S. Pagani, and T. Lassila. Accurate solution of Bayesian inverse uncertainty quantification problems combining reduced basis methods and reduction error models. SIAM/ASA Journal on Uncertainty Quantification, 4(1): 380–412, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Mukherjee:2020:OCB

[MRS^{+}20] Arpan Mukherjee, Rahul Rai, Puneet Singla, Tarunraj Singh, and Abani Patra. Overlapping clustering based technique for scalable uncertainty quantification in physical systems. SIAM/ASA Journal on Uncertainty Quantification, 8(3):827– 859, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Mishra:2016:NSS

[MRST16] Siddhartha Mishra, Nils Henrik Risebro, Christoph Schwab, and Svetlana Tokareva. Numerical solution of scalar conservation laws with random flux functions. SIAM/ASA Journal on Uncertainty Quantification, 4(1):552–591, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Mohamad:2015:PDE

[MS15] Mustafa A. Mohamad and Themistoklis P. Sapsis. Probabilistic description of extreme events in intermittently unstable dynamical systems excited by correlated stochastic processes. SIAM/ASA Journal on Uncertainty Quantification, 3 (1):709–736, ???? 2015. CO-DEN SJUQA3. ISSN 2166-2525.

Mai:2017:SMO

[MS17] Chu V. Mai and Bruno Sudret. Surrogate models for oscillatory systems using sparse polynomial chaos expansions and stochastic time warping. SIAM/ASA Journal on Uncertainty Quantification, 5(1):540– 571, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Mignolet:2020:CPC

[MS20] Marc Mignolet and Christian Soize. Compressed principal component analysis of non-Gaussian vectors. SIAM/ASA Journal on Uncertainty Quantification, 8(4):1261–1286, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Mohan:2014:VMA

[MSDO14] J. Jagalur Mohan, O. Sahni, A. Doostan, and A. A. Oberai. Variational multiscale analysis: The fine-scale Green’s function for stochastic partial differential equations. SIAM/ASA Journal on Uncertainty Quantification, 2(1):397–422, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Montagna:2016:CEN

[MT16] S. Montagna and S. T. Tokdar.

Computer emulation with nonstationary Gaussian processes. SIAM/ASA Journal on Uncertainty Quantification, 4(1):26– 47, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Muller:2019:BPC

[MUL19] Christopher Müller, Sebastian Ullmann, and Jens Lang. A Bramble–Pasciak conjugate gradient method for discrete Stokes equations with random viscosity. SIAM/ASA Journal on Uncertainty Quantification, 7(3):787–805, ???? 2019. CO-DEN SJUQA3. ISSN 2166-2525.

Minunno:2013:SPB

[MvOCP13] F. Minunno, M. van Oijen, D. R. Cameron, and J. S. Pereira. Selecting parameters for Bayesian calibration of a process-based model: a methodology based on canonical correlation analysis. SIAM/ ASA Journal on Uncertainty Quantification, 1(1):370–385, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Mahmood:2015:PBE

[MWP15] Asif Mahmood, Robert L. Wolpert, and E. Bruce Pitman. A physics-based emulator for the simulation of geophysical mass flows. SIAM/ASA Journal on Uncertainty Quantification, 3(1):562–585, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Nanty:2016:SMS

[NHM^{+}16] Simon Nanty, Céline Helbert, Amandine Marrel, Nadia Pérot, and Clémentine Prieur. Sampling, metamodeling, and sensitivity analysis of numerical simulators with functional stochastic inputs. SIAM/ASA Journal on Uncertainty Quantification, 4(1):636–659, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Newsum:2017:ERB

[NP17] Craig J. Newsum and Catherine E. Powell. Efficient reduced basis methods for saddle point problems with applications in groundwater flow. SIAM/ASA Journal on Uncertainty Quantification, 5(1):1248–1278, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Nusken:2019:CSS

[NP19] N. Nüsken and G. A. Pavliotis. Constructing sampling schemes via coupling: Markov semigroups and optimal transport. SIAM/ASA Journal on Uncertainty Quantification, 7 (1):324–382, ???? 2019. CO-DEN SJUQA3. ISSN 2166-2525.

Nane:2018:ASI

[NT18] Erkan Nane and Nguyen Huy Tuan. Approximate solutions of inverse problems for nonlinear space fractional diffusion equations with randomly perturbed data. SIAM/ASA Journal on Uncertainty Quantification, 6(1):302–338, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Nickl:2020:CRP

[NvdGW20] Richard Nickl, Sara van de Geer, and Sven Wang. Convergence rates for penalized least squares estimators in PDE constrained regression problems. SIAM/ASA Journal on Uncertainty Quantification, 8(1):374– 413, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Ozen:2016:DPC

[OB16] H. Cagan Ozen and Guillaume Bal. Dynamical polynomial chaos expansions and long time evolution of differential equations with random forcing. SIAM/ASA Journal on Uncertainty Quantification, 4(1):609–635, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Oughton:2016:HEM

[OC16] Rachel H. Oughton and Peter S. Craig. Hierarchical emulation: a method for modeling and comparing nested simulators. SIAM/ASA Journal on Uncertainty Quantification, 4(1):495–519, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Owen:2017:CSB

[OCMB17] N. E. Owen, P. Challenor, P. P. Menon, and S. Bennani. Comparison of surrogate-based uncertainty quantification methods for computationally expensive simulators. SIAM/ASA Journal on Uncertainty Quantification, 5(1):403–435, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Oughton:2022:IVE

[OGH22] Rachel H. Oughton, Michael Goldstein, and John C. P. Hemmings. Intermediate variable emulation: Using internal processes in simulators to build more informative emulators. SIAM/ASA Journal on Uncertainty Quantification, 10(1):268–293, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1370902.

Owen:2021:EEA

[OH21] Art B. Owen and Christopher Hoyt. Efficient estimation of the ANOVA mean dimension, with an application to neural net classification. SIAM/ASA Journal on Uncertainty Quantification, 9(2):708–730, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Osthus:2020:PCA

[OHK^{+}20] Dave Osthus, Jeffrey D. Hyman, Satish Karra, Nishant Panda, and Gowri Srinivasan. A probabilistic clustering approach for identifying primary subnetworks of discrete fracture networks with quantified uncertainty. SIAM/ASA Journal on Uncertainty Quantification, 8(2):573–600, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Oliver:2017:MRM

[Oli17] Dean S. Oliver. Metropolized randomized maximum likelihood for improved sampling from multimodal distributions. SIAM/ASA Journal on Uncertainty Quantification, 5(1):259– 277, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Owen:2017:SVM

[OP17] Art B. Owen and Clémentine Prieur. On Shapley value for measuring importance of dependent inputs. SIAM/ASA Journal on Uncertainty Quantification, 5(1):986–1002, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Omre:2021:BSI

[OR21] Henning Omre and Kjartan Rimstad. Bayesian spatial inversion and conjugate selection Gaussian prior models. SIAM/ASA Journal on Uncertainty Quantification, 9(2):420– 445, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

OMalley:2014:CPN

[OV14] D. O’Malley and V. V. Vesselinov. A combined probabilistic… nonprobabilistic decision analysis for contaminant remediation. SIAM/ASA Journal on Uncertainty Quantification, 2(1):607–621, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Osthus:2019:PUB

[OVHW19] Dave Osthus, Scott A. Vander Wiel, Nelson M. Hoffman, and Frederick J. Wysocki. Prediction uncertainties beyond the range of experience: a case study in inertial confinement fusion implosion experiments. SIAM/ASA Journal on Uncertainty Quantification, 7(2):604– 633, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Owen:2013:VCG

[Owe13] Art B. Owen. Variance components and generalized Sobol’ indices. SIAM/ASA Journal on Uncertainty Quantification, 1(1):19–41, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Owen:2014:SIS

[Owe14] Art B. Owen. Sobol’ indices and Shapley value. SIAM/ASA Journal on Uncertainty Quantification, 2(1):245–251, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Pulch:2019:SPS

[PA19] Roland Pulch and Florian Augustin. Stability preservation in stochastic Galerkin projections of dynamical systems. SIAM/ASA Journal on Uncertainty Quantification, 7(2):634– 651, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Pampell:2014:PDT

[PAS14] A. Pampell, A. B. Aceves, and G. Srinivasan. Predicting dynamic trends of the Atlantic meridional overturning circulation for transient and stochastic forcing effects. SIAM/ASA Journal on Uncertainty Quantification, 2(1):585–606, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Prescott:2020:MAB

[PB20] Thomas P. Prescott and Ruth E. Baker. Multifidelity approximate Bayesian computation. SIAM/ASA Journal on Uncertainty Quantification, 8 (1):114–138, ???? 2020. CO-DEN SJUQA3. ISSN 2166-2525.

Prescott:2021:MAB

[PB21] Thomas P. Prescott and Ruth E. Baker. Multifidelity approximate Bayesian computation with sequential Monte Carlo parameter sampling. SIAM/ASA Journal on Uncertainty Quantification, 9 (2):788–817, ???? 2021. CO-DEN SJUQA3. ISSN 2166-2525.

Patil:2023:ROM

[PB23] Prerna Patil and Hessam Babaee. Reduced-order modeling with time-dependent bases for PDEs with stochastic boundary conditions. SIAM/ ASA Journal on Uncertainty Quantification, 11(3):727–756, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1468097.

Peherstorfer:2019:MMC

[Peh19] Benjamin Peherstorfer. Multifidelity Monte Carlo estimation with adaptive low-fidelity models. SIAM/ASA Journal on Uncertainty Quantification, 7(2): 579–603, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Picheny:2013:NST

[PG13] Victor Picheny and David Ginsbourger. A nonstationary space-time Gaussian process model for partially converged simulations. SIAM/ASA Journal on Uncertainty Quantification, 1(1):57–78, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Pham:2022:EAC

[PG22] Trung Pham and Alex A.

Gorodetsky. Ensemble approximate control variate estimators: Applications to Multi-Fidelity importance sampling. SIAM/ASA Journal on Uncertainty Quantification, 10 (3):1250–1292, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1390426.

Patil:2022:OFU

[PKH22] Pratik Patil, Mikael Kuusela, and Jonathan Hobbs. Objective frequentist uncertainty quantification for atmospheric CO_{2} retrievals. SIAM/ASA Journal on Uncertainty Quantification, 10(3):827–859, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1356403.

Perry:2019:ASH

[PKNW19] Daniel J. Perry, Robert M.

Kirby, Akil Narayan, and Ross T. Whitaker. Allocation strategies for high fidelity models in the multifidelity regime. SIAM/ASA Journal on Uncertainty Quantification, 7(1): 203–231, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Peherstorfer:2018:MPC

[PKW18] Benjamin Peherstorfer, Boris Kramer, and Karen Willcox. Multifidelity preconditioning of the cross-entropy method for rare event simulation and failure probability estimation. SIAM/ASA Journal on Uncertainty Quantification, 6(2): 737–761, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Parno:2018:TMA

[PM18] Matthew D. Parno and Youssef M. Marzouk. Transport map accelerated Markov Chain Monte Carlo. SIAM/ASA Journal on Uncertainty Quantification, 6 (2):645–682, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Portone:2022:BIU

[PM22] Teresa Portone and Robert D. Moser. Bayesian inference of an uncertain generalized diffusion operator. SIAM/ ASA Journal on Uncertainty Quantification, 10(1):151–178, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M141659X.

Parno:2016:MSB

[PMM16] Matthew Parno, Tarek Moselhy, and Youssef Marzouk. A multiscale strategy for Bayesian inference using transport maps. SIAM/ASA Journal on Uncertainty Quantification, 4(1): 1160–1190, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Pagani:2017:ESP

[PMQ17] Stefano Pagani, Andrea Manzoni, and Alfio Quarteroni. Efficient state/parameter estimation in nonlinear unsteady PDEs by a reduced basis ensemble Kalman filter. SIAM/ ASA Journal on Uncertainty Quantification, 5(1):890–921, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Patel:2021:GBP

[PO21] Dhruv V. Patel and Assad A.

Oberai. GAN-based priors for quantifying uncertainty in supervised learning. SIAM/ASA Journal on Uncertainty Quantification, 9(3):1314–1343, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Pritchard:2023:TPL

[PP23] Nathaniel Pritchard and Vivak Patel. Towards practical largescale randomized iterative least squares solvers through uncertainty quantification. SIAM/ ASA Journal on Uncertainty Quantification, 11(3):996–1024, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M1515057.

Padonou:2016:PGP

[PR16] Esperan Padonou and Olivier Roustant. Polar Gaussian processes and experimental designs in circular domains. SIAM/ ASA Journal on Uncertainty Quantification, 4(1):1014–1033, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Plischke:2021:CSE

[PRB21] Elmar Plischke, Giovanni Rabitti, and Emanuele Borgonovo. Computing Shapley effects for sensitivity analysis. SIAM/ ASA Journal on Uncertainty Quantification, 9(4):1411–1437, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Papanicolaou:2017:DRS

[PS17] Andrew Papanicolaou and Konstantinos Spiliopoulos. Dimension reduction in statistical estimation of partially observed multiscale processes. SIAM/ ASA Journal on Uncertainty Quantification, 5(1):1220–1247, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Pembery:2020:HER

[PS20] O. R. Pembery and E. A. Spence. The Helmholtz equation in random media: Wellposedness and a priori bounds. SIAM/ASA Journal on Uncertainty Quantification, 8(1):58– 87, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Perrin:2014:PEO

[PSDF14] G. Perrin, C. Soize, D. Duhamel, and C. Funfschilling. A posteriori error and optimal reduced basis for stochastic processes defined by a finite set of realizations. SIAM/ASA Journal on Uncertainty Quantification, 2(1):745–762, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Park:2017:SRT

[PT17] M. Park and M. V. Tretyakov.

Stochastic resin transfer molding process. SIAM/ASA Journal on Uncertainty Quantification, 5(1):1110–1135, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Pulch:2013:SGM

[Pul13]  Roland Pulch.  Stochastic

Galerkin methods for analyzing equilibria of random dynamical systems. SIAM/ASA Journal on Uncertainty Quantification, 1(1):408–430, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Pronzato:2020:BQE

[PZ20] Luc Pronzato and Anatoly Zhigljavsky. Bayesian quadrature, energy minimization, and space-filling design. SIAM/ ASA Journal on Uncertainty Quantification, 8(3):959–1011, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Qian:2018:MMC

[QPO^{+}18] E. Qian, B. Peherstorfer, D. O’Malley, V. V. Vesselinov, and K. Willcox. Multifidelity Monte Carlo estimation of variance and sensitivity indices. SIAM/ASA Journal on Uncertainty Quantification, 6 (2):683–706, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Rao:2021:ECE

[RA21] Vishwas Rao and Mihai Anitescu. Efficient computation of extreme excursion probabilities for dynamical systems through Rice’s formula. SIAM/ASA Journal on Uncertainty Quantification, 9(2):731–762, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Rahman:2014:GAD

[Rah14] Sharif Rahman. A generalized ANOVA dimensional decomposition for dependent probability measures. SIAM/ASA Journal on Uncertainty Quantification, 2(1):670–697, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Rahman:2016:SI

[Rah16]  Sharif Rahman.  The f-

sensitivity index. SIAM/ASA Journal on Uncertainty Quantification, 4(1):130–162, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Rahman:2018:MPP

[Rah18]  Sharif Rahman.  Mathe-

matical properties of polynomial dimensional decomposition. SIAM/ASA Journal on Uncertainty Quantification, 6 (2):816–844, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Rahman:2020:SCE

[Rah20] Sharif Rahman. A spline chaos expansion. SIAM/ASA Journal on Uncertainty Quantification, 8(1):27–57, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Rabitti:2019:SOI

[RB19] Giovanni Rabitti and Emanuele Borgonovo. A Shapley–Owen index for interaction quantification. SIAM/ASA Journal on Uncertainty Quantification, 7 (3):1060–1075, ???? 2019. CO-DEN SJUQA3. ISSN 2166-2525.

Rohrbach:2022:RBA

[RDGS22] Paul B. Rohrbach, Sergey Dolgov, Lars Grasedyck, and Robert Scheichl. Rank bounds for approximating Gaussian densities in the tensor-train format. SIAM/ASA Journal on Uncertainty Quantification, 10(3):1191–1224, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1314653.

Ruzayqat:2022:LPF

[RErB^{+}22] Hamza Ruzayqat, Aimad Erraiy, Alexandros Beskos, Dan Crisan, Ajay Jasra, and Nikolas Kantas. A lagged particle filter for stable filtering of certain high-dimensional state-space models. SIAM/ASA Journal on Uncertainty Quantification, 10(3):1130–1161, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1450392.

Rumsey:2020:DMU

[RHBH20] Kellin Rumsey, Gabriel Huerta, Justin Brown, and Lauren Hund. Dealing with measurement uncertainties as nuisance parameters in Bayesian model calibration. SIAM/ASA Journal on Uncertainty Quantification, 8(4):1287–1309, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Ray:2015:BCC

[RHH^{+}15] J. Ray, Z. Hou, M. Huang, K. Sargsyan, and L. Swiler. Bayesian calibration of the community land model using surrogates. SIAM/ASA Journal on Uncertainty Quantification, 3(1):199–233, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Rahman:2022:SDD

[RJ22] Sharif Rahman and Ramin Jahanbin. A spline dimensional decomposition for uncertainty quantification in high dimensions. SIAM/ASA Journal on Uncertainty Quantification, 10(1):404–438, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1364175.

Roosta-Khorasani:2015:ASA

[RKSA15] Farbod Roosta-Khorasani, Gábor J.

Székely, and Uri M. Ascher. Assessing stochastic algorithms for large scale nonlinear least squares problems using extremal probabilities of linear combinations of gamma random variables. SIAM/ASA Journal on Uncertainty Quantification, 3(1):61–90, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Rim:2018:DIU

[RM18] Donsub Rim and Kyle T. Mandli. Displacement interpolation using monotone rearrangement. SIAM/ASA Journal on Uncertainty Quantification, 6 (4):1503–1531, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Rim:2018:TRM

[RML18] Donsub Rim, Scott Moe, and Randall J. LeVeque. Transport reversal for model reduction of hyperbolic partial differential equations. SIAM/ASA Journal on Uncertainty Quantification, 6(1):118–150, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Roustant:2020:GKG

[RPD^{+}20] Olivier Roustant, Espéran Padonou, Yves Deville, Aloïs Clément, Guillaume Perrin, Jean Giorla, and Henry Wynn. Group kernels for Gaussian process metamodels with categorical inputs. SIAM/ASA Journal on Uncertainty Quantification, 8(2):775–806, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Resseguier:2021:QTR

[RPMC21] Valentin Resseguier, Agustin M. Picard, Etienne Memin, and Bertrand Chapron. Quantifying truncation-related uncertainties in unsteady fluid dynamics reduced order models. SIAM/ASA Journal on Uncertainty Quantification, 9(3): 1152–1183, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Rao:2015:PEE

[RS15a] Vishwas Rao and Adrian Sandu. A posteriori error estimates for the solution of variational inverse problems. SIAM/ ASA Journal on Uncertainty Quantification, 3(1):737–761, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Regier:2015:MMU

[RS15b] Jeffrey C. Regier and Philip B. Stark. Mini-minimax uncertainty quantification for emulators. SIAM/ASA Journal on Uncertainty Quantification, 3 (1):686–708, ???? 2015. CO-DEN SJUQA3. ISSN 2166-2525.

Romer:2016:SMR

[RSW16] Ulrich Römer, Sebastian Schöps, and Thomas Weiland. Stochastic modeling and regularity of the nonlinear elliptic curl–curl equation. SIAM/ASA Journal on Uncertainty Quantification, 4(1):952–979, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Ren:2020:CIM

[RV20] Kui Ren and Sarah Vallélian.

Characterizing impacts of model uncertainties in quantitative photoacoustics. SIAM/ ASA Journal on Uncertainty Quantification, 8(2):636–667, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Reich:2021:FPP

[RW21] Sebastian Reich and Simon Weissmann. Fokker–Planck particle systems for Bayesian inference: Computational approaches. SIAM/ASA Journal on Uncertainty Quantification, 9(2):446–482, ???? 2021. CO-DEN SJUQA3. ISSN 2166-2525.

Sanz-Alonso:2018:ISN

[SA18] Daniel Sanz-Alonso. Importance sampling and necessary sample size: an information theory approach. SIAM/ASA Journal on Uncertainty Quantification, 6(2):867–879, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Sanz-Alonso:2015:LTA

[SAS15] Daniel Sanz-Alonso and Andrew M. Stuart. Long-time asymptotics of the filtering distribution for partially observed chaotic dynamical systems. SIAM/ASA Journal on Uncertainty Quantification, 3 (1):1200–1220, ???? 2015. CO-DEN SJUQA3. ISSN 2166-2525.

Sanz-Alonso:2022:FER

[SAY22] Daniel Sanz-Alonso and Ruiyi Yang. Finite element representations of Gaussian processes: Balancing numerical and statistical accuracy. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1323–1349, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M144788X.

Spiller:2014:AEC

[SBB^{+}14] Elaine T. Spiller, M. J. Bayarri, James O. Berger, Eliza S. Calder, Abani K. Patra, E. Bruce Pitman, and Robert L. Wolpert. Automating emulator construction for geophysical hazard maps. SIAM/ASA Journal on Uncertainty Quantification, 2(1): 126–152, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Saibaba:2019:EMB

[SBBA19] Arvind K. Saibaba, Johnathan Bardsley, D. Andrew Brown, and Alen Alexanderian. Efficient marginalization-based MCMC methods for hierarchical Bayesian inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 7(3): 1105–1131, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Sung:2022:CIC

[SBW22] Chih-Li Sung, Beau David Barber, and Berkley J. Walker. Calibration of inexact computer models with heteroscedastic errors. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1733–1752, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1417946.

Severino:2020:UQU

[SC20] Gerardo Severino and Salvatore Cuomo. Uncertainty quantification of unsteady flows generated by line-sources through heterogeneous geological formations. SIAM/ASA Journal on Uncertainty Quantification, 8 (2):807–825, ???? 2020. CO-DEN SJUQA3. ISSN 2166-2525.

Scarabosio:2022:DNN

[Sca22] Laura Scarabosio. Deep neural network surrogates for nonsmooth quantities of interest in shape uncertainty quantification. SIAM/ASA Journal on Uncertainty Quantification, 10(3):975–1011, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1393078.

Sinsbeck:2021:SDC

[SCN21] Michael Sinsbeck, Emily Cooke, and Wolfgang Nowak. Sequential design of computer experiments for the computation of Bayesian model evidence. SIAM/ASA Journal on Uncertainty Quantification, 9(1): 260–279, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Sousedik:2016:ISI

[SE16] Bedrich Sousedík and Howard C. Elman. Inverse subspace iteration for spectral stochastic finite element methods. SIAM/ ASA Journal on Uncertainty Quantification, 4(1):163–189, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Staber:2015:ASL

[SG15] B. Staber and J. Guilleminot. Approximate solutions of Lagrange multipliers for informationtheoretic random field models. SIAM/ASA Journal on Uncertainty Quantification, 3(1): 599–621, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Sun:2019:ESD

[SGH^{+}19] Furong Sun, Robert B. Gramacy, Benjamin Haaland, Earl Lawrence, and Andrew Walker. Emulating satellite drag from large simulation experiments. SIAM/ASA Journal on Uncertainty Quantification, 7(2):720– 759, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Shulkind:2018:EDN

[SHA18] Gal Shulkind, Lior Horesh, and Haim Avron. Experimental design for nonparametric correction of misspecified dynamical models. SIAM/ASA Journal on Uncertainty Quantification, 6(2):880–906, ???? 2018. CO-DEN SJUQA3. ISSN 2166-2525.

Schick:2014:NGM

[SHL14] M. Schick, V. Heuveline, and O. P. Le Maître. A Newton– Galerkin method for fluid flow exhibiting uncertain periodic dynamics. SIAM/ASA Journal on Uncertainty Quantification, 2(1):153–173, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Sharrock:2022:JOP

[SK22] Louis Sharrock and Nikolas Kantas. Joint online parameter estimation and optimal sensor placement for the partially observed stochastic advection– diffusion equation. SIAM/ASA Journal on Uncertainty Quantification, 10(1):55–95, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1375073.

Sousedik:2022:SGM

[SL22] Bedrich Sousedík and Kookjin Lee. Stochastic Galerkin methods for linear stability analysis of systems with parametric uncertainty. SIAM/ASA Journal on Uncertainty Quantification, 10(3):1101–1129, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1415595.

Spagnol:2019:GSA

[SLD19] Adrien Spagnol, Rodolphe Le Riche, and Sébastien Da Veigao. Global sensitivity analysis for optimization with variable selection. SIAM/ASA Journal on Uncertainty Quantification, 7(2):417–443, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Scott:2019:AZR

[SLH19] James M. Scott, M. Paul Laiu, and Cory D. Hauck. Analysis of the zero relaxation limit of hyperbolic balance laws with random initial data. SIAM/ASA Journal on Uncertainty Quantification, 7(3):806–837, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Sinsbeck:2017:SDC

[SN17] Michael Sinsbeck and Wolfgang Nowak. Sequential design of computer experiments for the solution of Bayesian inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 5(1):640–664, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Seshadri:2017:ESQ

[SNM17] Pranay Seshadri, Akil Narayan, and Sankaran Mahadevan. Effectively subsampled quadratures for least squares polynomial approximations. SIAM/ ASA Journal on Uncertainty Quantification, 5(1):1003–1023, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Song:2016:SEG

[SNS16] Eunhye Song, Barry L. Nelson, and Jeremy Staum. Shapley effects for global sensitivity analysis: Theory and computation. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):1060–1083, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Strong:2014:WMG

[SO14] Mark Strong and Jeremy E.

Oakley. When is a model good enough? Deriving the expected value of model improvement via specifying internal model discrepancies. SIAM/ASA Journal on Uncertainty Quantification, 2(1):106–125, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Soize:2015:PCE

[Soi15] C. Soize. Polynomial chaos expansion of a multimodal random vector. SIAM/ASA Journal on Uncertainty Quantification, 3(1):34–60, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Soize:2017:OPT

[Soi17] C. Soize. Optimal partition in terms of independent random vectors of any non-Gaussian vector defined by a set of realizations. SIAM/ASA Journal on Uncertainty Quantification, 5(1):176–211, ???? 2017. CODEN SJUQA3. ISSN 2166-

Silvester:2016:OSL

[SP16] David Silvester and Pranjal. An optimal solver for linear systems arising from stochastic FEM approximation of diffusion equations with random coefficients. SIAM/ASA Journal on Uncertainty Quantification, 4(1):298–311, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Sun:2021:RSD

[SPHJ21] Wenbo Sun, Matthew Plumlee, Jingwen Hu, and Jionghua (Judy) Jin. Robust system design with limited experimental data and an inexact simulation model. SIAM/ASA Journal on Uncertainty Quantification, 9(2): 483–506, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Spiliopoulos:2020:IGA

[Spi20] Konstantinos Spiliopoulos. Information geometry for approximate Bayesian computation. SIAM/ASA Journal on Uncertainty Quantification, 8(1):229– 260, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Spokoiny:2023:DFN

[Spo23] Vladimir Spokoiny. Dimension free nonasymptotic bounds on the accuracy of highdimensional Laplace approximation. SIAM/ASA Journal on Uncertainty Quantification, 11(3):1044–1068, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M1495688.

Stanhope:2017:RSI

[SRS17] S. Stanhope, J. E. Rubin, and D. Swigon. Robustness of solutions of the inverse problem for linear dynamical systems with uncertain data. SIAM/ASA Journal on Uncertainty Quantification, 5(1):572–597, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Stein:2013:FNH

[SS13] Jerome L. Stein and Seth Stein. Formulating natural hazard policies under uncertainty. SIAM/ASA Journal on Uncertainty Quantification, 1(1):42– 56, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Sturlaugson:2015:SAC

[SS15] Liessman Sturlaugson and John W. Sheppard. Sensitivity analysis of continuous time Bayesian network reliability models. SIAM/ASA Journal on Uncertainty Quantification, 3(1):346–369, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Singh:2013:PME

[SSJ^{+}13] K. Singh, A. Sandu, M. Jardak, K. W. Bowman, and M. Lee. A practical method to estimate information content in the context of 4D-var data assimilation. SIAM/ASA Journal on Uncertainty Quantification, 1 (1):106–138, ???? 2013. CO-DEN SJUQA3. ISSN 2166-2525.

Scheichl:2017:QMC

[SST17] R. Scheichl, A. M. Stuart, and A. L. Teckentrup. Quasi-Monte Carlo and multilevel Monte Carlo methods for computing posterior expectations in elliptic inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 5(1):493– 518, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Swigon:2019:IJD

[SSZR19] David Swigon, Shelby R. Stanhope, Sven Zenker, and Jonathan E. Rubin. On the importance of the Jacobian determinant in parameter inference for random parameter and random measurement error models. SIAM/ASA Journal on Uncertainty Quantification, 7(3): 975–1006, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Sinsbeck:2015:IDA

[ST15] Michael Sinsbeck and Daniel M. Tartakovsky. Impact of data assimilation on cost-accuracy tradeoff in multifidelity models. SIAM/ASA Journal on Uncertainty Quantification, 3(1): 954–968, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Su:2017:SIS

[ST17] Chen Su and Xuemin Tu. Sequential implicit sampling methods for Bayesian inverse problems. SIAM/ASA Journal on Uncertainty Quantification, 5(1):519–539, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Stark:2015:CVP

[Sta15] Philip B. Stark. Constraints versus priors. SIAM/ASA Journal on Uncertainty Quantification, 3(1):586–598, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Schillings:2023:EBG

[STW23] Claudia Schillings, Claudia Totzeck, and Philipp Wacker. Ensemble-based gradient inference for particle methods in optimization and sampling. SIAM/ASA Journal on Uncertainty Quantification, 11(3):757–787, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M1533281.

Schaden:2020:MBL

[SU20] Daniel Schaden and Elisabeth Ullmann. On multilevel best linear unbiased estimators. SIAM/ASA Journal on Uncertainty Quantification, 8(2): 601–635, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Schaden:2021:AAM

[SU21] Daniel Schaden and Elisabeth Ullmann. Asymptotic analysis of multilevel best linear unbiased estimators. SIAM/ASA Journal on Uncertainty Quantification, 9(3):953–978, ????

2021. CODEN SJUQA3. ISSN 2166-2525.

Spence:2023:WEP

[SW23] E. A. Spence and J. Wunsch. Wavenumber-explicit parametric holomorphy of Helmholtz solutions in the context of uncertainty quantification. SIAM/ASA Journal on Uncertainty Quantification, 11(2):567–590, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M1486170.

Salter:2022:QST

[SWGE22] James M. Salter, Daniel B. Williamson, Lauren J. Gregoire, and Tamsin L. Edwards. Quantifying spatiotemporal boundary condition uncertainty for the North American deglaciation. SIAM/ ASA Journal on Uncertainty Quantification, 10(2):717–744, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1409135.

Spiller:2023:ZPG

[SWTA23] Elaine T. Spiller, Robert L. Wolpert, Pablo Tierz, and Taylor G. Asher. The zero problem: Gaussian process emulators for range-constrained computer models. SIAM/ ASA Journal on Uncertainty Quantification, 11(2):540–566, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1467420.

Seshadri:2019:DRG

[SYP19] Pranay Seshadri, Shaowu Yuchi, and Geoffrey T. Parks. Dimension reduction via Gaussian ridge functions. SIAM/ ASA Journal on Uncertainty Quantification, 7(4):1301–1322, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Sawko:2022:EGC

[SZ22] Robert Sawko and Małgorzata J.

Zimo’n. Effective generation of compressed stationary Gaussian fields. SIAM/ ASA Journal on Uncertainty Quantification, 10(1):439–452, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1375541.

Schwab:2023:DLH

[SZ23] Christoph Schwab and Jakob Zech. Deep learning in high dimension: Neural network expression rates for analytic functions in L^{2}(R^{d}, γ_{d}). SIAM/ ASA Journal on Uncertainty Quantification, 11(1):199–234, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1462738.

Tan:2015:SBP

[Tan15] Matthias Hwai Yong Tan. Sequential Bayesian polynomial chaos model selection for estimation of sensitivity indices.

SIAM/ASA Journal on Uncertainty Quantification, 3(1):146– 168, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Tan:2019:GPM

[Tan19] Matthias H. Y. Tan. Gaussian process modeling of finite element models with functional inputs. SIAM/ASA Journal on Uncertainty Quantification, 7 (4):1133–1161, ???? 2019. CO-DEN SJUQA3. ISSN 2166-2525.

Tan:2020:BOE

[Tan20] Matthias H. Y. Tan. Bayesian optimization of expected quadratic loss for multiresponse computer experiments with internal noise. SIAM/ASA Journal on Uncertainty Quantification, 8(3):891–925, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Torlo:2018:SWR

[TBR18] Davide Torlo, Francesco Ballarin, and Gianluigi Rozza. Stabilized weighted reduced basis methods for parametrized advection dominated problems with random inputs. SIAM/ ASA Journal on Uncertainty Quantification, 6(4):1475–1502, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Teckentrup:2020:CGP

[Tec20] Aretha L. Teckentrup. Convergence of Gaussian process regression with estimated hyperparameters and applications in Bayesian inverse problems.

SIAM/ASA Journal on Uncertainty Quantification, 8(4): 1310–1337, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Thorarinsdottir:2013:UPD

[TGG13] Thordis L. Thorarinsdottir, Tilmann Gneiting, and Nadine Gissibl. Using proper divergence functions to evaluate climate models. SIAM/ASA Journal on Uncertainty Quantification, 1(1):522–534, ???? 2013. CODEN SJUQA3. ISSN 2166-2525.

Tsilifis:2017:EBE

[TGH17] Panagiotis Tsilifis, Roger G.

Ghanem, and Paris Hajali. Efficient Bayesian experimentation using an expected information gain lower bound. SIAM/ASA Journal on Uncertainty Quantification, 5(1):30– 62, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Travelletti:2023:UQE

[TGL23] Cédric Travelletti, David Ginsbourger, and Niklas Linde. Uncertainty quantification and experimental design for largescale linear inverse problems under Gaussian process priors. SIAM/ASA Journal on Uncertainty Quantification, 11(1):168–198, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1445028.

Thorbergsson:2018:EDP

[TH18] Leifur Thorbergsson and Giles Hooker. Experimental design for partially observed Markov decision processes. SIAM/ASA Journal on Uncertainty Quantification, 6(2):549–567, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Tang:2014:SGQ

[TI14] Gary Tang and Gianluca Iaccarino. Subsampled Gauss quadrature nodes for estimating polynomial chaos expansions. SIAM/ASA Journal on Uncertainty Quantification, 2 (1):423–443, ???? 2014. CO-DEN SJUQA3. ISSN 2166-2525.

Teckentrup:2015:MSC

[TJWG15] A. L. Teckentrup, P. Jantsch, C. G. Webster, and M. Gunzburger. A multilevel stochastic collocation method for partial differential equations with random input data. SIAM/ASA Journal on Uncertainty Quantification, 3(1):1046–1074, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Taghvaei:2020:DMB

[TMM20] Amirhossein Taghvaei, Prashant G. Mehta, and Sean P. Meyn. Diffusion map-based algorithm for gain function approximation in the feedback particle filter. SIAM/ASA Journal on Uncertainty Quantification, 8(3): 1090–1117, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Tong:2023:LDT

[TS23] Shanyin Tong and Georg Stadler. Large deviation theory-based adaptive importance sampling for rare events in high dimensions. SIAM/ ASA Journal on Uncertainty Quantification, 11(3):788–813, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/22M1524758.

Tipireddy:2018:SBA

[TST18] R. Tipireddy, P. Stinis, and A. M. Tartakovsky. Stochastic basis adaptation and spatial domain decomposition for partial differential equations with random coefficients. SIAM/ASA Journal on Uncertainty Quantification, 6(1):273–301, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Tuo:2019:ACM

[Tuo19] Rui Tuo. Adjustments to computer models via projected kernel calibration. SIAM/ASA Journal on Uncertainty Quantification, 7(2):553–578, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Tuo:2016:TFC

[TW16] Rui Tuo and C. F. Jeff Wu.

A theoretical framework for calibration in computer models: Parametrization, estimation and convergence properties. SIAM/ASA Journal on Uncertainty Quantification, 4 (1):767–795, ???? 2016. CO-DEN SJUQA3. ISSN 2166-2525.

Tuo:2020:IRC

[TWW20] Rui Tuo, Yan Wang, and C. F. Jeff Wu. On the improved rates of convergence for Matérn-type kernel ridge regression with application to calibration of computer models. SIAM/ASA Journal on Uncertainty Quantification, 8(4): 1522–1547, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Uribe:2022:HGS

[UBD^{+}22] Felipe Uribe, Johnathan M. Bardsley, Yiqiu Dong, Per Christian Hansen, and Nicolai A. B. Riis. A hybrid Gibbs sampler for edge-preserving tomographic reconstruction with uncertain view angles. SIAM/ASA Journal on Uncertainty Quantification, 10 (3):1293–1320, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1412268.

Ullmann:2015:MER

[UP15a] Elisabeth Ullmann and Iason Papaioannou. Multilevel estimation of rare events. SIAM/ ASA Journal on Uncertainty Quantification, 3(1):922–953, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Ullmann:2015:SLT

[UP15b] Elisabeth Ullmann and Catherine E. Powell. Solving logtransformed random diffusion problems by stochastic Galerkin mixed finite element methods. SIAM/ASA Journal on Uncertainty Quantification, 3(1):509–534, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

Uribe:2021:CEB

[UPMS21] Felipe Uribe, Iason Papaioannou, Youssef M. Marzouk, and Daniel Straub. Cross-entropy-based importance sampling with failure-informed dimension reduction for rare event simulation. SIAM/ASA Journal on Uncertainty Quantification, 9(2):818–847, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

vanAckooij:2017:SGF

[vAH17] Wim van Ackooij and R. Henrion. (sub-)gradient formulae for probability functions of random inequality systems under Gaussian distribution. SIAM/ASA Journal on Uncertainty Quantification, 5(1):63– 87, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Vidal-Codina:2016:EIM

[VCNGP16] F. Vidal-Codina, N. C. Nguyen, M. B. Giles, and J. Peraire. An empirical interpolation and model-variance reduction method for computing statistical outputs of parametrized stochastic partial differential equations. SIAM/ASA Journal on Uncertainty Quantification, 4(1):244–265, ???? 2016.

CODEN SJUQA3. ISSN 2166-2525.

vandenBos:2020:GNQ

[vdBSBvB20] Laurent van den Bos, Benjamin Sanderse, Wim Bierbooms, and Gerard van Bussel. Generating nested quadrature rules with positive weights based on arbitrary sample sets. SIAM/ASA Journal on Uncertainty Quantification, 8(1):139–169, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Vermiglio:2017:PCE

[Ver17]  Rossana Vermiglio.  Poly-

nomial chaos expansions for the stability analysis of uncertain delay differential equations. SIAM/ASA Journal on Uncertainty Quantification, 5 (1):278–303, ???? 2017. CO-DEN SJUQA3. ISSN 2166-2525.

Vernon:2019:KBE

[VJC19] Ian Vernon, Samuel E. Jackson, and Jonathan A. Cumming. Known boundary emulation of complex computer models. SIAM/ASA Journal on Uncertainty Quantification, 7(3): 838–876, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Vondrejc:2019:ACC

[VM19] Jaroslav Vondrejc and Hermann G. Matthies. Accurate computation of conditional expectation for highly nonlinear problems. SIAM/ASA Journal on Uncertainty Quantification, 7(4):1349–1368, ???? 2019.

CODEN SJUQA3. ISSN 2166-2525.

Vollmer:2015:DIM

[Vol15] Sebastian J. Vollmer. Dimensionindependent MCMC sampling for inverse problems with non-Gaussian priors. SIAM/ASA Journal on Uncertainty Quantification, 3(1):535–561, ???? 2015. CODEN SJUQA3. ISSN 2166-2525.

VanBarel:2019:ROP

[VV19] Andreas Van Barel and Stefan Vandewalle. Robust optimization of PDEs with random coefficients using a multilevel Monte Carlo method. SIAM/ASA Journal on Uncertainty Quantification, 7(1):174– 202, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Volodina:2020:DDN

[VW20] Victoria Volodina and Daniel Williamson. Diagnostics-driven nonstationary emulators using kernel mixtures. SIAM/ASA Journal on Uncertainty Quantification, 8(1):1–26, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

vanWyk:2015:PLN

[vWGBS15] Hans-Werner van Wyk, Max Gunzburger, John Burkhardt, and Miroslav Stoyanov. Powerlaw noises over general spatial domains and on nonstandard meshes. SIAM/ASA Journal on Uncertainty Quantification, 3(1):296–319, ???? 2015. CO-DEN SJUQA3. ISSN 2166-2525.

Wallstrom:2017:AMI

[Wal17] Timothy C. Wallstrom. On the application of McDiarmid’s inequality to complex systems. SIAM/ASA Journal on Uncertainty Quantification, 5(1):240– 245, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Wang:2022:PPK

[Wan22] Yan Wang. Penalized projected kernel calibration for computer models. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1652–1683, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1387614.

Williamson:2014:EBE

[WB14] Daniel Williamson and Adam T. Blaker. Evolving Bayesian emulators for structured chaotic time series, with application to large climate models. SIAM/ ASA Journal on Uncertainty Quantification, 2(1):1–28, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Wang:2016:ESC

[WB16] Xiaojing Wang and James O. Berger. Estimating shape constrained functions using Gaussian processes. SIAM/ASA Journal on Uncertainty Quantification, 4(1):1–25, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Wang:2015:PDF

[WBSC^{+}15] P. Wang, D. A. Barajas-Solano, E. Constantinescu, S. Abhyankar, D. Ghosh, B. F. Smith, Z. Huang, and A. M. Tartakovsky. Probabilistic density function method for stochastic ODEs of power systems with uncertain power input. SIAM/ASA Journal on Uncertainty Quantification, 3 (1):873–896, ???? 2015. CO-DEN SJUQA3. ISSN 2166-2525.

Wu:2023:FSC

[WCG23] Keyi Wu, Peng Chen, and Omar Ghattas. A fast and scalable computational framework for large-scale high-dimensional Bayesian optimal experimental design. SIAM/ASA Journal on Uncertainty Quantification, 11(1):235–261, ???? 2023. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1466499.

Wang:2022:PWG

[WCL22] Yifei Wang, Peng Chen, and Wuchen Li. Projected Wasserstein gradient descent for high-dimensional Bayesian inference. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1513–1532, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1454018.

Wilson:2022:VCM

[WGD22] Amy L. Wilson, Michael Goldstein, and Chris J. Dent. Varying coefficient models and design choice for Bayes linear emulation of complex computer models with limited model evaluations. SIAM/ ASA Journal on Uncertainty Quantification, 10(1):350–378, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1318560.

Wagner:2022:EKF

[WPU22] Fabian Wagner, I. Papaioannou, and E. Ullmann. The ensemble Kalman filter for rare event estimation. SIAM/ ASA Journal on Uncertainty Quantification, 10(1):317–349, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1404119.

Wang:2016:EGT

[WR16] Ting Wang and Muruhan Rathinam. Efficiency of the Girsanov transformation approach for parametric sensitivity analysis of stochastic chemical kinetics. SIAM/ASA Journal on Uncertainty Quantification, 4(1):1288–1322, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Wentworth:2016:PSV

[WSB16] Mami T. Wentworth, Ralph C. Smith, and H. T. Banks. Parameter selection and verification techniques based on global sensitivity analysis illustrated for an HIV model. SIAM/ASA Journal on Uncertainty Quantification, 4(1):266–297, ???? 2016. CODEN SJUQA3. ISSN 2166-2525.

Wen:2020:DMP

[WWLL20] Linjie Wen, Jiangqi Wu, Linjun Lu, and Jinglai Li. A defensive marginal particle filtering method for data assimilation. SIAM/ASA Journal on Uncertainty Quantification, 8 (3):1215–1235, ???? 2020. CO-DEN SJUQA3. ISSN 2166-2525.

Wang:2022:GPI

[WYHW22] Wenjia Wang, Xiaowei Yue, Benjamin Haaland, and C. F. Jeff Wu. Gaussian processes with input location error and applications to the composite parts assembly process. SIAM/ ASA Journal on Uncertainty Quantification, 10(2):619–650, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/20M1312447.

Xu:2014:RES

[XLL14] Gongjun Xu, Guang Lin, and Jingchen Liu. Rare-event simulation for the stochastic Korteweg–de Vries equation. SIAM/ASA Journal on Uncertainty Quantification, 2(1): 698–716, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Xie:2021:NBF

[XLWZ21] Wei Xie, Cheng Li, Yuefeng Wu, and Pu Zhang. A nonparametric Bayesian framework for uncertainty quantification in stochastic simulation. SIAM/ ASA Journal on Uncertainty Quantification, 9(4):1527–1552, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Xiao:2021:EEI

[XMLD21] Qian Xiao, Abhyuday Mandal, C. Devon Lin, and Xinwei Deng. EzGP: Easy-to-interpret Gaussian process models for computer experiments with both quantitative and qualitative factors. SIAM/ASA Journal on Uncertainty Quantification, 9(2):333–353, ???? 2021. CODEN SJUQA3. ISSN 2166-2525.

Xu:2017:MLE

[XS17] Wanting Xu and Michael L. Stein. Maximum likelihood estimation for a smooth Gaussian random field model. SIAM/ ASA Journal on Uncertainty Quantification, 5(1):138–175, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Xu:2019:MPC

[XSW19] Wanting Xu, Michael L. Stein, and Ian Wisher. Modeling and predicting chaotic circuit data. SIAM/ASA Journal on Uncertainty Quantification, 7(1):31– 52, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Yang:2018:SIR

[YLT18] Xiu Yang, Weixuan Li, and Alexandre Tartakovsky. Sliced-Inverse-Regression–Aided rotated compressive sensing method for uncertainty quantification. SIAM/ASA Journal on Uncertainty Quantification, 6(4):1532–1554, ???? 2018. CODEN SJUQA3. ISSN 2166-2525.

Yang:2020:RCM

[YWT20] Lun Yang, Peng Wang, and Daniel M. Tartakovsky. Resourceconstrained model selection for uncertainty propagation and data assimilation. SIAM/ASA Journal on Uncertainty Quantification, 8(3):1118–1138, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Yamada:2019:SOD

[YY19] Toshihiro Yamada and Kenta Yamamoto. Second order discretization of Bismut– Elworthy–Li formula: Application to sensitivity analysis. SIAM/ASA Journal on Uncertainty Quantification, 7(1): 143–173, ???? 2019. CODEN SJUQA3. ISSN 2166-2525.

Yu:2020:INE

[YZAJ20] Yang Yu, Ning Zhang, Daniel W.

Apley, and Wenxin Jiang. Including a nugget effect in lifted Brownian covariance models. SIAM/ASA Journal on Uncertainty Quantification, 8(4): 1338–1357, ???? 2020. CODEN SJUQA3. ISSN 2166-2525.

Zahr:2019:EGC

[ZCK19] Matthew J. Zahr, Kevin T. Carlberg, and Drew P. Kouri. An efficient, globally convergent method for optimization under uncertainty using adaptive model reduction and sparse grids. SIAM/ASA Journal on Uncertainty Quantification, 7 (3):877–912, ???? 2019. CO-DEN SJUQA3. ISSN 2166-2525.

Zhang:2022:EBI

[ZGS22] Jiahui Zhang, Anne Gelb, and Theresa Scarnati. Empirical Bayesian inference using a support informed prior. SIAM/ ASA Journal on Uncertainty Quantification, 10(2):745–774, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M140794X.

Zhao:2021:LBM

[ZHY21] Weifeng Zhao, Juntao Huang, and Wen-An Yong. Lattice Boltzmann method for stochastic convection–diffusion equations. SIAM/ASA Journal on Uncertainty Quantification, 9 (2):536–563, ???? 2021. CO-DEN SJUQA3. ISSN 2166-2525.

Zhou:2017:HAM

[ZHYL17] Qingping Zhou, Zixi Hu, Zhewei Yao, and Jinglai Li. A hybrid adaptive MCMC algorithm in function spaces. SIAM/ASA Journal on Uncertainty Quantification, 5(1):

621–639, ???? 2017. CODEN SJUQA3. ISSN 2166-2525.

Zou:2022:LAR

[ZKA22] Zilong Zou, Drew P. Kouri, and Wilkins Aquino. A locally adapted reduced-basis method for solving risk-averse PDE-constrained optimization problems. SIAM/ASA Journal on Uncertainty Quantification, 10(4):1629–1651, ???? 2022. CODEN SJUQA3. ISSN 2166-2525. URL https:

//epubs.siam.org/doi/10.
1137/21M1411342.

Zhang:2019:SDA

[ZLR19] Ru Zhang, C. Devon Lin, and Pritam Ranjan. A sequential design approach for calibrating dynamic computer simulators. SIAM/ASA Journal on Uncertainty Quantification, 7 (4):1245–1274, ???? 2019. CO-DEN SJUQA3. ISSN 2166-2525.

Zhu:2014:CAS

[ZNX14] Xueyu Zhu, Akil Narayan, and Dongbin Xiu. Computational aspects of stochastic collocation with multifidelity models. SIAM/ASA Journal on Uncertainty Quantification, 2(1):444– 463, ???? 2014. CODEN SJUQA3. ISSN 2166-2525.

Zhu:2021:ESS

[ZS21] Xujia Zhu and Bruno Sudret.

Emulation of stochastic simulators using generalized lambda models. SIAM/ASA Journal on Uncertainty Quantification, 9 (4):1345–1380, ???? 2021. CO-DEN SJUQA3. ISSN 2166-2525.